### v3

# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

Abstract.
The isoperimetric problem is a classic topic in geometric measure theory, yet critical questions regarding the characterization of optimal solutions—even asymptotically optimal ones—remain largely unresolved. In this paper, we investigate the large deviations asymptotics (governing sequences of exponentially small sets) for the isoperimetric problem on the product Riemannian manifold $\mathcal{M}^{n}$ endowed with the product probability measure $\nu^{\otimes n}$, where $\mathcal{M}$ is a Riemannian manifold satisfying the curvature-dimension condition $\mathrm{CD}(0,\infty)$. When the probability measure $\nu$} admits a finite moment generating function for squared distance, we establish an exact characterization of the large deviations asymptotics of the isoperimetric profile, which reveals a precise equivalence between this asymptotic isoperimetric inequality and an established strengthening of log-Sobolev inequality—namely, the nonlinear log-Sobolev inequality. It is observed that conditional typical sets (i.e., $\mathbf{f}$-coordinate half-spaces for some vector $\mathbf{f}$ of functions), a fundamental concept from information theory, form an asymptotically optimal solution to the isoperimetric problem. This class of subsets further yields an upper bound on the isoperimetric profile in the central limit regime (concerning constant-volume sets), expressed as the Gaussian isoperimetric profile times the square root of the spectral gap. Furthermore, we apply our results to establish quantitative relations among the optimal constants in isoperimetric, concentration and transport inequalities, additionally covering isoperimetry for subexponentially and superexponentially small sets. Our results provide a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers behave fundamentally differently across spaces with distinct geometric structures. We adopt a new proof idea in this paper which integrates tools from information theory, optimal transport, and geometric measure theory.
Key words and phrases:
Isoperimetric inequality, large deviations, log-Sobolev inequality, information theory, optimal transport, Wasserstein gradient flow, geometric measure theory
2020 Mathematics Subject Classification:
60E15, 52B60, 39B62, 60F10
L. Yu is with the School of Statistics and Data Science, LPMC, KLMDASR, and LEBPS, Nankai University, Tianjin 300071, China (e-mail: leiyu@nankai.edu.cn).
1. Introduction
In a metric-measure space, the isoperimetric problem seeks the subset of fixed measure that minimizes surface area—where surface area is rigorously defined via the boundary measure. This classic problem remains largely unresolved in both its exact and asymptotic formulations, with the explicit characterization of the isoperimetric profile (the minimal surface area corresponding to each fixed measure) and its asymptotic behavior posing a widely open question, save for a handful of specific cases: e.g., the standard Gaussian measure, log-concave measures on the real line, and Euclidean spheres endowed with constant density. In this paper, we study the isoperimetric problem across various asymptotic regimes, with particular emphasis on the large deviations regime that considers sequences of exponentially small sets. We formalize our core analytical setup as follows.
Let $(\mathcal{M},g)$ be a finite-dimensional smooth complete oriented connected Riemannian manifold with induced geodesic distance $d$. Let $\nu=e^{-V}\mathrm{vol}$ with $V\in C^{2}(\mathcal{M})$ be a reference Borel probability measure on $\mathcal{M}$ that is absolutely continuous with respect to the Riemannian volume $\mathrm{vol}$. We consider the $n$-fold product space $\mathcal{M}^{n}$, equipped with the product measure $\nu_{n}:=\nu^{\otimes n}$ and the product distance $d_{n}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$, where $\mathbf{x}=(x_{1},x_{2},...,x_{n})$.
For a set $A\subseteq\mathcal{M}^{n}$, denote its $r$-enlargement as
(1.1) $A^{r}:=\bigcup_{\mathbf{x}\in A}\{\mathbf{y}\in\mathcal{M}^{n}:d_{n}(\mathbf{x},\mathbf{y})\leq r\}.$
For a Borel set $A\subseteq\mathcal{M}^{n}$, its boundary measure (i.e., perimeter) is defined by
(1.2) $\nu_{n}^{+}(A):=\liminf_{r\downarrow 0}\frac{\nu_{n}(A^{r})-\nu_{n}(A)}{r}.$
In the isoperimetric problem, the objective is to minimize $\nu_{n}^{+}(A)$ over all sets $A$ of a given measure, i.e., to determine the isoperimetric profile: For $a\in[0,1]$,
(1.3) $I_{n}(a):=\inf_{A:\nu_{n}(A)=a}\nu_{n}^{+}(A),$
where $A$ is a Borel set. Note that there are several other ways in the literature to define the perimeter of a set, but they coincide when the set admits smooth boundary, and thus, lead to the same isoperimetric profile in our setup; see details in [10, 5, pp. 32-33].
Exactly characterizing the isoperimetric profile is in general extremely difficult, and remains widely open except for certain specific cases. For example, when $\nu$ is the standard Gaussian measure, the isoperimetric profile $I_{n}(a)=I_{\mathrm{G}}(a):=\phi(\Phi^{-1}(a))$ for all $n\geq 1$ with half-spaces as isoperimetric minimizers [61, 18]. Here, $\phi$ and $\Phi$ are respectively the pdf and cdf of the standard Gaussian. It is widely known that $I_{\mathrm{G}}(a)\sim a\sqrt{2\ln\frac{1}{a}}$ as $a\to 0$. That is, the asymptotic isoperimetric ratio for the Gaussian measure is
(1.4) $\lim_{a\to 0}\frac{I_{\mathrm{G}}(a)}{a\sqrt{\ln\frac{1}{a}}}=\sqrt{2}.$
Another example is log-concave measures on the real line (i.e., $n=1$), for which the isoperimetric profile is attained by intervals or half-lines [11]. Furthermore, in the Euclidean sphere of a given dimension equipped with constant density, the isoperimetric profile is known as well, which is attained by spherical caps.
As noted at the outset of this paper, our primary aim is to study the asymptotic behavior of the isoperimetric profile, in particular its large deviations asymptotics. We focus on this behavior in the regime where the probability measure of the set decays exponentially with increasing dimension, and to this end we identify the following limits: for $\alpha>0$,
(1.5) $\underline{\Lambda}(\alpha)$ $:=\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}},$
and $\overline{\Lambda}(\alpha)$ which is defined similarly but with $\liminf_{n\to\infty}$ replaced by $\limsup_{n\to\infty}$. We denote them by $\Lambda(\alpha)$ if they are equal.
1.1. Main Results
We next provide a detailed description of our main results.
1.1.1. Large Deviations Principle
Assumption 1 (Convexity).
A natural convexity condition on a manifold with density is the following curvature-dimension condition $\mathrm{CD}(0,\infty)$:
$\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq 0\;\;\text{ as 2-tensor fields},$
where $\mathrm{Ric}_{g}$ is the Ricci curvature and $\mathrm{Hess}_{g}V$ is the Hessian of $V$.
Since $I_{n}(a)/a$ is non-increasing in $a$ under Assumption 1 [45, Proposition 3.1], $\Lambda$ is non-decreasing.
Proposition 1.
Under Assumption 1, $\Lambda$ is non-decreasing.
Let $P_{2}^{\mathrm{ac}}(\mathcal{M})$ be the set of probability measures which are absolutely continuous with respect to $\nu$ and have finite second-order moments. We also need the concept of nonlinear log-Sobolev inequality and a generalized version of half-spaces.
Definition 1.
Given a probability measure $\nu$, for a convex function $\theta:[0,\infty]\to[0,\infty]$, the $\theta$-nonlinear log-Sobolev inequality is defined to be the following inequality
(1.6) $\theta(D(\mu\|\nu))\leq I(\mu\|\nu),\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}),$
where for $\mu=\rho\nu$,
$D(\mu\|\nu):=\int\rho\ln\rho d\nu\;\textrm{ and }\;I(\mu\|\nu):=\int\frac{|\nabla\rho|^{2}}{\rho}d\nu$
are respectively the (relative) entropy and Fisher (relative) information.
Remark 1.
The inequality in (1.6) holds for all $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ if and only if it holds for all $\mu\in P^{\mathrm{ac}}(\mathcal{M})$ (the set of probability measures which are absolutely continuous with respect to $\nu$) [54].
The nonlinear log-Sobolev inequality, also known as the entropy-energy inequality or the entropy isoperimetric inequality, can be seen as a strengthening of the standard (linear) log-Sobolev inequality since it provides more information about the tradeoff between entropy and Fisher information. This concept was investigated in several important works in the literature, e.g., Costa and Cover [20], F. Y. Wang [66], Bakry, Gentil, and Ledoux [7], as well as Polyanskiy and Samorodnitsky [55].
Any convex $\theta$ that admits a $\theta$-nonlinear log-Sobolev inequality must satisfy that $\theta(0)=0$ and $\theta$ is non-decreasing. In particular, when $\theta(t)=Kt$ for some constant $K>0$, then the $\theta$-nonlinear log-Sobolev inequality reduces to the standard log-Sobolev inequality. Define
$\Theta(\alpha):=\inf_{\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu)\geq\alpha}I(\mu\|\nu).$
Then, given a probability measure $\nu$, the optimal convex function $\theta$ in the $\theta$-nonlinear log-Sobolev inequality in (1.6) is the lower convex envelope of $\Theta$, given by
$\breve{\Theta}(\alpha)$ $:=\inf_{\begin{subarray}{c}\lambda\in[0,1],\alpha_{1},\alpha_{2}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{2}\geq\alpha\end{subarray}}\lambda\Theta(\alpha_{1})+(1-\lambda)\Theta(\alpha_{2})$
(1.7) $=\inf_{\begin{subarray}{c}\lambda\in[0,1],\mu_{1},\mu_{2}\in P_{2}^{\mathrm{ac}}(\mathcal{M}):\\ \lambda D(\mu_{1}\|\nu)+(1-\lambda)D(\mu_{2}\|\nu)\geq\alpha\end{subarray}}\lambda I(\mu_{1}\|\nu)+(1-\lambda)I(\mu_{2}\|\nu).$
It is obvious that $\breve{\Theta}(0)=0$ and $\breve{\Theta}$ is non-decreasing and convex.
The second concept we need is a generalized version of half-spaces which incorporates many important concepts as special cases, e.g., the concept of typical sets from information theory.
Definition 2.
Given $n\in\mathbb{N}$ and $\mathbf{f}=(f_{1},f_{2},...,f_{n})$ with each $f_{i}:\mathcal{M}\to\mathbb{R}$, the $\mathbf{f}$-coordinate half-space is defined as
(1.8) $H_{t}(\mathbf{f}):=\Big\{\mathbf{x}\in\mathcal{M}^{n}:\sum_{i=1}^{n}f_{i}(x_{i})\geq t\Big\}$
for a real number $t$. In particular, when $f_{i}=\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}$ for some probability measures $\mu_{i},i\in[n]$ and $t=\sum_{i=1}^{n}D(\mu_{i}\|\nu)-n\epsilon$ for some $\epsilon>0$, the $\mathbf{f}$-coordinate half-space reduces to the conditional (relative entropy) typical set of $(\mu_{i})_{i\in[n]}$ w.r.t. $\nu$ which is given by
$A_{n,\epsilon}((\mu_{i})_{i\in[n]}\|\nu):=\Big\{\mathbf{x}\in\mathcal{M}^{n}:\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})\geq\sum_{i=1}^{n}D(\mu_{i}\|\nu)-n\epsilon\Big\}.$
Remark 2.
The concept of $\mathbf{f}$-coordinate half-space reduces to the Orlicz ball if $f_{i}$ are Young functions (and t>0-t>0). However, we do not pose any restriction on functions $f_{i}$ here (except for measurability), and thus our concept is much more general and incorporates standard half-spaces as special cases.
Remark 3.
Here, $(\mu_{i})_{i\in[n]}$ can be seen as a conditional probability measure $\mu_{X|W}$ with $\mu_{X|W=i}=\mu_{i}$, $\frac{1}{n}\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})$ is known as the conditional information density of $\mu_{X|W}$ w.r.t. $\nu$ given $\mu_{W}\sim\mathrm{Unif}[n]$, and its expectation is $D(\mu_{X|W}\|\nu|\mu_{W}):=\frac{1}{n}\sum_{i=1}^{n}D(\mu_{i}\|\nu)$ which is known as the conditional relative entropy. That is why it is called the conditional typical set. In addition, the standard definition of the conditional relative entropy typical set is $\big\{\mathbf{x}\in\mathcal{M}^{n}:\big|\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})-\sum_{i=1}^{n}D(\mu_{i}\|\nu)\big|\leq n\epsilon\big\}$ [21, (11.206)]. Our definition here is slightly different from the standard one, and has benefit that it incorporates the isoperimetric minimizers for Gaussian and Lebesgue measures as special cases, which is explained below.
The concept of $\mathbf{f}$-coordinate half-space generalizes the standard half-spaces $\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}w_{i}x_{i}\leq t\}$ in Euclidean spaces, and the former reduces to the latter when $f_{i}(x_{i})=-w_{i}x_{i}$ and $t$ is rechosen, or essentially equivalently, set $\nu=\mathcal{N}(0,1),\mu_{i}=\mathcal{N}(w_{i},1)$ in the conditional typical set. It also generalizes the Euclidean balls $\left\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq t\right\}$ with $t>0$, and the former reduces to the latter when $f_{i}(x_{i})=-x_{i}^{2}$ and $t$ is rechosen, or essentially equivalently, set $\nu=\mathrm{Lebesgue},\mu=\mathcal{N}(0,c)$ with $c>0$ in the conditional typical set (although the Lebesgue measure is an infinite measure and not considered in this paper). Note that standard half-spaces and Euclidean balls are respectively isoperimetric minimizers for Gaussian measures and Lebesgue measures.
We now establish the large deviations principle for the isoperimetric problem, a result of particular interest as it reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. We also need the following assumption.
Assumption 2 (Full Exponential Integrability of Squared Distance).
For some $x_{0}\in\mathcal{M}$ (equivalently for any $x_{0}\in\mathcal{M}$),
$\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\forall\lambda\in(0,\infty).$
In other words, the probability measure $\nu$} admits a finite moment generating function for squared distance.
This assumption is satisfied if $\mathcal{M}$ is a bounded Riemann manifold or if $\mathcal{M}$ is a Euclidean space with $\nu$ having density $\propto e^{-C|x|^{\beta}}$ where $C>0$ and $\beta>2$.
Theorem 1 (Large Deviations Principle for Isoperimetry).
Under Assumptions 1 and 2, it holds that for any $\alpha>0$,
(1.9) $\Lambda(\alpha)=\sqrt{\breve{\Theta}(\alpha)}$}.
Moreover, a sequence of conditional typical sets $A_{n,\epsilon}((\mu_{1}^{*},...,\mu_{1}^{*},\mu_{2}^{*},...,\mu_{2}^{*})\|\nu)$ with $\mu_{1}^{*},\mu_{2}^{*}$ respectively appearing $\lfloor n\lambda^{*}\rfloor$ and $n-\lfloor n\lambda^{*}\rfloor$ times asymptotically attains the limit $\Lambda(\alpha)$ as $n\to\infty$ first and $\epsilon\to 0$ then, where $(\lambda^{*},\mu_{1}^{*},\mu_{2}^{*})$ is the optimal solution in (1.7). In addition, the inequality $\overline{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$} holds even without Assumption 2.
Remark 4.
Although this theorem does not include the Gaussian measure case and the Lebesgue measure case, for these two cases, the conditional typical sets with $\mu$ set to the extremizers (see [55]) respectively reduce to standard half-spaces and Euclidean balls, consistent with the already known fact that these two kinds of sets are isoperimetric extremizers for these two cases. Moreover, by this already known fact, Equation (1.9) still holds for these two cases, suggesting that Theorem 1 might hold for a much larger class of measures.
This theorem can be alternatively formulated as follows: For any convex $\theta:[0,\infty)\to[0,\infty)$,
${\displaystyle\Lambda(\alpha)\geq\sqrt{\theta(\alpha)}},\;\forall\alpha>0\quad\Longleftrightarrow\quad I(\mu\|\nu)\geq\theta(D(\mu\|\nu)),\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}).$
Equation (1.9) is also equivalent to that for any $\alpha>0$,
(1.10) $\lim_{n\to\infty}\frac{I_{n}(a_{n})}{I_{\mathrm{G}}(a_{n})}=\lim_{n\to\infty}\frac{I_{n}(a_{n})}{a_{n}\sqrt{2\ln\frac{1}{a_{n}}}}=\sqrt{\frac{\breve{\Theta}(\alpha)}{2\alpha}}$},
where $a_{n}=e^{-n\alpha}$ and $I_{\mathrm{G}}$ is the isoperimetric profile for the standard Gaussian measure. That is, as the volume vanishes exponentially, the asymptotic behavior of the isoperimetric profile for any product probability measure resembles the one for the standard Gaussian measure, but with different factors that are determined by the exponential convergence rate of the volume. One prototypical case occurs when $\Theta(\alpha)$ grows linearly in $\alpha$—for instance, $\Theta(\alpha)=2\alpha$ for Gaussian measures—implying that $I_{n}(a)$ behaves as $C\cdot a\sqrt{\ln\frac{1}{a}}$, with standard half-spaces likely serving as approximate extremizers for small volumes. A second prototypical case is when $\Theta(\alpha)$ grows exponentially in $\alpha$, as illustrated in the subsequent example, which implies that $I_{n}(a)$ behaves as $C\cdot\sqrt{n}\,a^{1-1/n}$, with geodesic balls acting as approximate extremizers for small volumes. Theorem 1 provides a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers of small volumes behave fundamentally differently across spaces with distinct geometric structures.
Example 1 (Exponential Growth of $\Theta$).
Let $D(t)=D(\mu_{t}\|\nu)$, $I(t)=I(\mu_{t}\|\nu)$, and $N(t)=\exp(-2D(\mu_{t}\|\nu))$ respectively be the relative entropy, Fisher information, and entropy power along the (generalized) heat flow $\partial_{t}\rho=L\rho$ with $L=\Delta-\nabla V\cdot\nabla$ and $\mu_{t}=\rho_{t}\nu$. It is easy to see that $\Theta(\alpha)$ grows exponentially as $\alpha\to\infty$, as long as $N(t)$ is concave in the time $t$ and $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$. This is because, in this case, $D(t)$ continuously decays to zero as $t\to\infty$ and $N^{\prime}(t)=2N(t)I(t)$ is nonincreasing, which implies that for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, we always have $N(0)I(0)\geq N(t_{0})I(t_{0})\geq\exp(-2\alpha_{0})\Theta(\alpha_{0})$ with $t_{0}$ being the solution to $D(t_{0})=\alpha_{0}$. Therefore, $I(0)\geq\Theta(\alpha_{0})\exp(2(D(0)-\alpha_{0}))$ for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, i.e., exponential growth of $\Theta$: $\Theta(\alpha)\geq\Theta(\alpha_{0})\exp(2(\alpha-\alpha_{0}))$ for all $\alpha\geq\alpha_{0}$. The concavity of the entropy power holds under the $\mathrm{CD}(0,m)$-condition with $m\in[n,\infty)$, as shown by S. Li and X. Li [42, Theorem 2.2]. The condition that $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$ is guaranteed as long as a standard log-Sobolev inequality (or a Poincaré inequality) is admitted. These two conditions are satisfied by the (normalized) unweighted measure on a bounded manifold with nonnegative Ricci curvature, which hence admits exponential growth of $\Theta$ and also exponential growth of $\Lambda$ (since Assumptions 1 and 2 are satisfied in this setting and Theorem 1 is valid). See other sufficient conditions to ensure exponential growth of $\Theta$, e.g. , in [7, Theorem 6.8.1][42, Theorem 2.5].
We next consider a simple but nontrivial case—the uniform measure on hypercube.
Example 2 (Hypercube).
Consider the hypercube $(0,1)^{n}$ with Euclidean metric and uniform measure. The standard log-Sobolev inequality and the Poincaré inequality with constants $K_{LS}=K_{P}=\pi^{2}$ ($K_{LS},K_{P}$ defined later in (1.12) and (1.18)) respectively hold in this setting [27], which implies $\lim_{\alpha\to 0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}=\pi^{2}$. Moreover, the $\mathrm{CD}(0,m)$-condition holds obviously. Thus, as mentioned in Example 1, we have exponential growth of $\Theta$, i.e., $\Theta(\alpha)\geq 2\pi^{2}\alpha_{0}\cdot e^{2(\alpha-\alpha_{0})}\geq\frac{\pi^{2}}{e}\cdot e^{2\alpha}$ (with $\alpha_{0}$ chosen $1/2$), which further implies exponential growth of $\Lambda$, i.e., $\Lambda(\alpha)\geq\sqrt{\frac{\pi^{2}}{e}}\cdot e^{\alpha}$. This lower bound is sharp in the exponent by verifying balls centered at the vertices of the hypercube, since for these balls with volume $e^{-n\alpha}$, their surface areas are $(1+o(1))\sqrt{\frac{\pi en}{2}}\cdot e^{\alpha(1-n)}$ as $n\to\infty$, giving $\Lambda(\alpha)\leq\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$. In fact, it was already known [58, 29] that given each $n$, for all small volumes, balls centered at the vertices are isoperimetric minimizers. However, it is unclear whether “small volume” here is large enough to include “exponentially small volume”. We conjecture the following precise relation holds: $\Lambda(\alpha)\sim\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$, or equivalently, $\breve{\Theta}(\alpha)\sim\frac{\pi e}{2}\cdot e^{2\alpha}$ as $\alpha\to\infty$, suggesting that half-normal distributions truncated on $(0,1)$ with sufficiently small variances approximately saturate nonlinear log-Sobolev inequalities for large $\alpha$. The factor $\frac{\pi^{2}}{e}$ obtained above is about $0.85\cdot\frac{\pi e}{2}$.
Theorem 1 reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. It is termed the large deviations principle for isoperimetry for two reasons. First, both this theorem and large deviations theory focus on sets of exponentially small measures. Second, isoperimetric minimizers have been proven to be expressed in terms of $\mathbf{f}$-coordinate half-spaces (or conditional typical sets) which concern the remote tails of a sum. These remote tails themselves are the core object of investigation in large deviations theory.
Showing the equivalence among inequalities from various fields is an interesting topic, since it will unveil the intrinsic relationships among diverse branches of mathematics. However, most of equivalences in the literature on this topic are known only in a weak sense. In particular, it was proved by Ledoux [36, 37] on Gaussian spaces or compact Riemannian manifolds and by Bobkov [14] on Euclidean spaces with any log-concave probability measures that for a fixed $n$,
$I_{n}(a)\geq a\sqrt{2K\ln\frac{1}{a}}$
holds for some $K>0$, if and only if the standard log-Sobolev inequality holds with some constant $K^{\prime}>0$. Theorem 1 provides an equivalence in a stronger sense for product spaces, which showing a precise equivalence between the isoperimetric inequality and (nonlinear) log-Sobolev inequality. Our equivalence will be applied in the following to compare the optimal constants in various inequalities.
1.1.2. Companion of Various Constants
While Theorem 1 is primarily tailored to subsets with probability measures decaying exponentially in $n$, it admits application to bounding the isoperimetric profile for subsets with probability measures decaying arbitrarily slowly. In other words, Theorem 1 can be used to quantitatively compare the log-Sobolev constant, the isoperimetric constant, and other constants for various inequalities. To this end, we define the following constants.
  • $K_{CD}$ denotes the optimal constant $K$ in the Ricci curvature-dimension condition
    (1.11) $\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq Kg.$
  • $K_{LS}$ denotes the optimal constant $K$ in the log-Sobolev inequality
    (1.12) $I(\mu\|\nu)\geq 2K\cdot D(\mu\|\nu),\;\forall\mu,$
    or equivalently,
    $K_{LS}=\frac{\breve{\Theta}^{\prime}(0)}{2}=\inf_{\alpha>0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\inf_{\alpha>0}\frac{\Theta(\alpha)}{2\alpha},$
    where $\breve{\Theta}^{\prime}(0)$ is the right derivative of $\breve{\Theta}$ at 0.
  • $K_{LS}^{+}$ denotes
    $K_{LS}^{+}:=\sup_{\alpha>0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to\infty}\frac{\breve{\Theta}(\alpha)}{2\alpha}.$
  • $K_{HC}$ denotes the optimal constant $K$ in the hypercontractivity inequality for the semigroup $T_{t}=e^{-tL}$ with $L=\Delta-\nabla V\cdot\nabla$,
    $\|T_{t}f\|_{q}\leq\|f\|_{p},\;\forall f,$
    for all $t\geq\frac{1}{2K}\ln\frac{q-1}{p-1}$, where the norms are taken w.r.t. $\nu$.
  • $K_{T}$ denotes the optimal constant $K$ in the transport inequality
    $\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}),\quad\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},$
    where
    $\mathsf{W}_{2}(\mu,\nu):=\left(\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{M}\times\mathcal{M}}d^{2}(x,y){\text{\rm d}}\pi(x,y)\right)^{1/2}$
    is the Wasserstein distance of order $2$ and $\Pi(\mu,\nu)$ is the set of couplings of $\mu,\nu$, i.e., the set of joint distributions with marginals equal to $\mu,\nu$.
  • $K_{T}^{+}$ denotes the optimal constant $K$ in the transport inequality that there is some constant $D_{0}\geq 0$ such that for all
    $\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},\quad\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})\textrm{ s.t. }D(\mu\|\nu)\geq D_{0}.$
  • $K_{C}$ denotes the optimal constant $K$ in the dimension-free Gaussian concentration inequality that for all $n\geq 1$ and $r\geq 0$, $\nu_{n}((A^{r})^{c})\leq e^{-\frac{Kr^{2}}{2}}$ for all $A$ such that $\nu_{n}(A)\geq 1/2$.
  • $K_{C}^{+}$ denotes the optimal constant $K$ in the asymptotic Gaussian concentration inequality that there is some constant $r_{0}\geq 0$ such that for all $n\geq 1$ and $r\geq\sqrt{n}r_{0}$, $\nu_{n}((A^{r})^{c})\leq e^{-\frac{Kr^{2}}{2}}$ for all $A$ such that $\nu_{n}(A)\geq 1/2$.
  • $K_{IS}$ denotes the optimal constant $K$ in the large deviations asymptotics of the isoperimetric inequalities
    $\liminf_{n\to\infty}\frac{I_{n}(a_{n})}{a_{n}\sqrt{2\ln\frac{1}{a_{n}}}}\geq\sqrt{K},\;\forall\alpha>0,$
    where $a_{n}=e^{-n\alpha}$.
  • $K_{IS}^{+}$ denotes the optimal constant $K$ in the following asymptotic isoperimetric inequality
    $\liminf_{a\to 0}\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq\sqrt{K},\;\forall n\geq 1.$
  • $K_{IS}^{-}$ denotes the constant in the dimension-free (also infinite-dimensional) isoperimetric inequality
    (1.13) $\liminf_{a\to 0}\frac{I_{\inf}(a)}{a\sqrt{2\ln\frac{1}{a}}}=\sqrt{K_{IS}^{-}},$
    where
    $I_{\inf}(a):=\lim_{n\to\infty}I_{n}(a)=\inf_{n\geq 1}I_{n}(a).$
From the definitions, the constants $K_{IS}$, $K_{IS}^{+}$, and $K_{IS}^{-}$ are the asymptotic isoperimetric ratios $\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}$ respectively for exponentially, superexponentially, and subexponentially small sequences of volumes. We now compare all optimal constants defined above.
Theorem 2.
Under Assumption 1 (i.e., $K_{CD}\geq 0$), the following hold.
  1. (1)
    It holds that
    (1.14) $K_{CD}\overset{\mathrm{(a)}}{\leq}\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\overset{\mathrm{(b)}}{\leq}$ $K_{IS}^{-}\overset{\mathrm{(c)}}{\leq}K_{IS}\overset{\mathrm{(d)}}{\leq}K_{LS}\overset{\mathrm{(e)}}{=}K_{HC}$
    $\overset{\mathrm{(f)}}{\leq}K_{T}\overset{\mathrm{(g)}}{=}K_{C}\overset{\mathrm{(h)}}{\leq}K_{C}^{+}\overset{\mathrm{(i)}}{=}K_{T}^{+},$
    and
    $K_{LS}\overset{\mathrm{(j)}}{\leq}K_{LS}^{+}\overset{\mathrm{(k)}}{\leq}K_{T}^{+},$
    where $x\wedge y=\min\{x,y\}$. Under the additional Assumption 2, the equality in (d) holds and moreover,
    $K_{LS}^{+}\overset{\mathrm{(l)}}{\leq}K_{IS}^{+}.$
    In addition, if $K_{CD}$ is equal to any one of $K_{IS}^{-},K_{IS},K_{LS},K_{HC},K_{T},K_{C}$, then all the inequalities in (a)–(g) of (1.14) become equalities.
  2. (2)
    For any log-concave probability measure $\nu$ on $\mathbb{R}$ with density $e^{-V}$ such that $\lim_{x\to\infty}V^{\prime\prime}(x)$ and $\lim_{x\to-\infty}V^{\prime\prime}(x)$ exist,
    $K_{IS,1}^{+}:=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq K_{LS}^{+}.$
For Statement 1, Inequality (a) is trivial, by noting that $0\leq K_{CD}/K_{C}\leq 1$. Inequality (c) is due to the fact that $n\mapsto I_{n}(a)$ is nonincreasing. Equality (d) is implied by Theorem 1. Equality (e) is a classic result of Gross [33]. The inequality $K_{LS}\leq K_{T}$ (or equivalently, Inequality (f)) was proven by Otto and Villani [54]. Equality (g) is also well known; see e.g. [65, Remark 22.23]. Inequalities (h) and (j) follow by definition. Equality (i) follows similarly to (g). Inequality (k) follows similarly to $K_{LS}\leq K_{T}$. Inequality (l) follows by Theorem 1, with a detailed proof given in Appendix A. So, it only remains to prove Inequality (b). This is done in Appendix A. In addition, a weaker version of (b), $\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\leq K_{LS}$, was already known; see Theorem 22.21 of [65]. However, we do not know if this is equivalent to the inequality (b) since we do not know if $K_{IS}^{-}=K_{IS}$ holds. Statement 2 is proven in Appendix A.
Example 3 (Hypercube).
We continue considering the hypercube described in Example 2. Statements in Example 2 imply that for this case,
$K_{CD}=0,\;K_{IS}=K_{LS}=K_{HC}=K_{T}=K_{C}=\pi^{2},\;K_{LS}^{+}=K_{C}^{+}=K_{T}^{+}=K_{IS}^{+}=\infty,$
where the density $1+\epsilon\cos\pi x$ asymptotically attains the constants $K_{LS}$ and $K_{T}$ as $\epsilon\to 0$. Moreover, it holds that $\sqrt{2\pi}\leq K_{IS}^{-}\leq\pi^{2}$; the lower bound will be explaind in Example 4. However, the exact value of $K_{IS}^{-}$ is still unknown.
Without Assumption 2, we have the following weaker result, as a consequence of Theorem 2.
Theorem 3.
Under Assumption 1, it holds that for any $\alpha>0$,
(1.15) $\sqrt{\frac{K_{LS}}{4}\alpha}\leq\sqrt{\frac{K_{T}}{4}\alpha}\leq\underline{\Lambda}(\alpha)\leq\overline{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$}.
1.1.3. Central Limit Regime
Let $I_{\infty}$ be the isoperimetric profile defined for the infinite dimensional product space $\mathcal{M}^{\infty}$ equipped with the product topology and the product measure $\nu_{\infty}:=\nu^{\otimes\infty}$, but the enlargement of a set is defined under $d_{\infty}(x,y)=\sqrt{\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}}$ (although it might diverse to infinity for two arbitrary points). Then, it is obvious that $I_{\inf}(a)\geq I_{\infty}(a)$. In the following, we show that this is actually an equality. That is, the isoperimetric profile $I_{\infty}$ on the infinite dimensional space is the pointwise limit of the isoperimetric profiles on finite dimensional spaces. The proof is provided in Appendix B.
Theorem 4.
Assume that $I_{\inf}$ is continuous (e.g., Assumption 1 holds). Then, $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in[0,1]$.
A well-known lower bound on $I_{\inf}$ is that
(1.16) $I_{\inf}(a)\geq K_{0}\cdot I_{\mathrm{G}}(a),\;\forall a\in[0,1],$
where $I_{\mathrm{G}}$ is the Gaussian isoperimetric profile and $K_{0}:=\inf_{a\in(0,1)}\frac{I_{\nu}(a)}{I_{\mathrm{G}}(a)}$ [13, 8]. An upper bound can be obtained by considering the $\mathbf{f}$-coordinate half-space given in (1.8), as shown in the following theorem. The proof is based on the coarea formula and the central limit theorem, which is given in Appendix C.
Theorem 5 (Upper Bound for Isoperimetry in Central Limit Regime).
Assume that $I_{\inf}$ is continuous (e.g., Assumption 1 holds). Then, for any $a\in[0,1]$,
(1.17) $I_{\inf}(a)\leq\sqrt{K_{P}}\cdot I_{\mathrm{G}}(a),$
where $K_{P}$ is the spectral gap (the reciprocal of the Poincaré constant) for $\nu$, i.e., the optimal constant in
(1.18) $K_{P}\cdot\mathrm{Var}_{\nu}(f)\leq\mathbb{E}_{\nu}\left[\|\nabla f\|^{2}\right]$
for all $f$ in the Sobolev space $W^{1,2}(\mathcal{M})$.
As a related result, the Poincaré constant and the Cheeger constant are known to be closely related; see e.g., [35]. The Cheeger constant characterizes the isoperimetric behavior exclusively for large subsets, whereas the inequality in (1.17) captures isoperimetry for all subsets without restriction.
The upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ in (1.17) can be used to bound $K_{IS}^{-}$, namely, obtaining that $K_{IS}^{-}\leq K_{P}$. However, this bound is worse than the upper bound $K_{LS}$, since in general $K_{LS}\leq K_{P}$. Note that by perturbation analysis, one can see that
$K_{P}=\liminf_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}.$
Thus, $K_{LS}=K_{P}$ if $\Theta(\alpha)\geq 2K_{P}\alpha$ (e.g., if $\Theta$ is convex), and otherwise, $K_{LS}<K_{P}$ (e.g., for exponential measures [65, Example 21.19]). For the latter case, the upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ is not tight for sufficiently small $a$. Combining the best known lower bound in (1.16), the upper bound in (1.17), and the trivial upper bound $I_{1}$ yields
$K_{0}I_{\mathrm{G}}(a)\leq I_{\inf}(a)\leq I_{1}(a)\wedge\sqrt{K_{P}}I_{\mathrm{G}}(a).$
Example 4 (Hypercube).
We continue considering the hypercube described in Example 2. Applying the inequalities above to the hypercube $(0,1)^{n}$ yields $\sqrt{2\pi}I_{\mathrm{G}}(a)\leq I_{\inf}(a)\leq 1\wedge\pi I_{\mathrm{G}}(a)$. As shown in [29], the lower bound is not tight.
1.2. Open Problems
As mentioned above, one major open question on the topic considered in this paper is the determination of $K_{IS}^{-}$.
Problem 1.
Under Assumption 1, is it true that $K_{IS}^{-}=K_{LS}$?
A more challenging problem is to exactly determine the function $I_{\inf}$ (or equivalently, $I_{\infty}$). This problem is a commonly recognized open one.
Problem 2.
Under Assumption 1, what is the explicit expression of $I_{\inf}$?
Another interesting topic is to remove Assumption 2 from Theorem 1. To this end, we propose the following conjecture.
Conjecture 1 (Pythagorean-Like Inequality).
Let $\nu_{1}=e^{-V_{1}}\mathrm{vol}_{1},\nu_{2}=e^{-V_{2}}\mathrm{vol}_{2}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy Assumption 1. Then,
(1.19) $\left(\frac{I_{\nu_{1}\otimes\nu_{2}}(a)}{a}\right)^{2}\geq\inf_{a_{1}a_{2}=a}\left(\frac{I_{\nu_{1}}(a_{1})}{a_{1}}\right)^{2}+\left(\frac{I_{\nu_{2}}(a_{2})}{a_{2}}\right)^{2},\;\forall a\in(0,1).$
Remark 5.
A weaker version of this conjecture was proven by Grigor’yan [32] and Morgan [50].
This conjecture can be equivalently formulated as the plane version, as explained in the following. Any probability measure $\nu$ satisfying Assumption 1 admits a concave isoperimetric profile $I_{\nu}$. For every such probability measure $\nu$, one can find a unique symmetric (w.r.t. $x=0$) log-concave probability measure $\tilde{\nu}$ on $\mathbb{R}$ such that $I_{\tilde{\nu}}=I_{\nu}$ [14]. Explicitly, the cdf $F$ of $\tilde{\nu}$ is exactly the inverse of the function
$G(y)=\int_{1/2}^{y}\frac{dz}{I_{\nu}(z)},$
which will induces $I_{\tilde{\nu}}=f\circ F^{-1}=I_{\nu}$, with $f=F^{\prime}$ denoting the pdf of $\nu$. Let $\nu_{1}=e^{-V_{1}}\mathrm{vol},\nu_{2}=e^{-V_{2}}\mathrm{vol}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy Assumption 1. Let $\tilde{\nu}_{1},\tilde{\nu}_{2}$ be the symmetric log-concave probability measures on $\mathbb{R}$ induced by $\nu_{1},\nu_{2}$ respectively such that $I_{\tilde{\nu}_{i}}=I_{\nu_{i}},i=1,2$. Denote $F_{i},f_{i}$ as the cdf and pdf of $\tilde{\nu}_{i}$, $i=1,2$. Define
(1.20) $J_{1}(a)$ $=\inf_{\sigma}\int_{\mathcal{M}_{1}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+|\nabla\sigma(x)|^{2}}{\text{\rm d}}\nu_{1}(x),$
$J_{2}(a)$ $=\inf_{\varphi}\int_{0}^{1}\sqrt{I_{\nu_{2}}^{2}(\varphi(t))+I_{\nu_{1}}^{2}(t)\varphi^{\prime}(t)^{2}}{\text{\rm d}}t,$
$J_{3}(a)$ $=\inf_{g}\int_{\mathbb{R}}\sqrt{1+(g^{\prime}(x))^{2}}f_{2}(g(x))f_{1}(x){\text{\rm d}}x,$
where the first infimum is taken over all smooth functions $\sigma:\mathcal{M}_{1}\to[0,1]$ such that $\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)=a$, the second infimum is taken over all smooth functions $\varphi:[0,1]\to[0,1]$ such that $\int_{0}^{1}\varphi(t){\text{\rm d}}t=a$, and the third infimum is taken over all smooth functions $g:\mathbb{R}\to\mathbb{R}$ such that $\int_{\mathbb{R}}F_{2}(g(x))f_{1}(x){\text{\rm d}}x=a$. The following proposition shows that the isoperimetric profile of $\nu_{1}\otimes\nu_{2}$ will decrease if we replace one or both of $\nu_{1},\nu_{2}$ to the corresponding $\tilde{\nu}_{1},\tilde{\nu}_{2}$. This proposition is an extension of the Gaussian version due to Bobkov [13] and the weighted Riemannian slab version due to [46] to Riemannian manifolds, and a simple proof is provided in Appendix D for completeness.
Proposition 2.
It holds that for all $a\in[0,1]$,
$I_{\nu_{1}\otimes\nu_{2}}(a)$ $\geq I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)=J_{1}(a)\geq I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)=J_{2}(a)=J_{3}(a).$
Remark 6.
By symmetry, $I_{\nu_{1}\otimes\nu_{2}}(a)\geq I_{\tilde{\nu}_{1}\otimes\nu_{2}}(a)\geq I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)$ also holds. In addition, this proposition can be used to derive dimension-dependent isoperimetric inequalities. Suppose that $\Upsilon_{i}:[0,1]\to[0,\infty),i\in\{1,2,...,n\}$ are functions such that $I_{\nu_{1}}\geq\Upsilon_{1}$ and for all $i\in\{2,3,...,n\}$,
$\int_{\mathcal{M}_{i}}\sqrt{\Upsilon_{i-1}^{2}(\sigma_{i}(x))+|\nabla\sigma_{i}(x)|^{2}}{\text{\rm d}}\nu_{i}(x)\geq\Upsilon_{i}(\int_{\mathcal{M}_{i}}\sigma_{i}(x){\text{\rm d}}\nu_{i}(x))$
holds for all smooth functions $\sigma_{i}:\mathcal{M}_{i}\to[0,1]$. Then, by repeatly applying Proposition 2, $I_{\nu_{1}\otimes\nu_{2}\otimes...\otimes\nu_{i}}\geq\Upsilon_{i},i\in\{1,2,...,n\}$.
This proposition implies that it suffices to restrict the inequality in Conjecture 1 to the Euclidean plane, yielding the following equivalent formulation of Conjecture 1.
Conjecture 2.
For any symmetric log-concave probability measures $\nu_{1},\nu_{2}$ on $\mathbb{R}$ (in which case, $I_{\nu_{i}}=f_{i}\circ F_{i}^{-1}$ with $f_{i}$ and $F_{i}$ denoting the pdf and cdf of $\nu_{i}$),
$\left(\frac{I_{\nu_{1}\otimes\nu_{2}}(a)}{a}\right)^{2}\geq\inf_{a_{1}a_{2}=a}\left(\frac{I_{\nu_{1}}(a_{1})}{a_{1}}\right)^{2}+\left(\frac{I_{\nu_{2}}(a_{2})}{a_{2}}\right)^{2},\;\forall a\in(0,1).$
We next show that under Conjecture 2, Theorem 1 still holds for one-dimensional Euclidean space without Assumption 2. The proof of the following proposition is given in Appendix E.
Proposition 3.
Assume that Conjecture 2 holds. Let $\nu$ be a log-concave probability measure on $\mathbb{R}$ with density $e^{-V}$ such that $\lim_{x\to\infty}V^{\prime\prime}(x)$ and $\lim_{x\to-\infty}V^{\prime\prime}(x)$ exist, Theorem 1 still holds.
1.3. Related Works
  • Isoperimetry
In the one-dimensional setting, Bobkov [11] established that half-lines are extremal sets in the isoperimetric inequality for log-concave probability measures. Extending to higher-dimensional structures, Bobkov and Houdré [15] further explored the isoperimetric constant and log-Sobolev inequalities in product spaces, laying a foundation for multi-dimensional isoperimetric analysis. The optimality of standard half-spaces— a core topic in isoperimetry— has been systematically studied by multiple scholars: Sudakov and Tsirel’son [61] initiated early investigations, Bobkov [11] deepened this topic by addressing the isoperimetric problem for log-concave product measures on $\mathbb{R}^{n}$ equipped with the uniform distance, and explicitly provided necessary and sufficient conditions for standard half-spaces to be extremal, and Barthe [9] investigated the extremal properties of central half-spaces (with measure $1/2$) for product probability measures on Riemannian manifolds equipped with the geodesic distance. It should be noting that the isoperimetric problem under the geodesic distance and the one under the uniform distance are significantly different. When equipped with the uniform distance, the asymptotics of the isoperimetric profile has been characterized by Bobkov in [12]. However, for the geodesic distance as in our setup, the asymptotics is still widely unknown except for special cases, e.g., the Gaussian setting. Roberto [59] offers a critical survey for a comprehensive overview of isoperimetry in product spaces. Furthermore, the isoperimery for small sets was already investigated in the literature; see, e.g., [41] for the continuous setting and [56, 53] for the discrete setting. In the discrete setting, the isoperimery for small sets is known as the small set expansion and has important applications in complexity theory [56]. It should be noted that characterizing the isoperimetric profile remains widely open, and previously so does its asymptotics.
  • Concentration of Measure
The concentration of measure in a probability metric space describes a key phenomenon: a slight enlargement of any measurable set with non-negligible probability will always result in a set with large probability. In functional analysis terms, this is equivalent to the fact that the value of any Lipschitz function is concentrated around its median. The modern study of this phenomenon began in the early 1970s, when V. Milman pioneered its application to the asymptotic geometry of Banach spaces. Subsequent decades saw in-depth exploration by scholars including Gromov, Maurey, Pisier, Schechtman, Talagrand, and Ledoux, solidifying its role in analysis and probability. A critical advancement came from Talagrand [62], who studied concentration of measure in product spaces equipped with product probability measures and derived a suite of sharp concentration inequalities. In information theory, this phenomenon is known as the blowing-up lemma [1, 44], which Gács, Ahlswede, and Körner leveraged to prove strong converses for two classic coding problems.
Concentration of measure has natural connections to isoperimetric inequalities. In fact, it can be seen as a variant version of isoperimetric problem concerning on sets with probability measure $1/2$ in which the boundary of a set is set to the $r$-neighbourhood of the set with $r>0$. A landmark methodological contribution in concentration of measure was made by Marton [44], who was the first to introduce information-theoretic techniques (especially transport-entropy inequalities) to the study of concentration of measure, yielding an elegant and concise proof for a dimension-free bound on concentration of measure. Talagrand extended this idea by developing a new transport-entropy inequality, adapting it to Gaussian measures and the Euclidean metric [63]; this argument has since become a standard reference, featured in prominent textbooks [38, 57, 64]. Gozlan and Léonard further refined this approach by replacing the “linear” transport-entropy inequality in Marton’s argument with a “nonlinear” version, obtaining a dimension-free sharp bound on concentration of measure—exponentially tight, as the exponent of the bound is asymptotically attained [31]. Prior to Gozlan and Léonard’s work, Alon et al. [4] are the first to characterize the sharp convergence exponent for concentration of measure, but limited to finite spaces. Gozlan [30] also used Marton’s framework to prove the equivalence between the Gaussian bound and Talagrand’s transport-entropy inequality. Additionally, Dembo [23] proposed a new class of transport-entropy inequalities, which he used them to recover and generalize several of Talagrand’s key results in [62]. Furthermore, E. Milman [45] established an equivalence between isoperimetry and concentration of measure. There are many existing works investigating isoperimetric inequalities or concentration of measure in the discrete setting, e.g., on various graphs, e.g., [43, 1, 17, 26].
Ahlswede and Zhang [3] and Yu [68] contributed to the study of isoperimetry with thick boundaries for exponentially small sets. This isoperimetric variant refines the concentration of measure phenomenon by relaxing the set size requirement to allow arbitrarily exponentially small measures, yet it behaves fundamentally differently from—and is far more intricate than—concentration of measure. This is because, in contrast to the latter, the isoperimetric variant admits no dimension-independent bounds. Ahlswede and Zhang [3] focused on finite spaces for which a key tool called the inherently typical subset lemma was developed by them. Yu [68] generalized their result to Polish spaces (complete, separable metric spaces), expanding the problem’s applicability, especially to the setting considered in this paper (i.e., the isoperimetric problem on a manifold with density), which plays a core role in our proof in the present paper.
  • Log-Sobolev Inequalities
Ledoux [37, 39] proposed a semigroup approach to derive isoperimetric inequalities from hypercontractivity (which is equivalent to log-Sobolev inequalities). The semigroup approach can be used to establish isoperimetric inequalities for the whole range of volumes, but the constant obtained via this method is in general suboptimal. In contrast, the constant in our work is asymptotically optimal. Otto and Villani [54] established that log-Sobolev inequalities imply transport-entropy inequalities, bridging log-Sobolev theory with concentration of measure. Building on this, Ledoux [40] (alongside Milman [45]) used a semigroup approach to derive isoperimetric inequalities from concentration inequalities (equivalently, transport-entropy inequalities). As a nonlinear log-Sobolev inequality for the standard heat semigroups (with Lebesgue measures as stationary measures), the entropic isoperimetric inequality was investigated in [20, 24] which is in fact equivalent to the standard log-Sobolev inequality for Ornstein–Uhlenbeck semigroups (with Gaussian measures as stationary measures). Nonlinear log-Sobolev inequalities for general heat (diffusion) semigroups were investigated by F. Y. Wang [66], Bakry, Gentil, and Ledoux [7] (called the entropy-energy inequality therein), as well as Polyanskiy and Samorodnitsky [55]. Specifically, F. Y. Wang [66] showed that nonlinear log-Sobolev inequalities imply nonlinear transport-entropy inequalities, extending Otto and Villani’s linear version to the nonlinear setting. Bakry, Gentil, and Ledoux [7] and Polyanskiy and Samorodnitsky [55] applied nonlinear log-Sobolev inequalities to strengthen hypercontractive inequalities. Though nonlinear Sobolev inequalities have been studied in the literature not directly to address asymptotic isoperimetry, an exact equivalence between these two concepts does exist—one we establish in the present work. Furthermore, the equivalence between isoperimetry and Sobolev-type inequalities was investigated in [16].
1.4. Proof Idea and Paper Organization
Log-Sobolev Variant Isoperimetry TransportIsoperimetry OT IT GMT IT + OT or GMT
Figure $1.$ Proof strategy of Theorem 1, where OT, IT, and GMT respectively refer to optimal transport, information theory, and geometric measure theory.
Ledoux’s semigroup method, as developed in [37, 39], seems insufficient for establishing Theorem 1. This limitation stems from the method’s core strategy: it reduces the isoperimetric problem to the more general noise stability problem. However, obtaining a sufficiently precise solution to the former demands an exact (or sufficiently precise, not merely asymptotic) solution to the latter—an outcome that remains elusive in general. While asymptotically optimal solutions for the noise stability problem are well-established [53, 34, 69, 67], a complete exact resolution has yet to be achieved. Accordingly, a novel approach is required to advance the proof of Theorem 1.
The heuristic reasoning behind our idea is as follows. Given an arbitrary set in a high-dimensional space, its (weighted) surface area does not admit any explicit formula and thus, is difficult to analyze. How can we convert the surface area to a simple formula? A possible way is to tailor the set to a new one that is “regular” in a certain sense, so that the surface area of the resulting set admits a simple formula. However, this is difficult to realize in a direct way, since the surface area is very sensitive to the shape of the set, and thus any subtle change of the set could lead to drastic change in surface area.
Our proof idea of Theorem 1 is as follows, which is new and totally different from the semigroup approach. Firstly, using a generalized Heintze-Karcher theorem [49] from geometric measure theory, we convert the standard isoperimetric problem into a variant version in which the “perimeter” is measured by the size of a thick boundary of the set. Secondly, using the inherently typical subset lemma [2] and its consequence [68] from information theory, we tailor the set into a “regular” one whose thick boundary admits a simple formula. The formula we obtain is expressed in terms of nonlinear transport-entropy inequality. The feasibility of this step is due to the fact that the size of thick boundary is robust to any slight change of the set. Lastly, using nonlinear transport inequality [66, 65] from optimal transport, we simplify the resulting formula and express it in terms of nonlinear log-Sobolev inequality, leading to the equivalence between the standard isoperimetric problem and the nonlinear log-Sobolev inequality.
Our specific strategy is illustrated in Fig. 1.1, which is described in the order opposite to the one given above. Specifically, using tools from optimal transport, in Section 2 we first deduce a specific form of nonlinear transport inequality from the nonlinear log-Sobolev inequality. By tools from information theory, we show in Section 3 that the nonlinear transport inequality in turn implies a variant isoperimetric inequality—the isoperimetric inequality with thick boundaries. Lastly, by tools from geometric measure theory, we show in Section 4 that the variant isoperimetric inequality implies the original isoperimetric inequality. Putting everything together, in Section 5 we deduce that the nonlinear log-Sobolev inequality implies the isoperimetric inequality. As for the other direction, we provide two proofs in Section 6: one is based on information theory and optimal transport, and the other is based on geometric measure theory.
We prove Theorem 2, Theorem 4, Theorem 5, Proposition 2, and Proposition 3 respectively in Appendices A–E.
2. From Log-Sobolev to Transport
2.1. From Log-Sobolev to Transport
In this section, we connect log-Sobolev inequality and transport-entropy inequality. Recall that $\mathcal{M}$ is a Riemannian manifold, equipped with a Borel reference measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ with $V\in C^{2}(\mathcal{M})$. Let $\mu$ be a probability measure on $\mathcal{M}$, whose density w.r.t. $\nu$ is $\rho$. Recall that the entropy and the Fisher information are respectively
$D(\mu\|\nu)$ $:=\int\rho\ln\rho{\text{\rm d}}\nu,$
$I(\mu\|\nu)$ $:=\int\frac{|\nabla\rho|^{2}}{\rho}{\text{\rm d}}\nu.$
It is well known that Sobolev-type inequalities (including log-Sobolev inequalities) imply Talagrand’s transport inequalities; this point was shown by Otto and Villani [54]. A nonlinear analogue of this result was given by F. Y. Wang [66, Theorem 2.1] which is proven by extending Otto and Villani’s idea. Specifically, F. Y. Wang showed that if
(2.1) $\theta(D(\lambda\|\nu))\leq I(\lambda\|\nu),\;\forall\lambda\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
for some function $\theta$, then
(2.2) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\int_{0}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r,\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}).$
In the present paper, we require a variant of F. Y. Wang’s result.
Theorem 6 (From Nonlinear Log-Sobolev to Nonlinear Transport ).
Assume that (2.1) holds for some function $\theta$. Then, for any $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ and any $0\leq s\leq D(\mu\|\nu)$, there is a probability measure $\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\pi\|\nu)=s$ and
(2.3) $\mathsf{W}_{2}\left(\mu,\pi\right)\leq\int_{s}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Proof.
Define a Laplace-type operator $L=\Delta-\nabla V\cdot\nabla$. Let $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ be a probability measure with density $h=\frac{{\text{\rm d}}\mu}{{\text{\rm d}}\nu}$ such that
[…]
The heat equation $\partial_{t}\rho=L\rho$ with initial data $\rho_{0}=h$ admits a solution $\rho_{t}$; see [54]. Let $\mu_{t}=\rho_{t}\nu$. The heat equation has intimate relations with the entropy and the Wasserstein metric in the sense that the heat equation is the gradient flow of the entropy functional $D$ with respect to the Wasserstein metric $\mathsf{W}_{2}$.
By an approximation argument in [54], for every $\mu\in P_{2}^{\mathrm{ac}}(M)$, there is a sequence of probability measures $\mu_{k}$ such that the density $h_{k}:={\text{\rm d}}\mu_{k}/{\text{\rm d}}\nu$ for each $k$ satisfies (2.4) and moreover, as $k\to\infty$,
(2.5) $\mathsf{W}_{2}\left(\mu_{k},\mu\right)$ $\to 0,$
(2.6) $D(\mu_{k}\|\nu)$ $\to D(\mu\|\nu).$
Given any $0<s<D(\mu\|\nu)$, it holds that $0<s<D(\mu_{k}\|\nu)$ for all sufficiently large $k$. Following the proof of F. Y. Wang [66, Theorem 2.1] (taking the integration $\int_{t}^{\infty}$ for a proper $t$, instead of $\int_{0}^{\infty}$ [66, (2.5)]), it is obtained that for each such $k$, there is a probability measure $\pi_{k}$ (the heat flow at $t$ with initial data $\mu_{k}$) such that $D(\pi_{k}\|\nu)=s$ and
(2.7) $\mathsf{W}_{2}\left(\mu_{k},\pi_{k}\right)\leq\int_{s}^{D(\mu_{k}\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Note that (2.6) implies that $\mu_{k}$ belongs to a sublevel set of the relative entropy $\pi\mapsto D(\pi\|\nu)$, and so does $\pi_{k}$. Since the sublevel set of the relative entropy is weakly compact, $\pi_{k}$ converges to some $\pi^{*}$ weakly. Since $\mu$ has finite second moment, (2.5) implies that the sequence $\mu_{k}$ has bounded second moment. In this case, $\pi_{k}\to\pi^{*}$ weakly is equivalent to $\mathsf{W}_{2}\left(\pi_{k},\pi^{*}\right)\to 0$. Taking limits for (2.7) and applying the triangle inequality yield
(2.8) $\mathsf{W}_{2}\left(\mu,\pi^{*}\right)\leq\int_{s}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Moreover, $D(\pi^{*}\|\nu)=s$. Such $\pi^{*}$ is what we want. ∎
Corollary 1 (From Nonlinear Log-Sobolev to Nonlinear Transport for Gaussians).
Let $\gamma$ be the standard Gaussian measure on $\mathbb{R}^{k}$. Let $X_{t}=e^{-t}X+\sqrt{1-e^{-2t}}Z$ be an Ornstein–Uhlenbeck process on $\mathbb{R}^{k}$ with distribution $\mu_{t}$, where $X\sim\mu$ and $Z\sim\gamma$ are independent. Then, the following hold.
  1. (1)
    For any $\mu\in P_{2}^{\mathrm{ac}}(\mathbb{R}^{k})$ satisfying (2.4) and any $t\geq 0$,
    $\sqrt{2D(\mu_{t}\|\gamma)}+\mathsf{W}_{2}(\mu_{t},\mu)\leq\sqrt{2D(\mu\|\gamma)}.$
  2. (2)
    For any $\mu\in P_{2}^{\mathrm{ac}}(\mathbb{R}^{k})$ and any $0\leq s\leq D(\mu\|\gamma)$, there is a probability measure $\pi$ such that $D(\pi\|\gamma)=s$ and
    $\sqrt{2s}+\mathsf{W}_{2}(\pi,\mu)\leq\sqrt{2D(\mu\|\gamma)}.$
For $s=0$, the inequality above reduces to the original version of Talagrand’s transport inequality.
The importance of Theorem 6 is as follows. Given $\tau\geq 0,\mu,\nu$, consider the following optimization problem:
(2.9) $\omega(\beta):=\inf_{\pi:\mathsf{W}_{2}(\pi,\mu)\leq\beta}D(\pi\|\nu),$
which characterizes the best tradeoff between $D(\pi\|\nu)$ and $\mathsf{W}_{2}(\pi,\mu)$ when we vary $\beta\geq 0$. In other words, within the Wasserstein ball of center $\mu$ and radius $\beta$, we aim at finding a distribution $\pi$ which minimizes $D(\pi\|\nu)$. By Sanov’s theorem in large deviations theory [25], $\omega(\beta)$ is the convergence rate of the probability that the empirical measure of i.i.d. random variables $X_{1},...,X_{n}$ with each $\sim\nu$ belongs to the Wasserstein ball of center $\mu$ and radius $\beta$. This probability naturally appears in concentration of measure and isoperimetric inequalities.
The Wasserstein gradient flow of the relative entropy starting from $\mu$ forms a feasible solution to (2.9). Specifically, the Wasserstein gradient flow is a curve that evolves to decrease $D(\pi\|\nu)$ most steeply with respect to the Wasserstein metric. If it stops until touching the Wasserstein sphere of center $\mu$ and radius $\beta$, we then obtain a feasible solution to (2.9) that is locally greedy. In other words, the Wasserstein gradient flow $\mu_{t}$ yields an upper bound on $\omega(\beta)$, although in general, $\mu_{t}$ is not the globally optimal. That is,
$\omega(\beta)$ $\leq D(\mu_{T(\beta)}\|\nu),$
where $T(\beta)$ is the time $t$ such that $\mathsf{W}_{2}(\mu_{t},\mu)=\beta$, corresponding to the escape time from the Wasserstein ball of center $\mu$ and radius $\beta$. Optimization problems like the one in (2.9) will be investigated in details in the next section.
2.2. Generalization to $\Phi$-Sobolev
Theorem 6 can be generalized to the general $\Phi$-entropy setting (also known as $U$-functional or $f$-divergence). Let $\Phi:[0,+\infty)\rightarrow\mathbb{R}$ be a continuous convex function, twice differentiable on $(0,+\infty)$. The $\Phi$-entropy is defined as
$D_{\Phi}(\mu\|\nu):=\int\Phi(\rho)d\nu-\Phi(1).$
When $\rho$ is locally Lipschitz, the Fisher information associated with the $\Phi$-entropy, named the $\Phi$-Fisher information, [65] is defined as
$I_{\Phi}(\mu\|\nu):=\int\rho\Phi^{\prime\prime}(\rho)^{2}|\nabla\rho|^{2}{\text{\rm d}}\nu=\int\frac{|\nabla p(\rho)|^{2}}{\rho}{\text{\rm d}}\nu=\int\rho|\nabla\Phi^{\prime}(\rho)|^{2}{\text{\rm d}}\nu,$
where $p(r)=r\Phi^{\prime}(r)-\Phi(r)$. When $\Phi(t)=t\ln t$, the $\Phi$-entropy and the $\Phi$-Fisher information respectively reduce to the entropy and the Fisher information. The class of $\Phi$-entropy includes many important special entropies, which have found important applications in mathematics, information theory, statistics, and physics [65].
A common requirement for the function $\Phi$ is the following, which guarantees the displacement convexity of the $\Phi$-entropy [65, Chapter 16].
Definition 3 (Displacement convexity classes [54]).
The class $\mathcal{DC}_{\infty}$ is defined as the set of continuous convex functions $\Phi:\mathbb{R}_{+}\to\mathbb{R}$, twice continuously differentiable on $(0,+\infty)$, such that $\Phi(0)=0$, and, with the notation
$p(r)=r\Phi^{\prime}(r)-\Phi(r),\quad p_{2}(r)=rp^{\prime}(r)-p(r),$
$\Phi$ satisfies any one of the following equivalent differential conditions:
  1. (1)
    $p_{2}\geq 0$;
  2. (2)
    $\frac{p(r)}{r}$ is a non-decreasing function of $r$;
  3. (3)
    $u(\delta):=e^{\delta}\Phi(e^{-\delta}),\;\delta\in\mathbb{R},$ is a convex function of $\delta$.
Theorem 6 can be easily generalized to the following $\Phi$-entropy version.
Theorem 7 (From Nonlinear $\Phi$-Sobolev to Nonlinear $\Phi$-Transport).
Let $\Phi\in\mathcal{DC}_{\infty}$. Assume that the $\theta$-nonlinear $\Phi$-Sobolev inequality
$\theta(D_{\Phi}(\pi\|\nu))\leq I_{\Phi}(\pi\|\nu),\;\forall\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
holds for some function $\theta:[0,\infty]\to[0,\infty]$. Further assume that the equation $\partial_{t}\rho=Lp(\rho)$ for an initial distribution $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ admits a solution $\mu_{t}=\rho_{t}\nu$ satisfying Assumptions (a) to (f) in Theorem 24.2 of [65] (called “smooth” in [65]), where $L=\Delta-\nabla V\cdot\nabla$ is a Laplace-type operator. Then, it holds that
(2.10) $\mathsf{W}_{2}\left(\mu,\mu_{t}\right)\leq\int_{D_{\Phi}(\mu_{t}\|\nu)}^{D_{\Phi}(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
In particular, letting $t\to\infty$ yields that
(2.11) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\int_{0}^{D_{\Phi}(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
3. From Transport to Variant Isoperimetry
3.1. From Transport to Variant Isoperimetry
We now consider a general setting for the isoperimetric problem. Let $\mathcal{X}$ and $\mathcal{Y}$ be Polish spaces (i.e., separable completely metrizable spaces, including complete Riemannian manifolds). Let $c:\mathcal{X}\times\mathcal{Y}\to[0,+\infty)$ be lower semi-continuous, which is called a cost function. Denote $\mathcal{X}^{n}$ (resp. $\mathcal{Y}^{n}$) as the $n$-fold product space of $\mathcal{X}$ (resp. $\mathcal{Y}$). Given $c$, we consider an additive cost function $c_{n}$ on $\mathcal{X}^{n}\times\mathcal{Y}^{n}$ given by
$c_{n}(x,y):=\sum_{i=1}^{n}c(x_{i},y_{i})\quad\mbox{for }(x,y)\in\mathcal{X}^{n}\times\mathcal{Y}^{n}.$
For a set $A\subseteq\mathcal{X}^{n}$, denote its $t$-enlargement under $c_{n}$ as
(3.1) $A^{c_{n}\leq t}:=\{\mathbf{y}\in\mathcal{Y}^{n}:c_{n}(\mathbf{x},\mathbf{y})\leq t,\mathbf{x}\in A\}.$
In particular, if $c=d^{2}$ on the Riemannian manifold $\mathcal{M}$, then $c_{n}^{1/2}$ corresponds to the product distance $d_{n}$. Let $\nu_{X}$ and $\nu_{Y}$ be probability measures on $\mathcal{X}$ and $\mathcal{Y}$ respectively, and let $\nu_{X}^{\otimes n}$ and $\nu_{Y}^{\otimes n}$ be the corresponding product measures.
Define the isoperimetric function as for $a\in[0,1],t\geq 0$,
(3.2) $\Gamma_{n}(a,t):=\inf_{A:\nu_{X}^{\otimes n}(A)\geq a}\nu_{Y}^{\otimes n}(A^{c_{n}\leq t}).$
We set $a=e^{-n\alpha}$ and $t=n\tau$. That is, we focus on exponentially small sets and enlarge them by a linear cost. Define the isoperimetric exponent as for $\alpha,\tau\geq 0$,
(3.3) $E_{n}(\alpha,\tau)$ $:=-\frac{1}{n}\ln\Gamma_{n}(e^{-n\alpha},n\tau).$
The optimal transport cost between two probability measures $\pi_{X}$ and $\pi_{Y}$ is defined as
(3.4) $\mathsf{C}(\pi_{X},\pi_{Y}):=\min_{\pi\in\Pi(\pi_{X},\pi_{Y})}\int c(x,y){\text{\rm d}}\pi(x,y),$
where $\Pi(\pi_{X},\pi_{Y})$ is the set of joint probability measures on $\mathcal{X}\times\mathcal{Y}$ with marginals equal to $\pi_{X},\pi_{Y}$ respectively. The existence of the minimizers are well-known; see, e.g., [64, Theorem 1.3]. The minimization problem in (3.4) is called the Monge–Kantorovich’s optimal transport problem [64]. If $\mathcal{X}=\mathcal{Y}$, and $d$ is a complete metric on $\mathcal{X}$ that induces the topology on this space (i.e., $(\mathcal{X},d)$ is a Polish metric space), and $c=d^{p}$ with $p\geq 1$, then $\mathsf{W}_{p}(\pi_{X},\pi_{Y}):=\mathsf{C}(\pi_{X},\pi_{Y})^{1/p}$ is the $p$-th Wasserstein metric between $\pi_{X}$ and $\pi_{Y}$.
For a probability measure $\pi_{W}$ and two transition probability $\pi_{X|W}$ and $\nu_{X|W}$, the conditional relative entropy of $\pi_{X|W}$ w.r.t. $\nu_{X|W}$ given $\pi_{W}$ is defined as
(3.5) $D(\pi_{X|W}\|\nu_{X|W}|\pi_{W}):=\int D(\pi_{X|W=w}\|\nu_{X|W=w}){\text{\rm d}}\pi_{W}(w).$
The conditional optimal transport cost between $\pi_{X|W}$ and $\pi_{Y|W}$ given $\pi_{W}$ is defined as
(3.6) $\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W}):=\int\mathsf{C}(\pi_{X|W=w},\pi_{Y|W=w}){\text{\rm d}}\pi_{W}(w).$
In particular, for $c=d^{p}$, we denote
(3.7) $\mathsf{W}_{p}(\pi_{X|W},\pi_{Y|W}|\pi_{W}):=\left(\int\mathsf{W}_{p}^{p}(\pi_{X|W=w},\pi_{Y|W=w}){\text{\rm d}}\pi_{W}(w)\right)^{1/p}.$
An important example is the following one.
Example 5 (Metric-Valued Cost).
$\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $p\geq 1$, i.e., $\mathsf{C}=\mathsf{W}_{p}^{p}$.
To characterize the asymptotic concentration exponent, we need the following assumption. Define the $(\mathcal{X},\epsilon)$-smooth cost function w.r.t. $c$ as
$c_{\mathcal{X},\epsilon}(x,y):=\inf_{x^{\prime}:d(x,x^{\prime})\leq\epsilon}c(x^{\prime},y).$
By definition, $c_{\mathcal{X},\epsilon}(x,y)\leq c_{\mathcal{X},0}(x,y)=c(x,y)$, and by the lower semicontinuity of $c$, $\lim_{\epsilon\downarrow 0}c_{\mathcal{X},\epsilon}(x,y)\geq c(x,y)$. So, $\lim_{\epsilon\downarrow 0}c_{\mathcal{X},\epsilon}(x,y)=c(x,y)$ pointwise.
Assumption 3.
We assume that there is a function $\delta(\epsilon):(0,\infty)\to(0,\infty)$ vanishing as $\epsilon\downarrow 0$ such that
(3.8) $c_{\mathcal{X},\epsilon}(x,y)\geq(1-\delta(\epsilon))c(x,y)-\delta(\epsilon),\;\forall(x,y).$
Assumption 3 is satisfied by Example 5. We now verify this point.
Proof.
Case of $\epsilon\leq d(x,y)$: For $x^{\prime}$ such that $d(x,x^{\prime})\leq\epsilon$, we have
(3.9) $d^{p}(x^{\prime},y)$ $\geq(d(x,y)-d(x^{\prime},x))^{p}\geq(d(x,y)-\epsilon)^{p}\geq\frac{1}{(1+\epsilon)^{p-1}}d^{p}(x,y)-\epsilon,$
where the last inequality follows by Jensen’s inequality:
$\frac{1}{1+\epsilon}\cdot(d(x,y)-\epsilon)^{p}+\frac{\epsilon}{1+\epsilon}\cdot 1^{p}\geq\left(\frac{d(x,y)}{1+\epsilon}\right)^{p}.$
We set $\delta(\epsilon)=\max\{1-\frac{1}{(1+\epsilon)^{p-1}},\epsilon\}$. Then, the inequalities in (3.9) imply (3.8).
Case of $\epsilon>d(x,y)$: As shown above, the inequality in (3.8) holds for $\epsilon=d(x,y)$. Note that the right hand side of (3.8) is increasing in $d(x,y)$ and the LHS is always zero for the case $d(x,y)\leq\epsilon$. So, for all $d(x,y)\leq\epsilon$, (3.8) still holds. ∎
Based on the above notation, we introduce the following quantity $\psi$ that is particularly important in large deviations analysis for isoperimetry.
$\psi(\alpha,\tau):=\psi(\alpha,\tau|\nu_{X},\nu_{Y})$ $:=\sup_{\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}$
(3.10) $\qquad\inf_{\pi_{Y|W}:\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\tau}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}),$
where $W$ is defined on $\{0,1\}$, and $P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ denotes the set of probability measures concentrated on compact sets. The infimum above actually represents a nonlinear transport inequality. By the convexity of optimal transport cost and the relative entropy, one can easily see that $\tau\mapsto\psi(\alpha,\tau)$ is convex. Define
(3.11) $\psi_{\epsilon,\kappa}(\alpha,\tau)$ $:=\psi_{\epsilon,\kappa}(\alpha,\tau|\nu_{X},\nu_{Y})$
(3.12) $:=\sup_{\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}$
$\qquad\inf_{\pi_{Y|XW}:\mathbb{E}_{\pi}[c(X,Y)]\leq\tau,\,\mathbb{E}_{\pi}[c^{1+\epsilon}(X,Y)]\leq\kappa}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}).$
Without changing the optimal value, the range of $W$ in (3.11) can be changed to any larger but finite set. This point can be proven by following similar proof steps as [68, Theorem 20].
Based on $\psi$ and $\psi_{\epsilon,\kappa}$, the asymptotic expression of $E_{n}$ is characterized in our previous work.
Theorem 8 (Asymptotic Exponent [68, Theorem 12]).
The following hold.
  1. (1)
    Assume that $\nu_{X}$ is compactly supported and Assumption 3 holds. Then, for any $\alpha\geq 0,\tau>0$, it holds that
    (3.13) $\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\lim_{\epsilon\downarrow 0,\kappa\uparrow\infty}\psi_{\epsilon,\kappa}(\alpha,\tau),$
    where the limit at the RHS can be replaced by $\inf_{\epsilon>0,\kappa>0}$.
  2. (2)
    Assume that $\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $1\leq p<\infty$. Then, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi:=\left\{(\alpha,\tau):\psi(\alpha,\tau)<\infty\right\}$, it holds that
    $\liminf_{n\to\infty}E_{n}(\alpha,\tau)\geq\psi(\alpha,\tau).$
Remark 7.
Statement 1 in this theorem follows from the original proof of Statement 1 in [68, Theorem 12] with two slight modifications. Firstly, we replace the power 2 in the equation chain in (56) of [68] by $q=1+\epsilon$, and then apply the Marcinkiewicz–Zygmund inequality to upper bound $\mathbb{E}_{\pi}\left[|c_{n}(x^{n},Y^{n})-n\mu|^{q}\right]$ in terms of $\mathbb{E}_{\pi}\left[c(X,Y)^{q}\right]$ (or $\kappa$). The Marcinkiewicz–Zygmund inequality here refers to that for $1<q<2$, $\mathbb{E}\left[|\sum_{i=1}^{n}X_{i}|^{q}\right]\leq B_{q}\mathbb{E}\left[(\sum_{i=1}^{n}|X_{i}|^{2})^{q/2}\right]\leq B_{q}\sum_{i=1}^{n}\mathbb{E}\left[|X_{i}|^{q}\right]$ for independent $X_{i}$’s such that $\mathbb{E}\left[X_{i}\right]=0$, where $B_{q}$ is a constant only depending on $q$. Secondly, Assumption 3 is modified in the present form in order to include Example 1 (with unbounded metrics). Statement 2 is also given in [68, Theorem 12].
Remark 8.
By checking the proof, Theorem 8 still holds if the constraint $\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ is replaced by $\pi_{XW}\in P(\mathcal{X}\times\{0,1\})$.
The inherently typical subset lemma in [2] plays a key role in the proof of the theorem above, which enables us to tailor a set into a new one that is “regular” in some sense, so that we can analyze the asymptotic expression of the boundary size of the set.
Why is the asymptotic exponent determined by the relative entropy and the Wasserstein metric? The intuition is as follows. For a distribution $\nu$, denote $B_{\epsilon]}(\nu)$ as the closed ball of center $\nu$ and radius $\epsilon$ under the Lévy–Prokhorov metric. Let $\mathsf{L}_{n}:x\in\mathcal{X}^{n}\mapsto\mathsf{L}_{x}\in P(\mathcal{X})$ be the empirical measure map, where $\mathsf{L}_{x}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ is the empirical measure of $x$, and $P(\mathcal{X})$ is the set of probability measures on $\mathcal{X}$. For a Polish space $\mathcal{X}$ and $\epsilon>0$, the empirically $\epsilon$-typical set of $\pi$ [47] is defined as
(3.14) $\mathcal{T}_{\epsilon}^{(n)}(\pi):=\mathsf{L}_{n}^{-1}(B_{\epsilon]}(\pi)).$
For an empirically $\epsilon$-typical set $A=\mathsf{L}_{n}^{-1}(B_{\epsilon}(\pi_{X}))$ (i.e., $A=\{\mathbf{x}\in\mathcal{X}^{n}:\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$), roughly speaking, by Sanov’s theorem it holds that $\nu_{n}(A)\approx e^{-nD(\pi_{X}\|\nu)}$ for $\epsilon$ chosen sufficiently small, which is no smaller than $e^{-n\alpha}$ if $D(\pi_{X}\|\nu)\leq\alpha$. The enlargement of this set $A^{c_{n}\leq n\tau}$ can be written as
$A^{c_{n}\leq n\tau}$ $=\{\mathbf{y}:\frac{1}{n}c_{n}(\mathbf{x},\mathbf{y})\leq\tau,\exists\mathbf{x}\in A\}$
$=\{\mathbf{y}:\mathbb{E}_{\mathsf{L}_{\mathbf{x},\mathbf{y}}}[c(X,Y)]\leq\tau,\mathsf{L}_{\mathbf{x}}\approx\pi_{X},\exists\mathbf{x}\}$
(3.15) $=\{\mathbf{y}:\mathsf{C}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})\leq\tau,\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$
$\approx\mathsf{L}^{-1}\{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau\},$
where (3.15) is due to Birkhoff’s theorem, $\min_{\sigma}\mathbb{E}_{\mathsf{L}_{\sigma(\mathbf{x}),\mathbf{y}}}[c(X,Y)]=\mathsf{C}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})$, with $\sigma$ denoting a permutation operator [64, Page 5]. By Sanov’s theorem again,
$\nu^{\otimes n}(A^{c_{n}\leq n\tau})\approx\exp\{-n\inf_{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau}D(\pi_{Y}\|\nu)\}.$
Optimizing over $\pi_{X}$ such that $D(\pi_{X}\|\nu)\leq\alpha$, we obtain the isoperimetric exponent for $A$,
$\sup_{\pi_{X}:D(\pi_{X}\|\nu)\leq\alpha}\inf_{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau}D(\pi_{Y}\|\nu).$
By product construction of such $A$’s (forming a conditional empirically typical set), we obtain the exponent $\psi(\alpha,\tau)$. That is, conditional empirically $\epsilon$-typical sets are optimal in attaining the optimal exponent.
Statement 1 in Theorem 8 holds for Example 5 since Assumption 3 is satisfied in this case.
Theorem 9 (Asymptotic Exponent for Metric-Valued Cost and Compactly Supported $\nu_{X}$).
Assume that $\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $1\leq p<\infty$. Assume that $\nu_{X}$ is compactly supported and $\mathbb{E}_{\nu_{Y}}[d(Y,x_{0})^{q}]<\infty$ for some $q>p$ and $x_{0}\in\mathcal{X}$. Then, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi$,
(3.16) $\lim_{n\to\infty}E_{n}(\alpha,\tau)=\psi(\alpha,\tau),$
where
$\psi(\alpha,\tau)=\psi(\alpha,\tau|\nu_{X},\nu_{Y})$ $=\sup_{\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}$
(3.17) $\qquad\inf_{\pi_{Y|W}:\mathsf{W}_{p}^{p}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\tau}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}).$
In particular, if further assume $\nu=\nu_{X}=\nu_{Y}$, then (3.16) holds for any $\alpha\geq 0,\tau\geq 0$.
Proof.
We prove this theorem by using Theorem 8. By Statement 2 of Theorem 8, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi:=\left\{(\alpha,\tau):\psi(\alpha,\tau)<\infty\right\}$, it holds that
$\liminf_{n\to\infty}E_{n}(\alpha,\tau)\geq\psi(\alpha,\tau).$
We next prove the upper bound. Note that the setting considered here satisfies the assumptions in Statement 1 of Theorem 8, which implies that the upper bound in (3.13) is still valid in this setting. Hence, it suffices to show that for sufficiently large $\kappa$,
$\psi_{\epsilon,\kappa}(\alpha,\tau)=\psi(\alpha,\tau),$
where $\epsilon=q/p-1>0$.
Let $F$ be the support of $\nu_{X}$ which is compact. Let
$P_{\kappa}:=\{\pi_{Y}:\mathbb{E}_{\pi_{Y}}[d(Y,x_{0})^{q}]\leq\kappa\}.$
We first give a dual formula for
$\theta_{\tau,\kappa}(\pi_{XW}):=\inf_{\pi_{Y|W}:\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\tau,\,\pi_{Y|W=w}\in P_{\kappa},\forall w}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}).$
Observe that
$\theta_{\tau,\kappa}(\pi_{XW})$ $=\inf_{\pi_{Y|W=w}\in P_{\kappa},\forall w}\sup_{\lambda\geq 0}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W})$
(3.18) $\qquad+\lambda(\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W})-\tau)$
$=\sup_{\lambda\geq 0}\inf_{\pi_{Y|W=w}\in P_{\kappa},\forall w}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W})$
(3.19) $\qquad+\lambda(\mathbb{E}_{\pi_{W}}[\mathsf{C}(\pi_{X|W}(\cdot|W),\pi_{Y|W}(\cdot|W))]-\tau)$
$=\sup_{\lambda\geq 0}\inf_{\pi_{Y|W=w}\in P_{\kappa},\forall w}\mathbb{E}_{\pi_{W}}[D(\pi_{Y|W}(\cdot|W)\|\nu_{Y})$
(3.20) $\qquad+\lambda(\sup_{f+g\leq c}\pi_{X|W}(f|W)+\pi_{Y|W}(g|W)-\tau)]$
$=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\inf_{\pi_{Y|W=w}\in P_{\kappa}}\sup_{f+g\leq c}D(\pi_{Y|W=w}\|\nu_{Y})$
(3.21) $\qquad+\lambda(\pi_{X|W=w}(f)+\pi_{Y|W=w}(g)-\tau)]$
$=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\sup_{f+g\leq c}\inf_{\pi_{Y|W=w}\in P_{\kappa}}D(\pi_{Y|W=w}\|\nu_{Y})$
(3.22) $\qquad+\lambda(\pi_{X|W=w}(f)+\pi_{Y|W=w}(g)-\tau)]$
$=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\sup_{f}\inf_{\gamma_{Y}\in P_{\kappa}}D(\gamma_{Y}\|\nu_{Y})$
(3.23) $\qquad+\lambda(\pi_{X|W=w}(f)+\gamma_{Y}(f^{c})-\tau)]$
where
  • the inf and sup are swapped in (3.19) which follows by the general minimax theorem [52, Theorem 5.2.2] together with the convexity of the relative entropy and the optimal transport cost functional;
  • (3.20) follows by the Kantorovich duality with $f,g$ denoting bounded continuous functions;
  • in (3.21) $\inf_{\pi_{Y|W}}$ is taken in a pointwise way;
  • the inf and sup are swapped in (3.22) which follows by the general minimax theorem [52, Theorem 5.2.2] by identifying that 1) the optimal value of the sup-inf in (3.22) is finite (since upper bounded by $\lambda(\mathsf{C}(\pi_{X|W=w},\nu_{Y})-\tau)$), and 2) by choosing $f,g$ as the zero function, the objective function turns to be $\pi_{Y|W=w}\mapsto D(\pi_{Y|W=w}\|\nu_{Y})-\lambda\tau$ whose sublevels are compact under the weak topology;
  • (3.23) follows since the optimal $g$ in (3.22) is
    $g(y)=f^{c}(y):=\inf_{x}c(x,y)-f(x).$
We consider the following optimization problem:
(3.24) $\inf_{\gamma_{Y}\in P_{\kappa}}D(\gamma_{Y}\|\nu_{Y})+\lambda\gamma_{Y}(g),$
where $g=f^{c}$ and $\lambda\geq 0$. Since adding $f$ by any constant will not change the objective function value in (3.23), without loss of generality, we assume $g(x_{0})=0$.
We now prove that the optimal value of the problem above remains the same for all $\kappa\geq\kappa_{0}$, where $\kappa_{0}$ is a constant independent of $\lambda,g$. To this end, we first bound $g$. By the triangle inequality, we have that
$g(y_{1})-g(y_{2})$ $\leq\sup_{x\in F}d^{p}(x,y_{1})-d^{p}(x,y_{2})\leq\left(d(x_{0},y_{1})+d(F)\right)^{p},$
and
$g(y_{1})-g(y_{2})$ $\geq\inf_{x\in F}d^{p}(x,y_{1})-d^{p}(x,y_{2})$
$\geq\left([d(x_{0},y_{1})-d(F)]^{+}\right)^{p}-\left(d(x_{0},y_{2})+d(F)\right)^{p},$
where $d(F)$ denotes the diameter of $F$ and $x]^{+}:=\max\{x,0\}$. Setting $y_{2}$ to the constant point $x_{0}$ yields that
$\left([d(x_{0},y)-d(F)]^{+}\right)^{p}-\left(d(F)\right)^{p}\leq g(y)-g(x_{0})$ $\leq\left(d(x_{0},y)+d(F)\right)^{p}.$
Since we assume $g(x_{0})=0$, we have
$g^{-}(y)\leq g(y)$ $\leq g^{+}(y),$
with
$g^{-}(y)$ $:=\left([d(x_{0},y)-d(F)]^{+}\right)^{p}-d(F)^{p},$
$g^{+}(y)$ $:=\left(d(x_{0},y)+d(F)\right)^{p}.$
The optimization problem in (3.23) is an information projection problem, whose solution is known to be $\gamma_{Y}^{(\lambda)}$ given by $\frac{{\text{\rm d}}\gamma_{Y}^{(\lambda)}}{{\text{\rm d}}\nu_{Y}}=\frac{e^{-\lambda g}}{\nu_{Y}(e^{-\lambda g})}$ [22]. Here for brevity, we use the notation $\nu(f)=\mathbb{E}_{\nu}[f]$ for a measure $\nu$ and a function $f$. For this distribution,
$\gamma_{Y}^{(\lambda)}(g^{q})=\frac{\nu_{Y}(g^{q}e^{-\lambda g})}{\nu_{Y}(e^{-\lambda g})}.$
Compute its derivative as
$\frac{{\text{\rm d}}}{{\text{\rm d}}\lambda}\gamma_{Y}^{(\lambda)}(g^{q})$ $=-\frac{\nu_{Y}(g^{q+1}e^{-\lambda g})\nu_{Y}(e^{-\lambda g})-\nu_{Y}(ge^{-\lambda g})\nu_{Y}(g^{q}e^{-\lambda g})}{\nu_{Y}(e^{-\lambda g})^{2}}$
$=-\left(\gamma_{Y}^{(\lambda)}(g^{q+1})-\gamma_{Y}^{(\lambda)}(g)\gamma_{Y}^{(\lambda)}(g^{q})\right)\leq 0,$
where the last inequality follows by the Harris inequality for the random variable $g(Y)$ with $Y\sim\gamma_{Y}^{(\lambda)}$ and increasing functions $t\mapsto t$ and $t\mapsto t^{q}$. Hence, $\gamma_{Y}^{(\lambda)}(g^{q})$ is decreasing in $\lambda\geq 0$. That is, $\gamma_{Y}^{(\lambda)}(g^{q})\leq\gamma_{Y}^{(0)}(g^{q})=\nu_{Y}(g^{q})\leq\nu_{Y}((g^{+})^{q})=:N.$ By the assumption $\nu_{Y}(d(x_{0},\cdot)^{q})<\infty$, it holds that $N<\infty$ since $g^{+}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$. So, $\gamma_{Y}^{(\lambda)}(g^{q})\leq N$ is bounded. Moreover, due to the fact that $g^{-}\leq g$, it holds that $\gamma_{Y}^{(\lambda)}((g^{-})^{q})\leq N$ is also bounded.
Since $g^{-}(y)^{q}\sim d(x_{0},y)^{q}$ as $d(x_{0},y)\to\infty$, we have $g^{-}(y)^{q}\geq(1-\epsilon)d(x_{0},y)^{q}-C$ for some positive constants $\epsilon,C$. Therefore,
$\gamma_{Y}^{(\lambda)}((g^{-})^{q})\geq(1-\epsilon)\gamma_{Y}^{(\lambda)}(d(x_{0},\cdot)^{q})-C,$
which implies that $\gamma_{Y}^{(\lambda)}(d(x_{0},\cdot)^{q})\leq\frac{N+C}{1-\epsilon}$ is bounded. Moreover, the upper bound $\frac{N+C}{1-\epsilon}$ is independent of $\pi_{XW},\lambda,f$. When $\kappa$ is set to any value greater than $\frac{N+C}{1-\epsilon}$, the optimal value $\theta(\tau,\kappa,\pi_{XW})$ does not depend on $\kappa$, i.e.,
$\theta(\tau,\kappa,\pi_{XW})$ $=\sup_{\lambda\geq 0}\sum_{w}\pi_{W}(w)[\sup_{f\in L^{1}(F)}\inf_{\gamma_{Y}}D(\gamma_{Y}\|\nu_{Y})$
$\qquad+\lambda(\pi_{X|W=w}(f)+\gamma_{Y}(f^{c})-\tau)]$
$=\theta(\tau,\infty,\pi_{XW}).$
The second part in the theorem follows since for the case $\nu=\nu_{X}=\nu_{Y}$, it is obvious that $\psi(\alpha,\tau)\leq\alpha$ for all $\alpha\geq 0,\tau\geq 0$, and hence $\mathrm{dom}\psi=\{\alpha\geq 0,\tau\geq 0\}$. Moreover, $\psi(\alpha,0)=\alpha$ for $\alpha\geq 0$, and $\psi(0,\tau)=0$ for $\tau\geq 0$, both of which coincide $\lim_{n\to\infty}E_{n}(\alpha,\tau)$. ∎
We next extend Theorem 9 to the case of non-compactly supported $\nu_{X}$.
Theorem 10 (Asymptotic Exponent for Metric-Valued Cost and Non-Compactly Supported $\nu_{X}$).
Assume that $\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $1\leq p<\infty$. Assume that
(3.25) $\mathbb{E}_{\nu_{X}}[e^{\lambda d^{p}(X,x_{0})}]$ $<\infty,\;\forall\lambda\in(0,\infty),$
$\mathbb{E}_{\nu_{Y}}[d^{q}(Y,x_{0})]$ $<\infty,$
for some $q>p$, and $x_{0}\in\mathcal{X}$. Then, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi$, (3.16) still holds. In particular, if further assume $\nu=\nu_{X}=\nu_{Y}$, then (3.16) holds for any $\alpha\geq 0,\tau\geq 0$.
Proof.
Our proof idea is as follows. By Sanov’s theorem, we may restrict most coordinates to compact sets without altering the exponential order of the set’s size, though the remaining coordinates cannot be constrained. Fortunately, the assumption in (3.25) ensures that the measure $\nu_{X}^{\otimes n}$ exhibits stronger concentration than Gaussian measures. This allows us to cover the majority of the mass of the remaining coordinates at only a negligible cost (i.e., in terms of distance). In other words, the isoperimetric exponent is determined by the one for compact sets which was already characterized in Theorem 9. We next present the detailed proof.
Since $\mathcal{X}=\mathcal{Y}$ is Polish, any probability measure on it is tight. So, for any $\epsilon\in(0,1)$, there is a compact set $F\subseteq\mathcal{X}$ such that $\nu_{X}(F^{c})\leq\epsilon$ and $\nu_{Y}(F^{c})\leq\epsilon$. Let $\mathbf{X}=(X_{1},...,X_{n})\sim\nu_{X}^{\otimes n}$ and $Z_{i}:=\mathbbm{1}_{B^{c}}(X_{i})$, $i\in[n]$. Then, $\mathbf{Z}\sim\mathrm{Bern}(\nu_{X}(F^{c}))^{\otimes n}$. By Sanov’s theorem, for any $\epsilon^{\prime}\in(\epsilon,1)$,
$\mathbb{P}\{\sum_{i=1}^{n}Z_{i}\geq n\epsilon^{\prime}\}\leq e^{-nD(\epsilon^{\prime}\|\nu_{X}(F^{c}))}\leq e^{-nD(\epsilon^{\prime}\|\epsilon)},$
where $D(\epsilon^{\prime}\|\epsilon)=\epsilon^{\prime}\ln\frac{\epsilon^{\prime}}{\epsilon}+(1-\epsilon^{\prime})\ln\frac{1-\epsilon^{\prime}}{1-\epsilon}$ is the binary relative entropy function, and the second inequality follows since $\epsilon\mapsto D(\epsilon^{\prime}\|\epsilon)$ is decreasing for $\epsilon<\epsilon^{\prime}$. Since $\epsilon\mapsto D(\epsilon^{\prime}\|\epsilon)$ goes to infinity as $\epsilon\downarrow 0$, given any $\epsilon^{\prime}>0$, we can choose $\epsilon$ small enough so that $D(\epsilon^{\prime}\|\epsilon)>\alpha$. For example, we can choose $\epsilon=\epsilon^{\prime}e^{-1/\epsilon^{\prime 2}}$ and choose $\epsilon^{\prime}$ small enough. For any measurable set $A$ such that $\nu_{X}^{\otimes n}(A)\geq e^{-n\alpha}$, it holds that
$\mathbb{P}\{\mathbf{X}\in A,\,\sum_{i=1}^{n}Z_{i}<n\epsilon^{\prime}\}$ $\geq\mathbb{P}\{\mathbf{X}\in A\}-\mathbb{P}\{\sum_{i=1}^{n}Z_{i}\geq n\epsilon^{\prime}\}\geq e^{-n\alpha}-e^{-nD(\epsilon^{\prime}\|\epsilon)}.$
Given any $\delta>0$, for all sufficiently large $n$,
(3.26) $\mathbb{P}\{\mathbf{X}\in A,\,\sum_{i=1}^{n}Z_{i}<n\epsilon^{\prime}\}\geq e^{-n(\alpha+\delta)}.$
For a subset $\mathcal{J}\subseteq[n]$, denote $C_{\mathcal{J}}$ as the event that $X_{i}\in F$ for $i\in\mathcal{J}$ and $X_{i}\in F^{c}$ for $i\in\mathcal{J}^{c}$. Then, (3.26) can be rewritten as
(3.27) $\mathbb{P}\{\mathbf{X}\in A\cap(\bigcup_{|\mathcal{J}|\geq n(1-\epsilon^{\prime})}C_{\mathcal{J}})\}\geq e^{-n(\alpha+o(1))}.$
On the other hand, there are ${n\choose\leq n\epsilon^{\prime}}:=\sum_{i=1}^{\left\lfloor n\epsilon^{\prime}\right\rfloor}{n\choose i}$ of sets $\mathcal{J}\subseteq[n]$ such that $|\mathcal{J}|\geq n(1-\epsilon^{\prime})$. Note that by Sanov’s theorem, ${n\choose\leq n\epsilon^{\prime}}\leq e^{nH(\epsilon^{\prime})}$, where $H(\epsilon^{\prime})=-\epsilon^{\prime}\ln\epsilon^{\prime}-(1-\epsilon^{\prime})\ln(1-\epsilon^{\prime})$ is the binary entropy function of $\epsilon^{\prime}$. Combining this with (3.27) yields that
(3.28) $\max_{|\mathcal{J}|\geq n(1-\epsilon^{\prime})}\mathbb{P}\{\mathbf{X}\in A\cap C_{\mathcal{J}}\}\geq e^{-n\alpha^{\prime}},$
where $\alpha^{\prime}:=\alpha+\delta+H(\epsilon^{\prime})$. Let $\mathcal{J}^{*}$ be the optimal $\mathcal{J}$ attaining the maximum in the above equation. Without loss of generality, we assume $\mathcal{J}^{*}=[n^{*}]$ for some $n^{*}\geq n^{\prime}:=\left\lceil n(1-\epsilon^{\prime})\right\rceil$. Let $n^{\prime\prime}=n-n^{\prime}$. We denote
(3.29) $\tilde{A}:=A\cap C_{\mathcal{J}^{*}}.$
In the following, we decompose a vector $\mathbf{x}=(\mathbf{u},\mathbf{v})\in\mathcal{X}^{n}$ into first $n^{\prime}$ components $\mathbf{u}\in\mathcal{X}^{n^{\prime}}$ and last $n^{\prime\prime}$ components $\mathbf{v}\in\mathcal{X}^{n^{\prime\prime}}$.
Denote for $\mathbf{v}_{0}=(x_{0},...,x_{0})\in\mathcal{X}^{n^{\prime\prime}}$,
(3.30) $B:=B_{r}(\mathbf{v}_{0})=\{\mathbf{v}\in\mathcal{X}^{n^{\prime\prime}}:d(\mathbf{v},\mathbf{v}_{0})\leq r\}.$
Then,
$\nu_{X}^{\otimes n^{\prime\prime}}(B^{c})$ $=\mathbb{P}\{\sum_{i=n^{\prime}+1}^{n}d^{p}(X_{i},x_{0})>r^{p}\}\leq\mathbb{E}_{\nu}[e^{\lambda d^{p}(X,x_{0})}]^{n^{\prime\prime}}e^{-\lambda r^{p}}=e^{-\lambda\left(r^{p}-n^{\prime\prime}\ell_{\lambda}\right)},$
where $\ell_{\lambda}:=\frac{1}{\lambda}\ln\mathbb{E}_{\nu}[e^{\lambda d^{p}(X,x_{0})}]$ is finite for all $\lambda>0$ by assumption.
Choose $r=(n\delta^{\prime})^{1/p}$. We then have
(3.31) $\nu_{X}^{\otimes n}(B^{c})$ $\leq e^{-\lambda n\left(\delta^{\prime}-\epsilon^{\prime}\ell_{\lambda}\right)}.$
For fixed $\delta^{\prime}>0$ and $\beta>\alpha^{\prime}$, we choose $\lambda$ large enough and $\epsilon^{\prime}$ small enough so that
(3.32) $\lambda\left(\delta^{\prime}-\epsilon^{\prime}\ell_{\lambda}\right)$ $>\beta\textrm{ and }\mathbb{E}_{\nu}[d^{p}(Y,x_{0})]<\frac{\delta^{\prime}}{\epsilon^{\prime}}.$
By the law of large numbers, $\nu_{Y}^{\otimes n^{\prime\prime}}(B)\to 1$ as $n\to\infty$.
For $\mathbf{u}\in\mathcal{X}^{n^{\prime}}$, denote $\tilde{A}_{\mathbf{u}}:=\{\mathbf{v}\in\mathcal{X}^{n^{\prime\prime}}:(\mathbf{u},\mathbf{v})\in\tilde{A}\}$, which is the projection of $\tilde{A}$ onto the last $n^{\prime\prime}$ components. Denote $\hat{A}:=\{\mathbf{u}\in\mathcal{X}^{n^{\prime}}:\nu_{X}^{\otimes n^{\prime\prime}}(\tilde{A}_{\mathbf{u}})\geq e^{-n\beta}\}\subseteq F^{n^{\prime}}$. Denote $\bar{A}:=\{(\mathbf{u},\mathbf{v})\in\tilde{A}:\mathbf{u}\in\hat{A}\}$. We then have
[…]
which implies
(3.33) $\nu_{X}^{\otimes n}(\bar{A})\geq e^{-n\alpha^{\prime}}-e^{-n\beta}=e^{-n(\alpha^{\prime}+o(1))},$
with $o(1)$ denoting a term that vanishes as $n\to\infty$.
From (3.31) and (3.32), we see that $B$ and $\tilde{A}_{\mathbf{u}}$ have nonempty intersection for each $\mathbf{u}\in\hat{A}$. Hence, $(\tilde{A}_{\mathbf{u}})^{2r}:=\{\mathbf{v}^{\prime}\in\mathcal{X}^{n^{\prime\prime}}:d(\mathbf{v}^{\prime},\mathbf{v})\leq 2r,\mathbf{v}\in\tilde{A}_{\mathbf{u}}\}$ covers the whole $B$, where $r$ is the radius in (3.30) and set to $(n\delta^{\prime})^{1/p}$. So,
$A^{c_{n}\leq t}\supseteq\tilde{A}^{c_{n}\leq t}$ $=\bigcup_{\mathbf{x}\in\tilde{A}}\{\mathbf{x}\}^{c_{n}\leq t}\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times A_{\mathbf{u}}^{c_{n^{\prime\prime}}\leq 2^{p}n\delta^{\prime}}\Big)$
$\supseteq\bigcup_{\mathbf{u}\in\hat{A}}\Big(\{\mathbf{u}\}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B\Big)=\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}\times B.$
Summarizing the above, it holds that
[…]
$\nu_{Y}^{\otimes n}(A^{c_{n}\leq t})$ $\geq\nu_{Y}^{\otimes n^{\prime}}(\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}})\nu_{Y}^{\otimes n^{\prime\prime}}(B)\geq(1-o(1))\nu_{Y}^{\otimes n^{\prime}}(\hat{A}^{c_{n^{\prime}}\leq t-2^{p}n\delta^{\prime}}),$
where $o(1)$ are terms that vanish as $n\to\infty$. Setting $t=n\tau$, we then have that
$\nu_{Y}^{\otimes n}(A^{n\tau})$ $\geq(1-o(1))\inf_{\hat{A}\subseteq F^{n^{\prime}}:\nu_{X}^{\otimes n^{\prime}}(\hat{A})\geq e^{-n^{\prime}\alpha^{\prime\prime}}}\nu_{Y}^{\otimes n^{\prime}}(\tilde{A}^{c_{n^{\prime}}\leq n^{\prime}\tau^{\prime}}),$
where $\tau^{\prime}:=\frac{\tau-2^{p}\delta^{\prime}}{1-\epsilon^{\prime}}$ and $\alpha^{\prime\prime}:=\frac{\alpha+\delta+H(\epsilon^{\prime})+o(1)}{1-\epsilon^{\prime}}$. That is,
$E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})$ $\leq E_{n^{\prime}}(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})+o(1),$
where $\nu_{X|F}:=\nu_{X|X\in F}:=\nu_{X}(\cdot|F)$ is the conditional probability measures of $\nu_{X}$ on $F$.
Since $F$ is compact, applying the upper bound on the isoperimetric exponent in Theorem 9, we obtain that
(3.34) $\limsup_{n\to\infty}E_{n}(\alpha,\tau)$ $\leq\limsup_{n\to\infty}E_{n^{\prime}}(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})$
$\leq\psi(\alpha^{\prime\prime}+\ln\nu_{X}(F),\tau^{\prime}|\nu_{X|F},\nu_{Y})\leq\psi(\alpha^{\prime\prime},\tau^{\prime}|\nu_{X},\nu_{Y}).$
Letting $\epsilon^{\prime}\to 0,\delta\to 0$, and $\delta^{\prime}\to 0$, (i.e., $\alpha^{\prime\prime}\to\alpha$ and $\tau^{\prime}\to\tau$), we further obtain that
$\limsup_{n\to\infty}E_{n}(\alpha,\tau)$ $\leq\lim_{\alpha^{\prime\prime}\downarrow\alpha,\tau^{\prime}\uparrow\tau}\psi(\alpha^{\prime\prime},\tau^{\prime})=\lim_{\alpha^{\prime\prime}\downarrow\alpha}\psi(\alpha^{\prime\prime},\tau),$
where the equality follows since $\psi$ is convex and hence continuous in $\tau>0$.
We now need the following lemma whose proof is provided later.
Lemma 1.
For $\tau>0$, $\alpha\mapsto\psi(\alpha,\tau)$ is right continuous.
By this lemma, we obtain the desired bound $\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\psi(\alpha,\tau)$. ∎
We now prove Lemma 1.
Proof of Lemma 1.
Since $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\geq\psi(\alpha,\tau)$, it suffices to prove $\lim_{\alpha^{\prime}\downarrow\alpha}\psi(\alpha^{\prime},\tau)\leq\psi(\alpha,\tau)$. We now use the standard lower semi-continuity and compactness arguments to prove this point.
Denote $\mathsf{C}=\mathsf{W}_{p}^{p}$. Denote
$g(\pi_{XW},\tau):=\min_{\pi_{Y|W}:\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\tau}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}),$
which is convex since both the relative entropy and the optimal transport cost functional are convex. The minimum is attained if $g(\pi_{XW},\tau)<\infty$, since we minimize a lower semi-continuous function over a (weakly) closed set. Notice that $\psi(\alpha,\tau)$ is defined as a supremum of $g(\pi_{XW},\tau)$. So, $\tau\mapsto\psi(\alpha,\tau)$ is convex for $\tau\geq 0$, and hence continuous for $\tau>0$ and $\psi(\alpha,\tau)<\infty$.
To prove that $\alpha\mapsto\psi(\alpha,\tau)$ is continuous for $\tau\geq 0$, it suffices to prove that $\psi(\alpha_{k},\tau)\to\psi(\alpha,\tau)$ for all sequences $\alpha_{k}$ such that $\alpha_{k}\downarrow\alpha$. Denote $\pi_{XW}^{(k)}$ as a distribution approximately attaining the supremization in $\psi(\alpha_{k},\tau)$ within a gap $1/k$. Denote $\pi_{Y|W}^{(k)}$ as a conditional distribution attaining $g(\pi_{XW}^{(k)},\tau)$, i.e., the infimization in $\psi(\alpha_{k},\tau)$. Since $\pi_{XW}^{(k)}$ is in $\{\pi_{XW}:D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha\}$, a weakly compact set (e.g., [25, Lemma 6.2.12]), by passing a subsequence, we can assume that $\pi_{XW}^{(k)}$ converges to some distribution $\pi_{XW}^{*}$. Denote $\pi_{Y|W}^{*}$ as a conditional distribution attaining $g(\pi_{XW}^{*},\tau-\delta)$. Since by considering the feasible solution $\pi_{X|W}$ for the minimization in $g(\pi_{XW},\tau)$, one can observe that $g(\pi_{XW},\tau)$ is bounded by the finite number $\alpha$. So, without loss of generality, we assume that $D(\pi_{Y|W=w}^{*}\|\nu)$ is finite for both $w=0$ and $1$.
By the triangle inequality,
$\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{Y|W}^{*}|\pi_{W}^{(k)})$ $\leq\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{X|W}^{*}|\pi_{W}^{(k)})+\mathsf{W}_{p}(\pi_{Y|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{(k)}).$
Since the Wasserstein metric induced by a bounded metric $d$ is compatible with the weak topology, it holds that as $k\to\infty$, $\mathsf{W}_{p}(\pi_{X|W=w}^{(k)},\pi_{X|W=w}^{*})\to 0$ for $w\in{\mathrm{supp}}(\pi_{W}^{*}):=\{w\in\{0,1\}:\pi_{W}^{*}(w)>0\}$. Moreover, $\mathsf{W}_{p}$ is bounded, and hence, $\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{X|W}^{*}|\pi_{W}^{(k)})\to\mathsf{W}_{p}(\pi_{X|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{*})=0$ and $\mathsf{W}_{p}(\pi_{Y|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{(k)})\to\mathsf{W}_{p}(\pi_{Y|W}^{*},\pi_{X|W}^{*}|\pi_{W}^{*})\leq\tau-\delta$, as $k\to\infty$. So, for sufficiently large $k$,
$\mathsf{W}_{p}(\pi_{X|W}^{(k)},\pi_{Y|W}^{*}|\pi_{W}^{(k)})$ $\leq\tau.$
That is, $\pi_{Y|W}^{*}$ is a feasible solution to $g(\pi_{XW}^{(k)},\tau)$, which implies $\psi(\alpha_{k},\tau)-\frac{1}{k}\leq g(\pi_{XW}^{(k)},\tau)\leq D(\pi_{Y|W}^{*}\|\nu_{Y}|\pi_{W}^{(k)})$. We in turn obtain that
(3.35) $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\limsup_{k\to\infty}D(\pi_{Y|W}^{*}\|\nu_{Y}|\pi_{W}^{(k)})=D(\pi_{Y|W}^{*}\|\nu_{Y}|\pi_{W}^{*})=g(\pi_{XW}^{*},\tau-\delta).$
On the other hand, by the lower semi-continuity of the relative entropy,
$D(\pi_{X|W}^{*}\|\nu_{X}|\pi_{W}^{*})\leq\limsup_{k\to\infty}D(\pi_{X|W}^{(k)}\|\nu_{X}|\pi_{W}^{(k)})\leq\alpha.$
Hence, $\pi_{XW}^{*}$ is a feasible solution to the following supremization:
$\psi(\alpha,\tau-\delta)=\sup_{\pi_{XW}:D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}g(\pi_{XW},\tau-\delta).$
So, $g(\pi_{XW}^{*},\tau-\delta)\leq\psi(\alpha,\tau-\delta)$, which, combined with (3.35), implies that $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau-\delta).$ Letting $\delta\downarrow 0$ and by the continuity of $\tau\mapsto\psi(\alpha,\tau)$, it holds that $\limsup_{k\to\infty}\psi(\alpha_{k},\tau)\leq\psi(\alpha,\tau)$. ∎
3.2. Interpretation from Concentration of Measure
The isoperimetric problem with thick boundary admits a natural interpretation from the perspective of concentration of measure. Concentration of measure principle states that for any set of not too small probability, slightly enlarging this set will always have probability close to one. Formally, let $r(a)$ be the minimum number such that the $r$-enlargement of any set of measure $a$ will always have probability close to one. Denote $a=e^{-n\alpha}$ and $r^{2}(a)=n\tau(\alpha)$. By our results, for sufficiently large $n$, $\tau(\alpha)$ is the minimum $\tau$ such that $\psi(\alpha,\tau)=0$, where $\mathsf{C}=\mathsf{W}_{2}^{2}$. Solving it yields $\tau\geq\mathsf{C}(\pi_{X|W},\nu|\pi_{W})$ for all $\pi_{XW}$ such that $D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha$. Hence, $\tau(\alpha)$ satisfies
$\tau(\alpha)$ $=\sup_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}\mathsf{C}(\pi_{X|W},\nu|\pi_{W}),$
which is the upper concave envelope of $\tau_{0}$ given below
$\tau_{0}(\alpha):$ $=\sup_{\pi:D(\pi\|\nu)\leq\alpha}\mathsf{C}(\pi,\nu).$
The curves $\tau$ and $\tau_{0}$ respectively characterize the best convex and nonconvex tradeoffs between the relative entropy and the Wasserstein metric. Thus, they determine the optimal nonlinear transport inequalities.
Recall that $K_{C}$ is the optimal transport constant in the following inequality: $\mathsf{W}_{2}(\pi,\nu)\leq\sqrt{\frac{2}{K_{C}}D(\pi\|\nu)},\;\forall\pi.$ Hence, $\tau(\alpha)\leq\frac{2\alpha}{K_{C}},\;\forall\alpha>0,$ and the constant $K_{C}$ cannot be improved. This implies that for sufficiently large $n$, enlarging any set of measure $\exp(-\frac{K_{C}r^{2}}{2})$ by the distance $r$ will always make it have probability close to one, and moreover, the constant $K_{C}$ cannot be improved. In contrast, another (but indeed equivalent) interpretation of $K_{C}$ in the concentration of measure is that for all $A$ such that $\nu_{n}(A)=1/2$, it always holds that $\nu_{n}((A^{r})^{c})\leq\exp(-\frac{K_{C}r^{2}}{2}),\;\forall r>0$ for some constant $C$. Here, the constant $K_{C}$ cannot be improved either.
4. From Variant Isoperimetry to Standard Isoperimetry
We now turn back to the standard isoperimetric problem in the Riemannian manifold. The “boundary” in the definition of perimeter in this setting is infinitely thin. We now connect the isoperimetric problem with thick “boundary” considered in the last section to this setting.
Recall that $\mathcal{M}$ is a $k$-dimensional Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol},V\in C^{2}(\mathcal{M})$ and the induced geodesic distance $d$. Recall the boundary measure defined in (1.2). The boundary measure can be alternatively expressed as
$\nu_{n}^{+}(A)$ $=\liminf_{r\downarrow 0}\frac{\nu_{n}(A^{r})-\nu_{n}(A)}{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}\cdot\frac{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}{r}$
(4.1) $=\nu_{n}(A)\liminf_{r\downarrow 0}\frac{\ln[\nu_{n}(A^{r})/\nu_{n}(A)]}{r}.$
However, to apply Theorem 14, we need keep $r$ to be large (in order of $\sqrt{n}$), and hence, we need remove $\liminf_{r\downarrow 0}$ in the last line. To this end, we need some nice properties of isoperimetric minimizers from geometric measure theory.
Here we adopt the notation used in [45]. For a smooth hypersurface $S$, the second fundamental form is by $\varPi_{S,p}^{\sigma}(u,v)=g(D_{u}v,v)$ for $u,v\in T_{p}S$, where $D$ is the covariant derivative. According to this definition, the second fundamental form of the sphere in Euclidean space with respect to the outer normal $\sigma$ is positive definite. The total curvature $H_{S}^{\sigma}(p)$ of $S$ at a point $p$ is the trace of the second fundamental form $\varPi_{S,p}^{\sigma}$. (The mean curvature is obtained if it is divided by the dimension of $S$.)
Definition 4.
The $\nu$-total curvature of a smooth hypersurface $S$ at a point $p\in S$ with respect to the unit normal vector $\sigma$, denoted $H_{S,\nu}^{\sigma}(p)$, is defined as
$H_{S,\nu}^{\sigma}(p):=H_{S}^{\sigma}(p)-\langle\nabla V(p),\sigma(p)\rangle.$
Bayle [10] and Morgan [49, 51] showed that isoperimetric minimizers always exist and the isoperimetric profile is concave. See an explicit explanation in [45, p. 216].
Theorem 11 (Morgan [48, 49, 51] and Bayle [10]).
Under Assumption 2, the following hold.
  1. (1)
    For any $a\in(0,1)$, there always exists an open isoperimetric minimizer $A$ in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a$ such that: (a) its boundary $\partial A$ consists of disjoint regular part and singular part; (b) the regular part is a $C^{2}$-smooth hypersurface (since $V\in C^{2}(\mathcal{M})$) which has a constant $\nu_{n}$-total curvature, denoted by $H_{\nu_{n}}(A)$; (c) the singular part has dimension strictly lower than that of the regular part, more precisely, lower than or equal to the dimension of the regular part minus $7$.
  2. (2)
    The isoperimetric profile $I_{n}$ is a concave function on $[0,1]$.
  3. (3)
    For the isoperimetric minimizer given in Statement 1, the constant $\nu_{n}$-total curvature $H_{\nu_{n}}(A)$ of the regular part of its boundary satisfies
    $\lim_{\epsilon\downarrow 0}\frac{I_{n}(a+\epsilon)-I_{n}(a)}{\epsilon}\leq H_{\nu_{n}}(A)\leq\lim_{\epsilon\downarrow 0}\frac{I_{n}(a)-I_{n}(a-\epsilon)}{\epsilon}.$
    In particular, if $I_{n}$ is differentiable at $a$, then $H_{\nu_{n}}(A)=I_{n}^{\prime}(a)$.
Theorem 12 (Generalized Heintze-Karcher due to Morgan [49, Theorem 2, Remark 3]).
Assume that for $\kappa\in\mathbb{R}$,
$\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq\kappa g.$
Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 11. Let $S$ denote the regular part of $\partial A$, and let $H_{\nu_{n}}(A)$ denote the constant $\nu_{n}$-total curvature of $S$ with respect to the outer unit normal vector field $\sigma$ on $S$. Then for any $r>0$,
$\nu_{n}(A^{r})-\nu_{n}(A)\leq\nu_{n}^{+}(A)\int_{0}^{r}\exp\left(H_{\nu_{n}}(A)t-\frac{\kappa}{2}t^{2}\right){\text{\rm d}}t.$
Note that the sign of the total curvature $H_{S}^{\sigma}(p)$ here is opposite to the one in [49, Theorem 2, Remark 3]. Based on the two theorems above, we build a bridge between the standard isoperimetric problem and the isoperimetric problem with thick boundary.
Theorem 13.
Assume that Assumption 1 holds. Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 11. Then, for any $r>0$,
$\frac{I_{n}(a)}{a}=\frac{\nu_{n}^{+}(A)}{\nu_{n}(A)}\geq\frac{1}{r}\ln\frac{\nu_{n}(A^{r})}{\nu_{n}(A)}.$
Proof.
Let $S$ denote the regular part of $\partial A$, and let $H_{\nu_{n}}(A)$ denote the constant $\nu_{n}$-total curvature of $S$ with respect to the outer unit normal vector field $\sigma$ on $S$. Let
$P(r)$ $:=\nu_{n}^{+}(A)\exp\left(H_{\nu_{n}}(A)r\right),$
$V(r)$ $:=\nu_{n}(A)+\int_{0}^{r}P(t){\text{\rm d}}t.$
Define g(r):=V(r)P(0)P(r)V(0),g(r):=V(r)P(0)-P(r)V(0), whose derivative satisfies
(4.2) $g^{\prime}(r)=P(r)P(0)-P^{\prime}(r)V(0)=P(r)P(0)-H_{\nu_{n}}(A)P(r)V(0).$
From Theorem 11, it is known that
$H_{\nu_{n}}(A)\leq\lim_{\epsilon\downarrow 0}\frac{I_{n}(a)-I_{n}(a-\epsilon)}{\epsilon},$
where $a=\nu_{n}(A)$. Since $I_{n}$ is concave,
$\lim_{\epsilon\downarrow 0}\frac{I_{n}(a)-I_{n}(a-\epsilon)}{\epsilon}\leq\frac{I_{n}(a)}{a}=\frac{P(0)}{V(0)}.$
Hence, $H_{\nu_{n}}(A)\leq\frac{P(0)}{V(0)},$ which, combined with (4.2), implies $g^{\prime}(r)\geq 0.$
Observe that $g(0)=0$. So, for all $r\geq 0$, it holds that $g(r)\geq 0,$ i.e., $\frac{P(r)}{V(r)}\leq\frac{P(0)}{V(0)}.$ Noting that $P(r)=V^{\prime}(r)$, we then have that $\ln\frac{V(r)}{V(0)}\leq\frac{P(0)}{V(0)}r.$ From Theorem 12, it is known that $V(r)\geq\nu_{n}(A^{r})$. We hence conclude that
$\frac{\nu^{+}(A)}{\nu(A)}=\frac{P(0)}{V(0)}\geq\frac{1}{r}\ln\frac{V(r)}{V(0)}\geq\frac{1}{r}\ln\frac{\nu_{n}(A^{r})}{\nu_{n}(A)}.$
5. Putting Everything Together: From Log-Sobolev to Standard Isoperimetry
In this section, we combine all the tools built in Sections 24 to prove the lower bound in our large deviations theorem, i.e., to deduce standard isoperimery from nonlinear log-Sobolev inequalities. To this end, we first deduce variant isoperimery from nonlinear log-Sobolev inequalities by combining Theorem 6 and Theorem 10. Define
$\Upsilon(s,\alpha)=\int_{s}^{\alpha}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Theorem 14 (From Nonlinear Log-Sobolev to Variant Isoperimetry).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that nonlinear log-Sobolev inequality
$\theta(D(\pi\|\nu))\leq I(\pi\|\nu),\;\forall\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
holds for some convex $\theta$. Assume $\Upsilon(0,\alpha)<\infty$ for some (thus all) $0<\alpha<\infty$. Let $E_{n}$ and $\psi$ be defined in Theorem 10 (for variant isoperimetry) with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. Then, under Assumption 2, for any $\alpha,\tau\geq 0$,
(5.1) $\lim_{n\to\infty}E_{n}(\alpha,\tau)\leq\psi(\alpha,\tau)$ $\leq\sup_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}>0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}\leq\alpha\end{subarray}}\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*},$
where $(s_{0}^{*},s_{1}^{*})=(0,0)$ if
(5.2) $\lambda\Upsilon^{2}(0,\alpha_{1})+(1-\lambda)\Upsilon^{2}(0,\alpha_{0})\leq\tau,$
and otherwise, $(s_{0}^{*},s_{1}^{*})=(s_{0}^{*}(\lambda,\alpha_{0},\alpha_{1}),s_{1}^{*}(\lambda,\alpha_{0},\alpha_{1}))$ is the unique solution to
(5.3) $\Upsilon(s_{0},\alpha_{0})$ $=\sqrt{\frac{\tau\theta(s_{0})}{\lambda\theta(s_{1})+(1-\lambda)\theta(s_{0})}},$
(5.4) $\Upsilon(s_{1},\alpha_{1})$ $=\sqrt{\frac{\tau\theta(s_{1})}{\lambda\theta(s_{1})+(1-\lambda)\theta(s_{0})}}.$
Moreover, the upper bound in (5.1) is strictly increasing in $\alpha$, and hence, the constraint $\lambda\alpha_{1}+(1-\lambda)\alpha_{0}\leq\alpha$ can be replaced by $\lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha$.
Proof.
Obviously, Assumption 2 implies $\mathbb{E}_{\nu}[d^{q}(X,x_{0})]<\infty$ for any $q\geq 1$. By Theorem 10, the first inequality in (5.1) is obtained. We next prove the second inequality.
Denote $\alpha_{i}=D(\pi_{X|W=i}\|\nu)$, $i\in\{0,1\}$. By Theorem 6, given $\pi_{X|W}$ satisfying $\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),i\in\{0,1\}$ and given every $s_{i}\leq\alpha_{i}$, there is $\pi_{Y|W}$ such that $\pi_{Y|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M})$, $s_{i}=D(\pi_{Y|W=i}\|\nu)$, and
$\mathsf{W}_{2}(\pi_{X|W=i},\pi_{Y|W=i})\leq\Upsilon(s_{i},\alpha_{i}),\;i\in\{0,1\}.$
Hence, it holds that
(5.5) $\psi(\alpha,\tau)$ $\leq\sup_{\begin{subarray}{c}\lambda\in[0,1],\alpha_{0},\alpha_{1}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}\leq\alpha\end{subarray}}\min_{\begin{subarray}{c}s_{0}\in[0,\alpha_{0}],s_{1}\in[0,\alpha_{1}]:\\ \lambda\Upsilon^{2}(s_{1},\alpha_{1})+(1-\lambda)\Upsilon^{2}(s_{0},\alpha_{0})\leq\tau\end{subarray}}\lambda s_{1}+(1-\lambda)s_{0}.$
Observe that when $\alpha$ increases, we can enlarge $s$ to maintain $\Upsilon(s,\alpha)$ unchanged. This implies the RHS above is strictly increasing in $\alpha$.
Observe that $\Upsilon^{2}(s,\alpha)$ is convex in $s$, which implies that given $\lambda\in[0,1],\alpha_{0},\alpha_{1}\geq 0$, the inner minimization in (5.5) is a convex optimization program. We then apply the Lagrangian multiplier method to obtain necessary conditions for optimal solutions. We first assume $\lambda\in(0,1)$. For this case, the optimal solution is $(0,0)$ if (5.2) is satisfied, and otherwise, any optimal solution $(s_{0},s_{1})$ must satisfy
(5.6) $\lambda\Upsilon^{2}(s_{1},\alpha_{1})+(1-\lambda)\Upsilon^{2}(s_{0},\alpha_{0})$ $=\tau,$
(5.7) $\frac{\Upsilon(s_{1},\alpha_{1})}{\sqrt{\theta(s_{1})}}$ $=\frac{\Upsilon(s_{0},\alpha_{0})}{\sqrt{\theta(s_{0})}}.$
These two equations can be rewritten as the ones in (5.3) and (5.4).
We now prove that there is a unique solution to the equations (5.6) and (5.7) (or (5.3) and (5.4)). Observe that given any $\alpha_{i}$, $s\mapsto\frac{\Upsilon(s,\alpha_{i})}{\sqrt{\theta(s)}}$ is decreasing in $s\in(0,\alpha_{i}]$. So, from (5.7), for each $s_{1}$, there is unique $s_{0}=s_{0}^{*}(s_{1})$ satisfying (5.7). Moreover, $s_{0}^{*}(s_{1})$ is increasing in $s_{1}$, which implies that
$s_{1}\mapsto\lambda\Upsilon^{2}(s_{1},\alpha_{1})+(1-\lambda)\Upsilon^{2}(s_{0}^{*}(s_{1}),\alpha_{0})$
is decreasing in $s_{1}\in(0,\alpha_{1}]$. So, there is a unique solution $s_{1}=s_{1}^{*}$ ensuring this function equal to $\tau$, which in turn implies that $(s_{1}^{*},s_{0}^{*}(s_{1}^{*})$) is the unique solution to the equations in (5.6) and (5.7). This yields the second inequality in (5.1) under the assumption $\lambda\in(0,1)$.
If $\lambda=1$, then the optimal $s_{1}^{*}$ for the inner minimization is $s_{1}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{1})$, and $s_{0}^{*}$ can be arbitrary, where $\Upsilon_{1}^{-1}$ is the inverse of $\Upsilon$ w.r.t. the first argument when the second is fixed to be $\alpha_{1}$. Similarly, if $\lambda=0$, then the optimal $s_{0}^{*}$ for the inner minimization is $s_{0}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{0})$ and $s_{1}^{*}$ can be arbitrary. For both these two cases, the upper bound in (5.5) cannot be greater than $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. By continuity, one can find that the most right hand side in (5.1) is greater than or equal to $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. So, the second inequality in (5.1) still holds for $\lambda=0$ or $1$. ∎
We now deduce standard isoperimery from nonlinear log-Sobolev inequalities. Define
(5.8) $\xi(\alpha):=\limsup_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}},$
where $\psi$ is defined in (3.17) but with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. As a consequence of Theorem 13 and Theorem 10 (or Theorem 14), we can lower bound the asymptotics of the isoperimetric profile $I_{n}$ in terms of $\xi$ (or $\breve{\Theta}$). Recall that $\breve{\Theta}$ defined in (1.7) is the function characterizing the optimal tradeoff in the nonlinear log-Sobolev inequality. We obtain the lower bound part of Theorem 1. Recall $\underline{\Lambda}$ defined in (1.5).
Theorem 15 (From Nonlinear Log-Sobolev to Standard Isoperimetry).
Under Assumptions 1 and 2, for any $\alpha>0$,
(5.9) $\underline{\Lambda}(\alpha)$ $\geq\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$
Proof.
Let $A_{n}$ be an isoperimetric minimizer of size $e^{-n\alpha}$ in the $n$-dimensional space. From Theorem 13, we see that for any $r>0$,
$\frac{\ln[\nu_{n}(A_{n}^{r})/\nu_{n}(A_{n})]}{r}\leq\frac{\nu_{n}^{+}(A_{n})}{\nu_{n}(A_{n})}=\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}}.$
Setting $r=\sqrt{n\tau}$ and then letting $n\to\infty$, we obtain for any $\tau>0$,
$\underline{\Lambda}(\alpha)$ $\geq\liminf_{n\to\infty}\frac{\ln[\nu_{n}(A_{n}^{\sqrt{n\tau}})/\nu_{n}(A_{n})]}{n\sqrt{\tau}}\geq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}},$
where the last inequality follows from Theorem 10. Letting $\tau\to 0$, we obtain the first inequality in (5.9).
We now prove the second inequality in (5.9). If $\underline{\Lambda}(\alpha)=\infty$, then there is nothing to prove. So, we assume $\underline{\Lambda}(\alpha)<\infty$. By the first inequality in (5.9), $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. Since $\psi(\alpha,\tau)\leq\alpha$, it holds that $\lim_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. On the other hand, by Theorem 14,
(5.10) $\psi(\alpha,\tau)$ $\leq\sup_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}>0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha\end{subarray}}\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*},$
where $(s_{0}^{*},s_{1}^{*})$ is $(0,0)$ or the unique solution to (5.3) and (5.4). Therefore, given any $0<\epsilon<\alpha$, for sufficiently small $\tau$, the constraint $\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon$ can be posed into the supremization above without changing the optimal value. So, $(s_{0}^{*},s_{1}^{*})=(0,0)$ is excluded. These yield the following bound
$\xi(\alpha)$ $\geq\inf_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha,\\ \lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon\end{subarray}}\frac{\lambda\left(\alpha_{1}-s_{1}^{*}\right)+(1-\lambda)\left(\alpha_{0}-s_{0}^{*}\right)}{\sqrt{\tau}},$
where $(s_{0}^{*},s_{1}^{*})$ is the unique solution to (5.3) and (5.4).
Observe that $\breve{\Theta}$ is increasing. The solution $(s_{0}^{*},s_{1}^{*})$ to (5.3) and (5.4) satisfies
$\frac{\alpha_{1}-s_{1}^{*}}{\sqrt{\breve{\Theta}(s_{1}^{*})}}\geq\int_{s_{1}^{*}}^{\alpha_{1}}\frac{1}{\sqrt{\breve{\Theta}(t)}}{\text{\rm d}}t$ $=\sqrt{\frac{\tau\breve{\Theta}(s_{1}^{*})}{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}},$
which implies
$\frac{\alpha_{1}-s_{1}^{*}}{\sqrt{\tau}}$ $\geq\frac{\breve{\Theta}(s_{1}^{*})}{\sqrt{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}}.$
Similarly,
$\frac{\alpha_{0}-s_{0}^{*}}{\sqrt{\tau}}\geq\frac{\breve{\Theta}(s_{0}^{*})}{\sqrt{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}}.$
Combining these two inequalities and utilizing the convexity of $\breve{\Theta}$ yield that
$\frac{\lambda\left(\alpha_{1}-s_{1}^{*}\right)+(1-\lambda)\left(\alpha_{0}-s_{0}^{*}\right)}{\sqrt{\tau}}$ $\geq\sqrt{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}$
$\geq\sqrt{\breve{\Theta}(\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*})}.$
Under the condition $\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon$, the last line above is further lower bounded by $\sqrt{\breve{\Theta}(\alpha-\epsilon)}$. Therefore, $\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha-\epsilon)}$. Letting $\epsilon\to 0$ yields $\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}$. ∎
6. The Other Direction: From Standard Isoperimetry to Log-Sobolev
We now prove the other direction (i.e., the upper bound) in Theorem 1. Recall $\xi$ defined in (5.8) and $\bar{\Lambda}$ defined around (1.5).
Theorem 16 (From Isoperimetry to Nonlinear Log-Sobolev).
Under Assumption 1, for any $\alpha>0$,
(6.1) $\bar{\Lambda}(\alpha)$ $\leq\lim_{t\downarrow\alpha}\xi(t)\leq\sqrt{\breve{\Theta}(\alpha)}.$
We provide two distinct proofs for this theorem. One is to use the tools of the variant isoperimetry and the famous HWI inequality. The other one is to apply the coarea formula and the large deviations theory.
First Proof of Theorem (16).
We first prove the first inequality in (6.1). Let $A$ be a set approximately attaining $\Gamma_{n}(e^{-n\alpha},n\tau)$ for $\nu=\nu_{X}=\nu_{Y}$ (given in (3.2)) within a factor $1-\epsilon$ where $\epsilon>0$, i.e., $A$ has probability $e^{-n\alpha}$ and $\nu_{n}(A^{\sqrt{n\tau}})\geq(1-\epsilon)\Gamma_{n}(e^{-n\alpha},n\tau)$. Let $r:=\sqrt{n\tau}$. Since $f(t):=\ln\nu_{n}(A^{t})$ is monotone, it is differentiable almost everywhere on $(0,r)$. As a basic property of monotone functions, $f^{\prime}$ is integrable over $[0,r]$ and $\int_{0}^{r}f^{\prime}(t){\text{\rm d}}t\leq f(r)-f(0)$ (see e.g., [60, Corollary 4 on p.113]), which implies that there is some $r_{0}\in(0,r)$ such that $\frac{f(r)-f(0)}{r}\geq f^{\prime}(r_{0}),$i.e.,
$\frac{\ln[\nu_{n}(A^{r})/\nu_{n}(A)]}{r}\geq\frac{\nu_{n}^{+}(A^{r_{0}})}{\nu_{n}(A^{r_{0}})}.$
Since $\nu_{n}^{+}(A^{r_{0}})\geq I_{n}(\nu_{n}(A^{r_{0}}))$ and $I_{n}(a)/a$ is non-increasing in $a$, it holds that
$\frac{\ln[\nu_{n}(A^{r})/\nu_{n}(A)]}{r}\geq\frac{I_{n}(\nu_{n}(A^{r_{0}}))}{\nu_{n}(A^{r_{0}})}\geq\frac{I_{n}(\nu_{n}(A^{r}))}{\nu_{n}(A^{r})}.$
Noting that $r=\sqrt{n\tau}$ and letting $n\to\infty$, by Statement 2 in Theorem 8, we obtain for any $\tau>0$,
$\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$ $\geq\lim_{n\to\infty}\frac{\ln[\nu_{n}(A^{\sqrt{n\tau}})/\nu_{n}(A)]}{n\sqrt{\tau}}\geq\limsup_{n\to\infty}\frac{I_{n}(a_{n,\tau})}{a_{n,\tau}\sqrt{n}},$
where $a_{n,\tau}=\nu_{n}(A^{\sqrt{n\tau}})$.
Since $A$ approximately attains $\Gamma_{n}(e^{-n\alpha},n\tau)$ within a factor $1-\epsilon$, by Statement 2 in Theorem 8 again, it holds that $\liminf_{n\to\infty}-\frac{1}{n}\ln a_{n,\tau}\geq\psi(\alpha,\tau).$ Hence, for any $\delta>0$, $\bar{\Lambda}(\psi(\alpha,\tau)-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$
We can assume that $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$, since otherwise, $\xi(\alpha)=\infty$ and nothing is needed to prove. Hence, given any $\delta>0$, it holds that $\psi(\alpha,\tau)>\alpha-\delta.$ for all sufficiently small $\tau$, which, combined with the inequality above, implies $\bar{\Lambda}(\alpha-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$ Letting $\tau\downarrow 0$ yields that
(6.2) $\bar{\Lambda}(\alpha-\delta)$ $\leq\xi(\alpha).$
Since $\delta>0$ is arbitrary, we obtain the first inequality in (6.1).
We next prove the second inequality in (6.1). Substituting the expression of $\psi(\alpha,\tau)$ into the bound above yields that
$\xi(\alpha)$ $=\liminf_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$
$\leq\lim_{\tau\downarrow 0}\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}$
$\qquad\sup_{\pi_{Y|W}:\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\sqrt{\tau}}\frac{\alpha-D(\pi_{Y|W}\|\nu|\pi_{W})}{\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})}$
$\leq\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})=\alpha}\lim_{\tau\downarrow 0}$
$\qquad\sup_{\pi_{Y|W}:\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\sqrt{\tau}}\frac{D(\pi_{X|W}\|\nu|\pi_{W})-D(\pi_{Y|W}\|\nu|\pi_{W})}{\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})},$
where in the last line, we swap the infimization and the limit, and we restrict the inequality constraint in the infimization to be an equality.
Denote $\alpha_{i}=D(\pi_{X|W=i}\|\nu),\beta_{i}=D(\pi_{Y|W=i}\|\nu)$, and $\gamma_{i}=\mathsf{W}_{2}(\pi_{X|W=i},\pi_{Y|W=i})$, $i\in\{0,1\}$. Then, by the Cauchy–Schwarz inequality,
$D(\pi_{X|W}\|\nu|\pi_{W})-D(\pi_{Y|W}\|\nu|\pi_{W})=\mathbb{E}_{\pi}[\alpha_{W}-\beta_{W}]$
$\qquad\leq\sqrt{\mathbb{E}_{\pi}[\left(\frac{\alpha_{W}-\beta_{W}}{\gamma_{W}}\right)^{2}]\mathbb{E}_{\pi}[\gamma_{W}^{2}]}.$
So,
(6.3) $\xi(\alpha)$ $\leq\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})=\alpha}\lim_{\tau\downarrow 0}\sup_{\pi_{Y|W}:\mathbb{E}_{\pi}[\gamma_{W}^{2}]\leq\tau}\sqrt{\mathbb{E}_{\pi}[\left(\frac{\alpha_{W}-\beta_{W}}{\gamma_{W}}\right)^{2}]}.$
Since $\pi_{Y|W=i}=\pi_{X|W=i}$ is a feasible solution to the supremum above, we have $\beta_{i}\leq\alpha_{i},i\in\{0,1\}$.
Define the descending slope of the relative entropy w.r.t. the Wasserstein metric as
$|\partial^{-}D|(\mu\|\nu)$ $:=\limsup_{\pi\to\mu}\frac{[D(\mu\|\nu)-D(\pi\|\nu)]^{+}}{\mathsf{W}_{2}(\pi,\mu)}$
$=\lim_{\tau\downarrow 0}\sup_{\pi:0<\mathsf{W}_{2}(\pi,\mu)\leq\tau}\frac{[D(\mu\|\nu)-D(\pi\|\nu)]^{+}}{\mathsf{W}_{2}(\pi,\mu)}.$
Using this notation, we rewrite (6.3) as
$\xi(\alpha)$ $\leq\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})=\alpha}\sqrt{\mathbb{E}_{W\sim\pi_{W}}[\left(|\partial^{-}D|(\pi_{X|W}\|\nu)\right)^{2}]}.$
We now need the HWI inequality for the Riemannian manifold setting [65, Corollary 20.13]. The original version for Euclidean spaces is due to Otto and Villani [54].
Theorem 17 (HWI Inequality [65, Corollary 20.13]).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that Assumption 1 holds. Then, for any two probability measures $\mu_{0}=\rho_{0}\nu,\mu_{1}=\rho_{1}\nu$ in $P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\mu_{1}\|\nu)<\infty$ and $\rho_{0}$ is Lipschitz, it holds that
$D(\mu_{0}\|\nu)-D(\mu_{1}\|\nu)\leq\mathsf{W}_{2}(\mu_{0},\mu_{1})\sqrt{I(\mu_{0}\|\nu)}.$
This theorem implies $|\partial^{-}D|(\mu\|\nu)\leq\sqrt{I(\mu\|\nu)}.$ In fact, this is an equality; see [6, Theorem 9.3]. We restrict $\pi_{X|W=i},i\in\{0,1\}$ to be in $P_{2}^{\mathrm{ac}}(\mathcal{M})$. Therefore,
$\xi(\alpha)$ $\leq\inf_{\begin{subarray}{c}\pi_{XW}:\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),\forall i,\\ D(\pi_{X|W}\|\nu|\pi_{W})=\alpha\end{subarray}}\sqrt{I(\pi_{X|W}\|\nu|\pi_{W})}=\sqrt{\breve{\Theta}(\alpha)}.$
Combining this with (6.2) yields that $\bar{\Lambda}(\alpha-\delta)\leq\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$ That is, for $\alpha,\delta>0$, $\bar{\Lambda}(\alpha)\leq\xi(\alpha+\delta)\leq\sqrt{\breve{\Theta}(\alpha+\delta)}.$ Since $\breve{\Theta}(t)$ is continuous in $t>0$, letting $\delta\downarrow 0$ yields $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$
Second Proof of Theorem (16).
By an approximation argument, we may assume that the probability measures $\mu$ in the nonlinear log-Sobolev inequality have smooth density $f:={\text{\rm d}}\mu/{\text{\rm d}}\nu$. Obviously, $\int f\,d\nu=1$. Let $f^{\otimes n}$ be the $n$-fold product of $f$ with itself, which satisfies ${\text{\rm d}}\mu_{n}/{\text{\rm d}}\nu_{n}=f^{\otimes n}$ with $\mu_{n}=\mu^{\otimes n}$. Then, the coarea formula (see e.g., [19, Exercise III.12]) states:
(6.4) $\int|\nabla f^{\otimes n}|\,d\nu_{n}=\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$
where $\varphi_{n}(\mathbf{x})=e^{-\sum_{i=1}^{n}V(x_{i})}$ be the density function of $\nu_{n}$ w.r.t. the volume form on the manifold $\mathcal{M}^{n}$, $C_{s}=\{\mathbf{x}\in\mathcal{M}^{n}:f^{\otimes n}(\mathbf{x})=s\}$ be the level set of the function $f^{\otimes n}$, ${\text{\rm d}}\mathrm{\mathrm{area}}$ be the induced volume form (Hausdorff measure of dimension equal to the dimension of $\mathcal{M}^{n}$ minus $1$) on the level set $C_{s}$, and $|\nabla f^{\otimes n}(\mathbf{x})|$ be the norm of the gradient of $f^{\otimes n}$ at the point $\mathbf{x}\in\mathcal{M}^{n}$. For brevity, we denote $D:=D(\mu\|\nu)$.
Observe that
$\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s\geq\int_{e^{n(D-\delta)}}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$
(6.5) $\qquad=\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}^{+}\{f^{\otimes n}\geq s\}{\text{\rm d}}s\geq\eta_{n}\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s,$
where
$\eta_{n}:=\inf_{\begin{subarray}{c}s\geq e^{n(D-\delta)}:\\ \nu_{n}\{f^{\otimes n}\geq s\}>0\end{subarray}}\frac{\nu_{n}^{+}\{f^{\otimes n}\geq s\}}{\nu_{n}\{f^{\otimes n}\geq s\}}.$
Note that
$\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s=\int_{0}^{\infty}\int 1\{f^{\otimes n}\geq s,\,s\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}{\text{\rm d}}s$
(6.6) $\qquad=\int(f^{\otimes n}-e^{n(D-\delta)})\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}=\gamma_{1,n}-\gamma_{2,n},$
where
$\gamma_{1,n}$ $:=\int f^{\otimes n}\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}=\mu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\},$
and
$\gamma_{2,n}:=e^{n(D-\delta)}\nu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\}.$
By the law of large numbers, $\gamma_{1,n}\to 1$ as $n\to\infty$. By the large deviations theory (or the asymptotic equipartition property in information theory),
$\limsup_{n\to\infty}-\frac{1}{n}\ln\gamma_{2,n}=\sup_{t>0}t\left(D_{1-t}(\mu\|\nu)-(D(\mu\|\nu)-\delta)\right)>0,$
where $D_{1-t}(\mu\|\nu)=-\frac{1}{t}\ln\int f^{1-t}d\nu,\;f={\text{\rm d}}\mu/{\text{\rm d}}\nu$ is the (1t)(1-t)-Rényi divergence of $\mu$ w.r.t. $\nu$. Therefore, substituting these into (6.6) yields $\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s\to 1,\textrm{ as }n\to\infty,$ which, combined with (6.5), implies
$\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$ $\geq(1+o(1))\eta_{n}.$
We now lower bound $\eta_{n}$. Observe that
$\eta_{n}$ $\geq\inf_{\begin{subarray}{c}s\geq e^{n(D-\delta)}:\\ \nu_{n}\{f^{\otimes n}\geq s\}>0\end{subarray}}\frac{I_{n}(\nu_{n}\{f^{\otimes n}\geq s\})}{\nu_{n}\{f^{\otimes n}\geq s\}}=\frac{I_{n}(a_{n})}{a_{n}},$
where $a_{n}:=\nu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\}$, and the equality follows since $I_{n}(a)/a$ is non-increasing in $a$. Hence,
$\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$ $\geq(1+o(1))\frac{I_{n}(a_{n})}{a_{n}}.$
One can see that
$a_{n}$ $=\int_{\frac{{\text{\rm d}}\mu_{n}}{{\text{\rm d}}\nu_{n}}\geq e^{n(D-\delta)}}{\text{\rm d}}\nu_{n}\leq\int_{\frac{{\text{\rm d}}\mu_{n}}{{\text{\rm d}}\nu_{n}}\geq e^{n(D-\delta)}}e^{-n(D-\delta)}{\text{\rm d}}\mu_{n}\leq e^{-n(D-\delta)}=:b_{n}.$
By again the fact that $I_{n}(a)/a$ is non-increasing in $a$, we have
$\limsup_{n\to\infty}\frac{I_{n}(a_{n})}{a_{n}\sqrt{n}}\geq\limsup_{n\to\infty}\frac{I_{n}(b_{n})}{b_{n}\sqrt{n}}=\bar{\Lambda}(D-\delta).$
Therefore, for any $\delta>0$,
(6.7) $\limsup_{n\to\infty}\frac{1}{\sqrt{n}}\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$ $\geq\bar{\Lambda}(D(\mu\|\nu)-\delta).$
On the other hand, by the Cauchy–Schwarz inequality,
(6.8) $\int|\nabla f^{\otimes n}|\,{\text{\rm d}}\nu_{n}\leq\sqrt{\int f^{\otimes n}\,{\text{\rm d}}\nu_{n}\,\cdot\int\frac{|\nabla f^{\otimes n}|^{2}}{f^{\otimes n}}\,{\text{\rm d}}\nu_{n}}=\sqrt{nI(\mu\|\nu)}.$
Combining (6.4), (6.7), and (6.8) yields that $\sqrt{I(\mu\|\nu)}\geq\bar{\Lambda}(D(\mu\|\nu)-\delta),\,\forall\mu.$ Noting that $\bar{\Lambda}$ is non-decreasing, we have $\sqrt{\Theta(\alpha)}\geq\bar{\Lambda}(\alpha-\delta).$ That is,
(6.9) $\bar{\Lambda}(\alpha)\leq\sqrt{\Theta(\alpha+\delta)}.$
By substitution $\nu\leftarrow\nu^{\otimes k}$, we define
$\bar{\Lambda}_{k}(\alpha):=\limsup_{n\to\infty}\frac{I_{kn}(e^{-kn\alpha})}{e^{-kn\alpha}\sqrt{n}}.$
Noting that both $n\mapsto I_{n}(a)$ and $a\mapsto I_{n}(a)/a$ are non-increasing, it holds that for $kn\leq n^{\prime}<k(n+1)$,
$\frac{I_{k(n+1)}(e^{-kn\alpha})}{e^{-kn\alpha}\sqrt{k(n+1)}}\leq\frac{I_{n^{\prime}}(e^{-n^{\prime}\alpha})}{e^{-n^{\prime}\alpha}\sqrt{n^{\prime}}}\leq\frac{I_{kn}(e^{-k(n+1)\alpha})}{e^{-k(n+1)\alpha}\sqrt{kn}}.$
Taking $n\to\infty$ yields that for any $\delta>0$, $\frac{\bar{\Lambda}_{k}(\alpha-\delta)}{\sqrt{k}}\leq\bar{\Lambda}(\alpha)\leq\frac{\bar{\Lambda}_{k}(\alpha+\delta)}{\sqrt{k}}.$ Applying (6.9) to $(\mathcal{M}^{k},\nu^{\otimes k})$ yields that for all $k$,
(6.10) $\bar{\Lambda}(\alpha)\leq\frac{\bar{\Lambda}_{k}(\alpha+\delta)}{\sqrt{k}}\leq\sqrt{\frac{\Theta_{k}(k(\alpha+2\delta))}{k}},$
where $\Theta_{k}$ is defined for $(\mathcal{M}^{k},\nu^{\otimes k})$ and given by
$\Theta_{k}(\alpha):=\inf_{\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu^{\otimes k})\geq\alpha}I(\mu\|\nu^{\otimes k}).$
We now claim that
(6.11) $\lim_{k\to\infty}\frac{\Theta_{k}(k\beta)}{k}=\breve{\Theta}(\beta).$
This is because, on one hand, by the tensorization property of nonlinear log-Sobolev inequality, for all $k$ and $\beta\geq 0$, $\frac{1}{k}\Theta_{k}(k\beta)\geq\breve{\Theta}(\beta);$ on the other hand, by setting the function $f$ in the $k$-dimensional nonlinear log-Sobolev inequality to be the product form $f(x)=\prod_{i=1}^{j}f_{1}(x_{i})\prod_{i=j+1}^{k}f_{2}(x_{i})$ for some functions $f_{1}$ and $f_{2}$ defined on $\mathcal{M}$, the lower bound $\breve{\Theta}$ can be asymptotically approached.
Combining (6.10) and (6.11) yields $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha+2\delta)}$. Letting $\delta\downarrow 0$ yields the desired result $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$. ∎
Combining this theorem with Theorem 15 yields that $\underline{\Lambda}(\alpha)=\bar{\Lambda}(\alpha)=\sqrt{\breve{\Theta}(\alpha)},$ completing the proof of our main result, Theorem 1.
Appendix A: Proof of Theorem 2
In this section, we now prove the remaining unproven parts of Theorem 2.
Appendix A-1. Inequality (b) in Statement 1
We first prove Inequality (b) in Statement 1 of Theorem 2. Due to Theorem 12, or more precisely, due to (3.1) and (4.2) in [45] but with $K_{CD}\geq 0$, it holds that for any $n$,
$\frac{1}{2}-a\leq I_{n}(a)\int_{0}^{r(a)}\exp\left(\frac{I_{n}(a)}{a}t-\frac{K_{CD}}{2}t^{2}\right){\text{\rm d}}t,$
where $r(a)=\sqrt{\frac{2}{K_{C}}\ln\frac{1}{a}}$. Taking limits as $n\to\infty$, we have
(6.12) $\frac{1}{2}-a\leq I_{\inf}(a)\int_{0}^{r(a)}\exp\left(\frac{I_{\inf}(a)}{a}t-\frac{K_{CD}}{2}t^{2}\right){\text{\rm d}}t.$
By definition, $\sqrt{K_{IS}^{-}}=\liminf_{a\to 0}\frac{I_{\inf}(a)}{a\sqrt{2\ln\frac{1}{a}}}$, which implies that for any $\epsilon>0$, there is a decreasing sequence $a_{k}$ converging to 0 as $k\to\infty$ such that for all sufficiently large $k$,
(6.13) $I_{\inf}(a_{k})\leq a_{k}\sqrt{2(K_{IS}^{-}+\epsilon)\ln\frac{1}{a_{k}}}.$
Substituting $a_{k}$ and the upper bound in (6.13) into (6.12) yields
$\frac{1}{2a_{k}}-1\leq\sqrt{2(K_{IS}^{-}+\epsilon)\ln\frac{1}{a_{k}}}\int_{0}^{r(a_{k})}\exp\left(\sqrt{2(K_{IS}^{-}+\epsilon)\ln\frac{1}{a_{k}}}t-\frac{K_{CD}}{2}t^{2}\right){\text{\rm d}}t.$
Denote $\alpha_{k}=\ln\frac{1}{a_{k}}$, which diverges to $\infty$. Substituting it to the inequality above and performing change of variables, we then have
$\frac{1}{2}e^{\alpha_{k}}-1$ $\leq\sqrt{2(K_{IS}^{-}+\epsilon)}\alpha_{k}\int_{0}^{\sqrt{\frac{2}{K_{C}}}}\exp\left(\sqrt{2(K_{IS}^{-}+\epsilon)}\alpha_{k}t-\frac{K_{CD}}{2}\alpha_{k}t^{2}\right){\text{\rm d}}t.$
Note that $K_{C}\geq K_{IS}^{-}\geq K_{CD}$, which implies $K_{C}\left(K_{IS}^{-}+\epsilon\right)>K_{CD}$. By Laplace’s method, under this condition, the last line above is
$\sim\frac{\sqrt{\left(K_{IS}^{-}+\epsilon\right)K_{C}}}{\sqrt{\left(K_{IS}^{-}+\epsilon\right)K_{C}}-K_{CD}}\exp\left((2\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{C}}}-\frac{K_{CD}}{K_{C}})\alpha_{k}\right)$
as $k\to\infty$. It is asymptotically larger than $\frac{1}{2}e^{\alpha_{k}}-1$ only if $2\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{C}}}-\frac{K_{CD}}{K_{C}}\geq 1$, i.e., $K_{IS}^{-}\geq\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}$ with letting $\epsilon\to 0$.
Appendix A-2. Inequality (l)
Inequality (l) is exactly the following inequality: for any $n\geq 1$,
(6.14) $\liminf_{a\to 0}\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq\sqrt{K_{LS}^{+}}.$
Given any $n$, to prove (6.14), it suffices to show that for any sequence $\{a_{i}\}_{i\in\mathbb{N}}$ of positive numbers that converges to 0, it holds that
(6.15) $\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}\geq\sqrt{K_{LS}^{+}}.$
Let $\alpha>0$. For each $a_{i}$, one can find an integer $n_{i}$ such that $e^{-(n_{i}+1)\alpha}<a_{i}\leq e^{-n_{i}\alpha}$. As $i\to\infty$, it holds that $n_{i}\to\infty$. So, by the monotonicity of $t\mapsto\frac{I_{n}(t)}{t}$ [45, Proposition 3.1] and $t\mapsto\sqrt{\ln\frac{1}{t}}$,
$\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}\geq\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2(n_{i}+1)\alpha}}.$
Taking $\liminf_{i\to 0}$, we obtain
$\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}$ $\geq\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2(n_{i}+1)\alpha}}=\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}$
$\geq\liminf_{i\to 0}\frac{I_{n_{i}}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}\geq\sqrt{\frac{\breve{\Theta}(\alpha)}{2\alpha}},$
where the second inequality is due to that $I_{n}(a)\geq I_{m}(a)$ for all $m\geq n$ and any $a$. Since $\alpha>0$ is arbitrary, we obtain (6.15).
Appendix A-3. Statement 2
We now prove Statement 2 in Theorem 2. For brevity, we denote $K_{1}:=K_{IS,1}^{+}=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}$. For any log-concave probability measure $\nu$ on $\mathbb{R}$, Bobkov [11] showed that $I(a)=f(F^{-1}(a))\wedge f(F^{-1}(1-a))$.
By L’Hospital’s rule, as $a\to 0$,
$\frac{f(F^{-1}(a))}{I_{G}(a)}$ $=\frac{f(F^{-1}(a))}{\phi(\Phi^{-1}(a))}\sim\frac{V^{\prime}(F^{-1}(a))}{\Phi^{-1}(a)}\sim\frac{V^{\prime\prime}(F^{-1}(a))\phi(\Phi^{-1}(a))}{f(F^{-1}(a))}$
which implies $\left(\frac{f(F^{-1}(a))}{I_{G}(a)}\right)^{2}\sim V^{\prime\prime}(F^{-1}(a)).$ Similarly, $\left(\frac{f(F^{-1}(1-a))}{I_{G}(a)}\right)^{2}\sim V^{\prime\prime}(F^{-1}(1-a)).$ Hence, $\left(\frac{I(a)}{I_{G}(a)}\right)^{2}\sim V^{\prime\prime}(F^{-1}(a))\wedge V^{\prime\prime}(F^{-1}(1-a)),$ which implies $K_{1}=\lim_{x\to\infty}V^{\prime\prime}(x)\wedge\lim_{x\to-\infty}V^{\prime\prime}(x).$
Without loss of generality, we assume $\lim_{x\to\infty}V^{\prime\prime}(x)\geq\lim_{x\to-\infty}V^{\prime\prime}(x)$. So, $K_{1}=\lim_{x\to-\infty}V^{\prime\prime}(x)$, which implies $(1-\epsilon)K_{1}\leq V^{\prime\prime}(x)\leq(1+\epsilon)K_{1}$ for all $x\leq x_{0}$ and hence,
(6.16) $V^{\prime}(x_{0})+(1-\epsilon)K_{1}(x-x_{0})\leq V^{\prime}(x)\leq V^{\prime}(x_{0})+(1+\epsilon)K_{1}(x-x_{0}),$
and
$V(x_{0})+V^{\prime}(x_{0})(x-x_{0})+\frac{(1-\epsilon)K_{1}}{2}(x-x_{0})^{2}$
(6.17) $\leq$ $V(x)\leq V(x_{0})+V^{\prime}(x_{0})(x-x_{0})+\frac{(1+\epsilon)K_{1}}{2}(x-x_{0})^{2},$
for all $x\leq x_{0}$. Construct a test function $\rho(x)=\frac{g(x-b)}{f(x)}=\frac{1}{\sqrt{2\pi/K_{1}}}e^{V(x)-\frac{K_{1}(x-b)^{2}}{2}}$, where $g$ is the density of $\mathcal{N}(0,1/K_{1})$. Substituting this test function to the log-Sobolev inequality yields that
(6.18) $\int V(x+b)g(x){\text{\rm d}}x-\frac{1}{2}\ln\frac{2\pi e}{K_{1}}\leq\frac{1}{2K}\int(V^{\prime}(x+b)-K_{1}x)^{2}g(x){\text{\rm d}}x.$
Substituting (6.16) and (6.17) into (6.18) and noting that both the parts $\int_{x_{0}}^{\infty}$ of the integrals at two sides are negligible when $b$ is chosen to approach -$\infty$, we obtain
$\int\left(V(x_{0})+V^{\prime}(x_{0})(x+b-x_{0})+\frac{(1-\epsilon)K_{1}}{2}(x+b-x_{0})^{2}\right)g(x){\text{\rm d}}x-\frac{1}{2}\ln\frac{2\pi e}{K_{1}}$
$\qquad\leq\frac{1}{2K}\int\left(V^{\prime}(x_{0})+(1+\epsilon)K_{1}(x+b-x_{0})-K_{1}x\right)^{2}g(x){\text{\rm d}}x+o_{b}(1),$
i.e.,
$V(x_{0})+V^{\prime}(x_{0})(b-x_{0})+\frac{(1-\epsilon)K_{1}}{2}\left(\frac{1}{K_{1}}+(b-x_{0})^{2}\right)-\frac{1}{2}\ln\frac{2\pi e}{K_{1}}$
(6.19) $\qquad\leq\frac{1}{2K}\left(\epsilon^{2}+\left(V^{\prime}(x_{0})+(1+\epsilon)K_{1}(b-x_{0})\right)^{2}\right)+o_{b}(1),$
where $o_{b}(1)$ is a term vanishing as $b\to$-$\infty$. Letting $b\to$-$\infty$ and comparing the coefficients of $b^{2}$ at two sides, we obtain $\frac{(1-\epsilon)K}{(1+\epsilon)^{2}}\leq K_{1}.$ Noting that when $b\to$-$\infty$, the relative entropy (i.e., the LHS of (6.19)) approaches $\infty$. So, the constant $K$ can be set to $K_{LS}^{+}$. Letting $\epsilon\to 0$ yields the desired result.
Appendix B: Proof of Theorem 4
We next compare the dimension-free isoperimetric profile $I_{\inf}$ and the infinite-dimensional isoperimetric profile $I_{\infty}$. Recall that the dimension-free isoperimetric profile $I_{\inf}(a):=\lim_{n\to\infty}I_{n}(a)=\inf_{n\geq 1}I_{n}(a),$ and $I_{\infty}$ is the isoperimetric profile for the space $(\mathcal{M}^{\infty},\nu^{\otimes\infty})$.
Theorem 18.
Let $(\mathcal{X},d,\nu)$ be a locally compact Polish probability metric space. For $n\in\mathbb{N}\cup\{\infty\}$, let
(6.20) $I_{n}(a):=\inf_{\mathrm{closed}\,A:\nu_{n}(A)=a}\nu_{n}^{+}(A),\;a\in[0,1]$
be the isoperimetric profile for the $n$-product space $(\mathcal{X}^{n},\nu^{\otimes n})$, where the enlargement of a set is defined under $d_{n}(\mathbf{x},\mathbf{y}):=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$. If $I_{\inf}$ given above is continuous at $a$, then $I_{\inf}(a)=I_{\infty}(a).$
Remark 9.
We restrict the set $A$ in (6.20) to be closed, in order to ensure that $A^{r}\to A$ and hence $\nu_{n}(A^{r})\to\nu_{n}(A)$ as $r\downarrow 0$. This restriction usually does not affect the isoperimetric profile, e.g., for the smooth Riemannian manifold setting with an absolutely continuous reference measure [48].
In our Riemannian manifold setting, under Assumption 1, the isoperimetric profiles $I_{n},n\in\mathbb{N}$ are all concave, which implies their pointwise minimum $I_{\inf}$ is concave as well and hence also continuous on $(0,1)$. So, Theorem 18 implies $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in(0,1)$. Moreover, obviously $I_{\inf}(a)=I_{\infty}(a)=0$ for $a\in\{0,1\}$. Thus, $I_{\inf}$ is continuous on $[0,1]$. Theorem 4 is a consequence of Theorem 18. Hence it remains to prove Theorem 18.
Appendix B-1. Proof of Theorem 18
In the following, we also use $\mathcal{X}^{\mathbb{N}}$ to denote $\mathcal{X}^{\infty}$, the countably infinite product of $\mathcal{X}$ with itself. Recall that $\nu_{n}=\nu^{\otimes n}$ with $n\in\mathbb{N}\cup\{\infty\}$.
Observe that $I_{\infty}(a)\leq I_{\inf}(a)$, since for any $n$ and $B\subseteq\mathcal{X}^{n}$, $B\times\mathcal{X}^{\mathbb{N}}$ in $\mathcal{X}^{\mathbb{N}}$ has the same measure and the same boundary as those of $B$ in $\mathcal{X}^{n}$. We next prove that $I_{\infty}(a)\geq I_{\inf}(a)$. We follow similar proof steps for Gaussian measures in [28, Theorem 2.2.4]. We first make the following claim.
Claim 1.
Let $\pi_{n}\colon\mathcal{X}^{\mathbb{N}}\to\mathcal{X}^{n}$ be the projection $\pi_{n}(\mathbf{x})=\pi_{n}(x_{k}:k\in\mathbb{N})=(x_{1},\dots,x_{n})$. Then it holds that: (a) $\nu_{n}=\nu_{\infty}\circ\pi_{n}^{-1}$, (b) if $K\subseteq\mathcal{X}^{\mathbb{N}}$ is compact in the product topology, then $K=\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K))$, (c) $K^{r}:=\bigcup_{\mathbf{x}\in K}B_{r]}(\mathbf{x})$ is compact in the product topology if $K$ is, where $B_{r]}(\mathbf{x})$ is the closed ball with center $\mathbf{x}$ and radius $r$ under the metric $d_{\infty}$.
Proof of Claim 1.
(a) is obvious.
(b) Inclusion $K\subseteq\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))$: By definition, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$ for any $\mathbf{x}\in K$, so $\mathbf{x}\in\pi_{n}^{-1}(\pi_{n}(K))$ for all $n$. Thus, $\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
Inclusion $\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))\subseteq K$: Suppose $\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K))$. Then for each $n$, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$, so there exists $\mathbf{x}^{(n)}\in K$ such that $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$. Since $K$ is compact in the product topology of $\mathcal{X}^{\mathbb{N}}$, it is closed and sequentially compact. The sequence $(\mathbf{x}^{(n)})$ lies in $K$. The projections $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$ imply that the first $n$ coordinates of $\mathbf{x}^{(n)}$ match those of $\mathbf{x}$. This implies $\mathbf{x}^{(n)}\to\mathbf{x}$ in the product topology. Since $K$ is closed, $\mathbf{x}\in K$. Hence, $K=\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
(c) We first note that $d_{\infty}$ might be infinity between two points, and it does not induce the product topology (topology of coordinate-wise convergence) on $\mathcal{X}^{\mathbb{N}}$. In fact, the convergence in $d_{\infty}$ implies the convergence in product topology.
Let $K\subseteq\prod_{i=1}^{\infty}X_{i}$ be compact, and define
$C_{i}:=\pi_{i}(K)\subseteq X_{i},$
which is compact in $X_{i}$ since $\pi_{i}$ is continuous. For any $\mathbf{y}\in K^{r}$, there exists $\mathbf{x}\in K$ such that $\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}\leq r^{2}.$ Hence, for each coordinate $i$, $d(y_{i},C_{i})\leq d(y_{i},x_{i})\leq r,$ because $x_{i}\in C_{i}$. This implies
$y_{i}\in C_{i}^{r}:=B_{r]}(C_{i}):=\{z\in X_{i}:d(z,C_{i})\leq r\},$
which is compact since $\mathcal{X}$ is locally compact. Therefore, $K^{r}\subseteq\prod_{i=1}^{\infty}C_{i}^{r},$ and by Tychonoff’s theorem, $\prod_{i=1}^{\infty}C_{i}^{r}$ is compact in the product topology.
We now prove that $K^{r}$ is closed in the product topology. Fix $\mathbf{x}\in K$. For each $N\in\mathbb{N}$, define
$B_{r]}^{(N)}(\mathbf{x}):=\left\{\mathbf{y}:\sum_{i=1}^{N}d(x_{i},y_{i})^{2}\leq r^{2}\right\}\textrm{ and }K_{N}^{r}:=\bigcup_{\mathbf{x}\in K}B_{r]}^{(N)}(\mathbf{x}).$
Obviously, $K_{N}^{r}$ is non-increasing in $N$, and $K^{r}\subseteq K_{N}^{r}$. So, $K^{r}\subseteq G:=\bigcap_{N=1}^{\infty}K_{N}^{r}$.
On the other hand, let $\mathbf{y}\in G$. It holds that for each $N$, there is $\mathbf{x}^{(N)}\in K$ such that $\sum_{i=1}^{N}d(x_{i}^{(N)},y_{i})^{2}\leq r^{2}$. Since $K$ is compact in the product topology, there is a subsequence $\mathbf{x}^{(N_{k})},k\in\mathbb{N}$ which converges in the product topology to some point $\mathbf{z}\in K$. By the definition of product topology, for each $i$, $d(x_{i}^{(N_{k})},y_{i})\to d(z_{i},y_{i})$ as $k\to\infty$. Given any $m\in\mathbb{N}$, for sufficiently large $k$ such that $N_{k}\geq m$, by triangle inequality and Minkowski inequality, it holds that
$\sum_{i=1}^{m}d(z_{i},y_{i})^{2}\leq\sum_{i=1}^{m}(d(x_{i}^{(N_{k})},y_{i})+d(x_{i}^{(N_{k})},z_{i}))^{2}$
$\qquad\leq\Big(\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},y_{i})^{2}}+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}}\Big)^{2}\leq\Big(r+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}}\Big)^{2},$
Taking limits for both sides as $k\to\infty$, we obtain that $\sum_{i=1}^{m}d(z_{i},y_{i})^{2}\leq r^{2}.$ Letting $m\to\infty$ yields $\sum_{i=1}^{\infty}d(z_{i},y_{i})^{2}\leq r^{2}.$ That is, $\mathbf{y}\in K^{r}$. So, $G\subseteq K^{r}$. Hence, $K^{r}=G=\bigcap_{N=1}^{\infty}K_{N}^{r}$.
By a similar convergence argument, the set $K_{N}^{r}$ is closed in the product topology. Note that $K^{r}=\bigcap_{N=1}^{\infty}K_{N}^{r}$ an intersection of closed sets, hence $K^{r}$ is closed in the product topology. Since $K^{r}$ is closed in the compact set $\prod_{i=1}^{\infty}C_{i}^{r}$, it follows that $K^{r}$ is compact. ∎
We now turn back to prove $I_{\infty}(a)\geq I_{\inf}(a)$. Let $A\subseteq\mathcal{X}^{\mathbb{N}}$ be a compact set. Let $A_{n}=\pi_{n}^{-1}(\pi_{n}(A))$. Then, by the claim above, $A=\lim_{n\to\infty}A_{n}$, and $A^{r}$ is compact as well. By the claim above again, we have $A^{r}=\lim_{n\to\infty}B_{n}$, where $B_{n}=\pi_{n}^{-1}(\pi_{n}(A^{r}))=\pi_{n}^{-1}(\pi_{n}(A)^{r})=A_{n}^{r}$. So, by the continuity of probability measures,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $=\lim_{n\to\infty}\frac{\nu_{\infty}(A_{n}^{r})-\nu_{\infty}(A_{n})}{r}=\lim_{n\to\infty}\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r},$
where $E_{n}:=\pi_{n}(A)$.
Since $f(t):=\nu_{n}(E_{n}^{r})$ is monotone, it is differentiable almost everywhere on $(0,r)$. Hence, by the mean value theorem, there is some $t_{n}\in(0,r)$ such that $f$ is differentiable at $t_{n}$, and moreover, $\frac{f(r)-f(0)}{r}\geq f^{\prime}(t_{n}),$ i.e., $\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r}\geq\nu_{n}^{+}(E_{n}^{t_{n}}).$ Hence,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $\geq\limsup_{n\to\infty}\nu_{n}^{+}(E_{n}^{t_{n}})\geq\limsup_{n\to\infty}I_{n}(\nu_{n}(E_{n}^{t_{n}}))\geq\limsup_{n\to\infty}I_{\inf}(\nu_{n}(E_{n}^{t_{n}})).$
Observe that $\nu_{\infty}(A_{n})=\nu_{n}(E_{n})\leq\nu_{n}(E_{n}^{t_{n}})\leq\nu_{n}(E_{n}^{r})=\nu_{\infty}(B_{n}).$ Taking $n\to\infty$, $\nu_{\infty}(A)\leq\limsup_{n\to\infty}\nu_{n}(E_{n}^{t_{n}})\leq\nu_{\infty}(A^{r}).$ Therefore,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $\geq\inf_{\nu_{\infty}(A)\leq a\leq\nu_{\infty}(A^{r})}I_{\inf}(a).$
Taking $r\downarrow 0$ and by the continuity of $I_{\inf}$ at $a$, $\nu_{\infty}^{+}(A)\geq I_{\inf}(a).$ So, Theorem 18 holds for compact sets.
We now extend the theorem to closed sets. Let $A\subseteq\mathcal{X}^{\mathbb{N}}$ be a closed set. Since $\mathcal{X}^{\mathbb{N}}$ equipped with the product topology is Polish, it follows that $\nu_{\infty}$ is tight. That is, given any $\epsilon>0$, there is a compact set $K_{\epsilon}\subseteq\mathcal{X}^{\mathbb{N}}$ such that $\nu_{\infty}(K_{\epsilon}^{c})\leq\epsilon$. Let $A_{\epsilon}=A\cap K_{\epsilon}$ which is compact. So, $\nu_{\infty}(A_{\epsilon})\leq\nu_{\infty}(A)\leq\nu_{\infty}(A_{\epsilon})+\epsilon$ and $\nu_{\infty}(A_{\epsilon}^{r})\leq\nu_{\infty}(A^{r})\leq\nu_{\infty}(A_{\epsilon}^{r})+\epsilon$. So,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}\geq\frac{\nu_{\infty}(A_{\epsilon}^{r})-\nu_{\infty}(A_{\epsilon})-\epsilon}{r}$
$\qquad\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A_{\epsilon})\leq a\leq\nu_{\infty}(A_{\epsilon}^{r})}I_{\inf}(a)\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A)-\epsilon\leq a\leq\nu_{\infty}(A^{r})}I_{\inf}(a).$
Letting first $\epsilon\to 0$ and then $r\to 0$, by the continuity of $I_{\inf}$ at $a$, we obtain $\nu_{\infty}^{+}(A)\geq I_{\inf}(a).$ Hence, $I_{\infty}(a)\geq I_{\inf}(a)$.
Appendix C: Proof of Theorem 5
By an approximation argument of Otto and Villani [54, Appendix A], without affecting the optimal constant in (1.18), it suffices to consider $f\in W^{1,2}(\mathcal{M})$ satisfying the smoothness and boundedness conditions in (2.4). Without loss of generality, we can further assume $\mathbb{E}_{\nu}\left[f\right]=0$ and $\mathbb{E}_{\nu}\left[f^{2}\right]=1$, since otherwise we can do normalization on $f$ which neither affects the optimal constant in (1.18). Among all this kind of functions, in the following we let $f$ be the approximately optimal function attaining the optimal constant in (1.18), i.e., $\mathbb{E}[|\nabla f|^{2}]\leq(K_{P}+\delta)\cdot\mathrm{Var}_{\nu}(f)=K_{P}+\delta$ for $\delta>0$.
For such $f$, we consider the set $A_{t}=H_{t\sqrt{n}}(f)=\left\{\mathbf{x}:g_{n}(\mathbf{x})\leq t\right\}$, where $g_{n}(\mathbf{x})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f(x_{i})$. We next determine the asymptotics of the volume and the perimeter of this set.
By the coarea formula,
(6.21) $\int_{|g_{n}-t|\leq\delta}|\nabla g_{n}|\,d\nu_{n}=\int_{t-\delta}^{t+\delta}\left(\int_{g_{n}=s}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$
where $\varphi_{n}(\mathbf{x})=e^{-\sum_{i=1}^{n}V(x_{i})}$ is the density of $\nu_{n}$, ${\text{\rm d}}\mathrm{\mathrm{area}}$ is the induced volume form on the level set $\{g_{n}=s\}$, and $|\nabla g_{n}(\mathbf{x})|$ be the norm of the gradient of $g_{n}$ at the point $\mathbf{x}\in\mathcal{M}^{n}$. Thus, by the mean value theorem, the RHS above is larger than or equal to
(6.22) $2\delta\cdot\int_{g_{n}=t_{n}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\geq 2\delta\cdot\nu_{n}^{+}(A_{t_{n}}),$
for some $t_{n}\in[t-\delta,t+\delta]$. Denoting $\mathbf{X}\sim\nu_{n}$, the LHS above is equal to
(6.23) $\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}\times\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right].$
By the central limit theorem, as $n\to\infty$,
(6.24) $\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}\to\Phi(t+\delta)-\Phi(t-\delta).$
We next determine the limit of $\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right]$. Note that $|\nabla g_{n}(\mathbf{X})|^{2}=\frac{1}{n}\sum_{i=1}^{n}|\nabla f|^{2}(X_{i})$. Let $\beta:=\mathbb{E}[|\nabla f|^{2}]\leq K_{P}+\delta.$ By the weak law of large number, $\lim_{n\to\infty}\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\}\to 1.$ By basic probabilistic inequalities, for $\delta>0$,
$\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta,|g_{n}(\mathbf{X})-t|\leq\delta\}$ $\geq\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}-\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|>\delta\},$
which implies
$\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\big||g_{n}(\mathbf{X})-t|\leq\delta\}$ $\geq 1-\frac{\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|>\delta\}}{\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}}.$
By taking limit as $n\to\infty$, $\lim_{n\to\infty}\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\big||g_{n}(\mathbf{X})-t|\leq\delta\}=1.$ Since $|\nabla f|^{2}$ is bounded by assumption (and thus, so is $|\nabla g_{n}|^{2}$), we obtain
(6.25) $\lim_{n\to\infty}\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right]$ $\leq\sqrt{\beta+\delta}\leq\sqrt{K_{P}+\delta}+\delta.$
Combining (6.21), (6.22), (6.23), (6.24), and (6.25) yields
$\limsup_{n\to\infty}\nu_{n}^{+}(A_{t_{n}})\leq\frac{\Phi(t+\delta)-\Phi(t-\delta)}{2\delta}\left(\sqrt{K_{P}+\delta}+\delta\right).$
Letting $\delta\to 0$ yields
(6.26) $\limsup_{\delta\to 0}\limsup_{n\to\infty}\nu_{n}^{+}(A_{t_{n}})\leq\sqrt{K_{P}}\phi(t).$
By the central limit theorem, $g_{n}(\mathbf{X})$ converges in distribution to $Z\sim\mathcal{N}(0,1)$. Thus, the volume $\nu_{n}(A_{t_{n}})$ is asymptotically sandwiched between $\Phi(t-\delta)$ and $\Phi(t+\delta)$ as $n\to\infty$, which implies
(6.27) $\lim_{\delta\to 0}\lim_{n\to\infty}\nu_{n}(A_{t_{n}})=\Phi(t).$
Equations (6.26) and (6.27) implies that there is a sequence of sets $B_{n}\subseteq\mathcal{M}^{n}$ such that $\lim_{n\to\infty}\nu_{n}^{+}(B_{n})\leq\sqrt{K_{P}}\phi(t)$ and $\lim_{n\to\infty}\nu_{n}(B_{n})=\Phi(t)$. By continuity of $I_{\inf}$, it holds that $I_{\inf}(a)\leq\sqrt{K_{P}}\cdot I_{\mathrm{G}}(a)$.
Appendix D: Proof of Proposition 2
Proof of $I_{\nu_{1}\otimes\nu_{2}}(a)\geq J_{1}(a)$: The isoperimetric profile in our setup can be equivalently defined as the infimum of the perimeters of the sets with a given volume and smooth boundaries [10, pp. 32-33]. So, it suffices to consider an arbitrary set $A$ of measure $a$ with smooth boundary. Let $\phi_{n}:[0,\infty)\to[0,1]$ be a smooth function such that $\phi_{n}(0)=0$ and $\phi_{n}(t)=1$ for all $t\geq 1/n$. Let $f_{n}(\mathbf{x})=\phi_{n}(d(\mathbf{x},A))$, where $d$ denotes the geodesic distance on $\mathcal{M}_{1}\times\mathcal{M}_{2}$. Then, $f_{n}$ is smooth. Moreover, by the coarea formula, $\lim_{n\to\infty}\mathbb{E}|\nabla f_{n}|=(\nu_{1}\otimes\nu_{2})^{+}(A).$ Hence, it suffices to prove that $\lim_{n\to\infty}\mathbb{E}|\nabla f_{n}|\geq J_{1}(a).$
Denote $a_{n}:=\mathbb{E}[f_{n}]$ which converges $a$ as $n\to\infty$. We can also estimate $\mathbb{E}|\nabla f_{n}|$ in the following way:
$\quad\mathbb{E}|\nabla f_{n}|=\int\int\sqrt{|\nabla_{x}f_{n}|^{2}+|\nabla_{y}f_{n}|^{2}}{\text{\rm d}}\nu_{2}{\text{\rm d}}\nu_{1}$
(6.28) $\geq\int\sqrt{\left(\int|\nabla_{x}f_{n}(x,y)|{\text{\rm d}}\nu_{2}(y)\right)^{2}+\left(\int|\nabla_{y}f_{n}(x,y)|{\text{\rm d}}\nu_{2}(y)\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
(6.29) $=\int\sqrt{\Big|\nabla_{x}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+\left(\int_{0}^{1}\nu_{2}^{+}\{y:f_{n}(x,y)>t\}{\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
$\geq\int\sqrt{\Big|\nabla_{x}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+\left(\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
$=\frac{a_{n}}{a}\int\sqrt{\Big|\nabla_{x}\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+\left(\frac{a}{a_{n}}\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
(6.30) $\geq\frac{a_{n}}{a}\int\sqrt{\Big|\nabla_{x}\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))^{2}}{\text{\rm d}}\nu_{1}(x)$
$\qquad-\frac{a_{n}}{a}\int\Big|I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))-\frac{a}{a_{n}}\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\Big|{\text{\rm d}}\nu_{1}(x)$
(6.31) $\geq\frac{a_{n}}{a}J_{1}(a)-\int\Big|\frac{a_{n}}{a}I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))-\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\Big|{\text{\rm d}}\nu_{1}(x),$
where (6.28) follows by the convexity of $(x,y)\mapsto\sqrt{x^{2}+y^{2}}$, (6.29) follows by the coarea formula (noting that $\nu_{2}$ is absolutely continuous w.r.t. Riemannian volume and $f_{n}$ is smooth), (6.30) follows by the triangle inequality $|(x,y)|\geq|(x,z)|-|(0,y-z)|$, and (6.31) follows by the definition of $J_{1}$ in (1.20) and noting that $\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)$ is a feasible solution to the infimization in (1.20).
By the dominated convergence theorem and the continuity of probability measures, $\lim_{n\to\infty}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)=\nu_{2}(A_{x}),$ and
$\lim_{n\to\infty}\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t=\int_{0}^{1}\lim_{n\to\infty}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t$
$\qquad=\int_{0}^{1}I_{\nu_{2}}(\lim_{n\to\infty}\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t=\int_{0}^{1}I_{\nu_{2}}(\nu_{2}(A_{x})){\text{\rm d}}t=I_{\nu_{2}}(\nu_{2}(A_{x})),$
where $A_{x}:=\{y:\exists x\textrm{ s.t. }(x,y)\in A\}$.
Thus, letting $n\to\infty$, we get from (6.31) that $\lim_{n\to\infty}\mathbb{E}|\nabla f_{n}|\geq J_{1}(a)$, which implies $I_{\nu_{1}\otimes\nu_{2}}(a)\geq J_{1}(a)$.
Proof of $J_{1}(a)=I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)$: By considering $\nu_{2}$ as symmetric log-concave probability measures, i.e., $\nu_{2}=\tilde{\nu}_{2}$, we obtain $I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)\geq J_{1}(a)$.
We next prove $I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)\leq J_{1}(a)$. Let $\sigma:\mathcal{M}_{1}\to[0,1]$ be a feasible solution to $J_{1}(a)$, i.e., it is smooth and $\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)=a$. Let $g(x):=F_{2}^{-1}\circ\sigma(x),$ and
$A:=\{(x,y)\in\mathcal{M}_{1}\times\mathbb{R}:y\leq g(x)=F_{2}^{-1}\circ\sigma(x)\}.$
By definition,
$a=\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)$ $=\int F_{2}(g(x)){\text{\rm d}}\nu_{1}(x)=\nu_{1}\otimes\tilde{\nu}_{2}(A).$
Noting that $\nabla\sigma(x)=f_{2}(g(x))\nabla g(x)$, we have that
$\int_{\mathcal{M}_{1}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+|\nabla\sigma(x)|^{2}}{\text{\rm d}}\nu_{1}(x)=\int_{\mathcal{M}_{1}}\sqrt{f_{2}(g(x))^{2}+f_{2}(g(x))^{2}|\nabla g(x)|^{2}}{\text{\rm d}}\nu_{1}(x)$
$\qquad=\int_{\mathcal{M}_{1}}\sqrt{1+|\nabla g(x)|^{2}}f_{2}(g(x)){\text{\rm d}}\nu_{1}(x)=\left(\nu_{1}\otimes\tilde{\nu}_{2}\right)^{+}(A)\geq I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a),$
where the inequality follows by definition of the isoperimetric profile $I_{\nu_{1}\otimes\tilde{\nu}_{2}}$. Therefore, $J_{1}(a)\geq I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)$.
Proof of $I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)\geq I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)=J_{2}(a)=J_{3}(a)$: Currently we have obtain $I_{\nu_{1}\otimes\nu_{2}}(a)\geq I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)$ and by symmetry, $I_{\nu_{1}\otimes\nu_{2}}(a)\geq I_{\tilde{\nu}_{1}\otimes\nu_{2}}(a)$. Substituting $\nu_{i}$ with $\tilde{\nu}_{i}$ yields that both $I_{\tilde{\nu}_{1}\otimes\nu_{2}}(a)$ and $I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)$ are $\geq I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)$. Moreover,
$I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)=\inf_{\sigma:\int_{\mathbb{R}}\sigma(x){\text{\rm d}}F_{1}(x)=a}\int_{\mathbb{R}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+(\sigma^{\prime}(x))^{2}}{\text{\rm d}}F_{1}(x).$
Let $\varphi:=\sigma\circ F_{1}^{-1}$. By definition,
$a=\int_{\mathbb{R}}\sigma(x){\text{\rm d}}F_{1}(x)$ $=\int_{0}^{1}\sigma\circ F_{1}^{-1}(t){\text{\rm d}}t=\int_{0}^{1}\varphi(t){\text{\rm d}}t.$
Noting that $\sigma^{\prime}(x)=\varphi^{\prime}(F_{1}(x))f_{1}(x)$, we have that
$\int_{\mathbb{R}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+(\sigma^{\prime}(x))^{2}}{\text{\rm d}}F_{1}(x)$ $=\int_{\mathbb{R}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+(\varphi^{\prime}(F_{1}(x))f_{1}(x))^{2}}{\text{\rm d}}F_{1}(x)$
$=\int_{0}^{1}\sqrt{I_{\nu_{2}}^{2}(\varphi(t))+(\varphi^{\prime}(t)I_{\nu_{1}}(x))^{2}}{\text{\rm d}}t.$
Therefore, $I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)=J_{2}(a)$.
The identity $J_{2}(a)=J_{3}(a)$ can be proven similarly by substitution $g(x)=F_{2}^{-1}(\varphi(F_{1}(x))).$
Appendix E: Proof of Proposition 3
In this section, we now prove Proposition 3. By the induction method, (1.19) implies the $k$-dimensional version.
Proposition 4.
Assume that Conjecture 2 holds. Then, for any probability measures $\nu_{1},...,\nu_{k}$ satisfying Assumption 1, it holds that
(6.32) $\left(\frac{I_{\nu_{1}\otimes\nu_{2}\otimes...\otimes\nu_{k}}(a)}{a}\right)^{2}\geq\inf_{a_{1}a_{2}...a_{k}=a}\left(\frac{I_{\nu_{1}}(a_{1})}{a_{1}}\right)^{2}+\left(\frac{I_{\nu_{2}}(a_{2})}{a_{2}}\right)^{2}+...+\left(\frac{I_{\nu_{k}}(a_{k})}{a_{k}}\right)^{2},\;\forall a.$
Let $A$ be an isoperimetric minimizer of probability $e^{-n\alpha}$ in the standard sense. The proof of Theorem 10 up to (3.29) still works for this case. Recall $\tilde{A}$ defined in (3.29), whose probability satisfies $e^{-n\alpha^{\prime}}\leq\nu_{n}(\tilde{A})\leq e^{-n\alpha}$ with $\alpha^{\prime}:=\alpha+\delta^{\prime}$. Here, $\delta^{\prime}=\delta+H(\epsilon^{\prime})$ can be chosen arbitrarily small. Note that here $\nu_{n}:=\nu^{\otimes n}$.
Observe that
(6.33) $\frac{\nu_{n}^{+}(A)}{\nu_{n}(A)\sqrt{n}}\geq\frac{1}{n\sqrt{\tau}}\ln\frac{\nu_{n}(A^{\sqrt{n\tau}})}{\nu_{n}(A)}\geq\frac{1}{n\sqrt{\tau}}\ln\frac{\nu_{n}(\tilde{A}^{\sqrt{n\tau}})}{\nu_{n}(\tilde{A})}-\frac{\delta^{\prime}}{\sqrt{\tau}}\geq\frac{\nu_{n}^{+}(\tilde{A}^{r_{n,\tau}})}{\nu_{n}(\tilde{A}^{r_{n,\tau}})\sqrt{n}}-\frac{\delta^{\prime}}{\sqrt{\tau}},$
where the last inequality follows by the mean value theorem, and $r_{n,\tau}$ is some number such that $0\leq r_{n,\tau}\leq\sqrt{n\tau}$.
Denoting $a_{n,\tau}=\nu_{n}(\tilde{A}^{n,\tau})$ and $\alpha_{n,\tau}=-\frac{1}{n}\ln a_{n,\tau}$ and recalling that $n^{\prime}=\left\lceil n(1-\epsilon^{\prime})\right\rceil$ and $n^{\prime\prime}=n-n^{\prime}$, by Proposition 4, it holds that
$\left(\frac{\nu_{n}^{+}(\tilde{A}^{r_{n,\tau}})}{\nu_{n}(\tilde{A}^{r_{n,\tau}})}\right)^{2}$ $\geq\inf_{a_{0}a_{1}...a_{n^{\prime\prime}}=a_{n,\tau}}\left(\frac{I_{n^{\prime}}(a_{0})}{a_{0}}\right)^{2}+\left(\frac{I_{1}(a_{1})}{a_{1}}\right)^{2}+...+\left(\frac{I_{1}(a_{n^{\prime\prime}})}{a_{n^{\prime\prime}}}\right)^{2}$
$=\inf_{a_{0}a_{1}^{n^{\prime\prime}}=a_{n,\tau}}\left(\frac{I_{n^{\prime}}(a_{0})}{a_{0}}\right)^{2}+n^{\prime\prime}\left(\frac{I_{1}(a_{1})}{a_{1}}\right)^{2}$
$=\inf_{0\leq\alpha_{0}\leq\alpha_{n,\tau}}\left(\frac{I_{n^{\prime}}(e^{-n\alpha_{0}})}{e^{-n\alpha_{0}}}\right)^{2}+n^{\prime\prime}\left(\frac{I_{1}(e^{-n(\alpha_{n,\tau}-\alpha_{0})/n^{\prime\prime}})}{e^{-n(\alpha_{n,\tau}-\alpha_{0})/n^{\prime\prime}}}\right)^{2}.$
Note that $\frac{I_{n^{\prime}}(e^{-n\alpha_{0}})}{e^{-n\alpha_{0}}}\geq(1+o(1))\sqrt{n^{\prime}\breve{\Theta}(\alpha_{0})}$. Hence, we obtain that
$\left(\frac{\nu_{n}^{+}(\tilde{A}^{r_{n,\tau}})}{\nu_{n}(\tilde{A}^{r_{n,\tau}})\sqrt{n}}\right)^{2}$
$\geq(1+o(1))\inf_{0\leq\alpha_{0}\leq\alpha_{n,\tau}}\left\{\breve{\Theta}(\alpha_{0})+(\alpha_{n,\tau}-\alpha_{0})\left(\frac{I_{1}(e^{-n(\alpha_{n,\tau}-\alpha_{0})/n^{\prime\prime}})}{e^{-n(\alpha_{n,\tau}-\alpha_{0})/n^{\prime\prime}}\sqrt{n(\alpha_{n,\tau}-\alpha_{0})/n^{\prime\prime}}}\right)^{2}\right\}$
$\to\inf_{0\leq\alpha_{0}\leq\alpha_{\tau}}\breve{\Theta}(\alpha_{0})+2K_{1}(\alpha_{\tau}-\alpha_{0})\,=\breve{\Theta}(\alpha_{\tau}),$
where $\alpha_{\tau}:=\liminf_{n\to\infty}\alpha_{n,\tau}$, $K_{1}:=K_{IS,1}^{+}=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}$, and the equality follows since from Statement 3 of Theorem 2, it is known that $K_{1}\geq K_{IS}^{+}\geq K_{LS}^{+}\geq\frac{\breve{\Theta}^{\prime}(\alpha_{\tau})}{2}$.
Substituting these into (6.33) and taking limits as $n\to\infty$, we obtain that
(6.34) $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}(\alpha_{\tau})}-\frac{\delta^{\prime}}{\sqrt{\tau}}.$
We now estimate $\alpha_{\tau}$. By (6.33),
$\frac{1}{n}\ln\frac{\nu_{n}(\tilde{A}^{\sqrt{n\tau}})}{\nu_{n}(\tilde{A})}\leq\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}\sqrt{\tau}+\delta^{\prime}.$
Taking limits as $n\to\infty$ and applying the upper bound in (1.15), we obtain that
$\limsup_{n\to\infty}\frac{1}{n}\ln\frac{\nu_{n}(\tilde{A}^{\sqrt{n\tau}})}{\nu_{n}(\tilde{A})}\leq\sqrt{\tau\breve{\Theta}(\alpha)}+\delta^{\prime}.$
That is, $\alpha_{\tau}\geq\alpha-(\sqrt{\tau\breve{\Theta}(\alpha)}+\delta^{\prime}).$ Substituting this into (6.34) yields that $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}\left(\alpha-(\sqrt{\tau\breve{\Theta}(\alpha)}+\delta^{\prime})\right)}-\frac{\delta^{\prime}}{\sqrt{\tau}}.$ Letting $\delta^{\prime}\to 0$ first and $\tau\to 0$ then, $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$
Acknowledgements
This work was supported by the National Key Research and Development Program of China under grant 2023YFA1009604, the NSFC under grant 62101286, and the Fundamental Research Funds for the Central Universities of China (Nankai University) under grant 054-63253112.