### v4

# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

Abstract.
The isoperimetric problem is a classic topic in geometric measure theory, yet critical questions regarding the characterization of optimal solutions—even asymptotically optimal ones—remain largely unresolved. In this paper, we investigate the large deviations asymptotics (governing sequences of exponentially small sets) for the isoperimetric problem on the product Riemannian manifold $\mathcal{M}^{n}$ endowed with the product probability measure $\nu^{\otimes n}$, where $(\mathcal{M},\nu)$ is a weighted Riemannian manifold with nonnegative Bakry–Émery–Ricci curvature. We establish an exact characterization of the large deviations asymptotics of the isoperimetric profile, which reveals a precise equivalence between this asymptotic isoperimetric inequality and an established strengthening of log-Sobolev inequality—namely, the nonlinear log-Sobolev inequality. It is observed that conditional typical sets (i.e., $\mathbf{f}$-coordinate half-spaces for some vector $\mathbf{f}$ of functions), a fundamental concept from information theory, form an asymptotically optimal solution to the isoperimetric problem. This class of subsets further yields an upper bound on the isoperimetric profile in the central limit regime (concerning sets of constant weighted volume). Although our results are stated for product spaces, they imply certain isoperimetric inequalities for non-product spaces, e.g., they can be used to recover the weaker equivalence established by Ledoux and Bobkov for arbitrary non-product spaces or used to establish quantitative relations among the optimal constants in isoperimetric, concentration and transport inequalities for product or non-product spaces. Our results provide a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers behave fundamentally differently across spaces with distinct geometric structures. Our proof idea is a new framework which integrates tools from information theory, optimal transport, and geometric measure theory.
Key words and phrases:
Isoperimetric inequality, large deviations, log-Sobolev inequality, information theory, optimal transport, Wasserstein gradient flow, geometric measure theory
L. Yu is with the School of Statistics and Data Science, LPMC, KLMDASR, and LEBPS, Nankai University, Tianjin 300071, China (e-mail: leiyu@nankai.edu.cn).
1. Introduction
In a metric-measure space, the isoperimetric problem seeks the subset of fixed measure that minimizes surface area—where surface area is rigorously defined via the boundary measure. This classic problem remains largely unresolved in both its exact and asymptotic formulations, with the explicit characterization of the isoperimetric profile (the minimal surface area corresponding to each fixed measure) and its asymptotic behavior posing a widely open question, save for a handful of specific cases: e.g., the standard Gaussian measure, log-concave measures on the real line, and Euclidean spheres endowed with constant density. In this paper, we study the isoperimetric problem across various asymptotic regimes, with particular emphasis on the large deviations regime that considers sequences of exponentially small sets. We formalize our core analytical setup as follows.
Let $(\mathcal{M},g)$ be a $k$-dimensional smooth complete oriented connected Riemannian manifold with induced geodesic distance $d$. We assume $k\geq 1$ and $(\mathcal{M},g)=(\mathbb{R},|\cdot|)$ if $k=1$. Let $\nu=e^{-V}\mathrm{vol}$ with $V\in C^{2}(\mathcal{M})$ be a reference Borel probability measure on $\mathcal{M}$ that is absolutely continuous with respect to the Riemannian volume $\mathrm{vol}$. We consider the $n$-fold product space $\mathcal{M}^{n}$, equipped with the product measure $\nu_{n}:=\nu^{\otimes n}$ and the product distance $d_{n}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$, where $\mathbf{x}=(x_{1},x_{2},...,x_{n})$.
For a set $A\subseteq\mathcal{M}^{n}$, denote its $r$-enlargement as
(1.1) $A^{r}:=\bigcup_{\mathbf{x}\in A}\{\mathbf{y}\in\mathcal{M}^{n}:d_{n}(\mathbf{x},\mathbf{y})\leq r\}.$
For a Borel set $A\subseteq\mathcal{M}^{n}$, its boundary measure (i.e., perimeter) is defined by
(1.2) $\nu_{n}^{+}(A):=\liminf_{r\downarrow 0}\frac{\nu_{n}(A^{r})-\nu_{n}(A)}{r}.$
In the isoperimetric problem, the objective is to minimize $\nu_{n}^{+}(A)$ over all sets $A$ of a given measure, i.e., to determine the isoperimetric profile: For $a\in[0,1]$,
(1.3) $I_{n}(a):=\inf_{A:\nu_{n}(A)=a}\nu_{n}^{+}(A),$
where $A$ is a Borel set. Note that there are several other ways in the literature to define the perimeter of a set, but they coincide when the set admits smooth boundary, and thus, lead to the same isoperimetric profile in our setup; see details in [15, 5, pp. 32-33].
Exactly characterizing the isoperimetric profile is in general extremely difficult, and remains widely open except for certain specific cases. For example, when $\nu$ is the standard Gaussian measure, the isoperimetric profile $I_{n}(a)=I_{\mathrm{G}}(a):=\phi(\Phi^{-1}(a))$ for all $n\geq 1$ with half-spaces as isoperimetric minimizers [68, 24]. Here, $\phi$ and $\Phi$ are respectively the pdf and cdf of the standard Gaussian. It is widely known that $I_{\mathrm{G}}(a)\sim a\sqrt{2\ln\frac{1}{a}}$ as $a\to 0$. That is, the asymptotic isoperimetric ratio for the Gaussian measure is
(1.4) $\lim_{a\to 0}\frac{I_{\mathrm{G}}(a)}{a\sqrt{\ln\frac{1}{a}}}=\sqrt{2}.$
Another example is log-concave measures on the real line (i.e., $n=1$), for which the isoperimetric profile is attained by intervals or half-lines [17]. Furthermore, in the Euclidean sphere of a given dimension equipped with constant density, the isoperimetric profile is known as well, which is attained by spherical caps.
As noted at the outset of this paper, our primary aim is to study the asymptotic behavior of the isoperimetric profile, in particular its large deviations asymptotics. We focus on this behavior in the regime where the probability measure of the set decays exponentially with increasing dimension, and to this end we identify the following limits: for $\alpha>0$,
(1.5) $\underline{\Lambda}(\alpha)$ $:=\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}},$
and $\overline{\Lambda}(\alpha)$ which is defined similarly but with $\liminf_{n\to\infty}$ replaced by $\limsup_{n\to\infty}$. We denote them by $\Lambda(\alpha)$ if they are equal.
1.1. Main Results
We next provide a detailed description of our main results.
1.1.1. Large Deviations Principle
A natural convexity condition on a manifold with density is the curvature-dimension condition $\mathrm{CD}(0,\infty)$.
Condition 1 (Nonnegative Bakry–Émery–Ricci Curvature).
We assume that
$\mathrm{Ric}_{V}:=\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq 0\;\;\text{ as 2-tensor fields},$
where $\mathrm{Ric}_{g}$ is the Ricci curvature and $\mathrm{Hess}_{g}V$ is the Hessian of $V$.
Since $I_{n}(a)/a$ is non-increasing in $a$ under Condition 1 (see e.g. [52, Proposition 3.1]), $\Lambda$ is non-decreasing.
Proposition 1.
Under Condition 1, $\Lambda$ is non-decreasing.
Let $\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$ be the set of probability measures which are absolutely continuous with respect to $\nu$ and have finite second-order moments. We also need the concept of nonlinear log-Sobolev inequality and a generalized version of half-spaces.
Definition 1.
For a convex function $\theta:[0,\infty]\to[0,\infty]$, we say a probability measure $\nu$ admits the $\theta$-nonlinear log-Sobolev inequality if
(1.6) $\theta(D(\mu\|\nu))\leq I(\mu\|\nu),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),$
where for $\mu=\rho\nu$,
$D(\mu\|\nu):=\int\rho\ln\rho{\text{\rm d}}\nu\;\textrm{ and }\;I(\mu\|\nu):=\int\frac{|\nabla\rho|^{2}}{\rho}{\text{\rm d}}\nu$
are respectively the (relative) entropy and Fisher (relative) information.
Remark 1.
The inequality in (1.6) holds for all $\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$ if and only if it holds for all $\mu\in\mathscr{P}^{\mathrm{ac}}(\mathcal{M})$ (the set of probability measures which are absolutely continuous with respect to $\nu$) [58].
The nonlinear log-Sobolev inequality, also known as the entropy-energy inequality or the entropy isoperimetric inequality, can be seen as a strengthening of the standard (linear) log-Sobolev inequality since it provides more information about the tradeoff between entropy and Fisher information. This concept was investigated in several important works in the literature, e.g., Costa and Cover [26], Bakry [9], F. Y. Wang [75], Bakry, Gentil, and Ledoux [10], as well as Polyanskiy and Samorodnitsky [59].
For a convex $\theta$, a $\theta$-nonlinear log-Sobolev inequality holds only if $\theta(0)=0$ and $\theta$ is non-decreasing. In particular, when $\theta(t)=Kt$ for some constant $K>0$, then the $\theta$-nonlinear log-Sobolev inequality reduces to the standard log-Sobolev inequality. Define
$\Theta(\alpha):=\inf_{\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu)\geq\alpha}I(\mu\|\nu).$
Then, given the probability measure $\nu$, the optimal convex function $\theta$ in the $\theta$-nonlinear log-Sobolev inequality in (1.6) is the lower convex envelope of $\Theta$, given by
$\breve{\Theta}(\alpha)$ $:=\inf_{\begin{subarray}{c}\lambda\in[0,1],\alpha_{1},\alpha_{2}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{2}\geq\alpha\end{subarray}}\lambda\Theta(\alpha_{1})+(1-\lambda)\Theta(\alpha_{2})$
(1.7) $=\inf_{\begin{subarray}{c}\lambda\in[0,1],\mu_{1},\mu_{2}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}):\\ \lambda D(\mu_{1}\|\nu)+(1-\lambda)D(\mu_{2}\|\nu)\geq\alpha\end{subarray}}\lambda I(\mu_{1}\|\nu)+(1-\lambda)I(\mu_{2}\|\nu).$
It is obvious that $\breve{\Theta}(0)=0$ and $\breve{\Theta}$ is non-decreasing and convex.
The second concept we need is a generalized version of half-spaces which incorporates several important concepts as special cases, e.g., the concept of typical sets from information theory.
Definition 2.
Given $n\in\mathbb{N}$ and $\mathbf{f}=(f_{1},f_{2},...,f_{n})$ with each $f_{i}:\mathcal{M}\to\mathbb{R}$, the $\mathbf{f}$-coordinate half-space is defined as
(1.8) $H_{t}(\mathbf{f}):=\Big\{\mathbf{x}\in\mathcal{M}^{n}:\sum_{i=1}^{n}f_{i}(x_{i})\geq t\Big\}$
for a real number $t$. In particular, when $f_{i}=\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}$ for some probability measures $\mu_{i},i\in[n]$ and $t=\sum_{i=1}^{n}D(\mu_{i}\|\nu)-n\epsilon$ for some $\epsilon>0$, the $\mathbf{f}$-coordinate half-space reduces to the conditional (relative entropy) typical set of $(\mu_{i})_{i\in[n]}$ w.r.t. $\nu$ which is given by
$A_{n,\epsilon}((\mu_{i})_{i\in[n]}\|\nu):=\Big\{\mathbf{x}\in\mathcal{M}^{n}:\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})\geq\sum_{i=1}^{n}D(\mu_{i}\|\nu)-n\epsilon\Big\}.$
Remark 2.
The concept of $\mathbf{f}$-coordinate half-space reduces to the Orlicz ball if $f_{i}$ are Young functions (and t>0-t>0). However, we do not pose any restriction on functions $f_{i}$ here (except for measurability), and thus our concept is more general and incorporates standard half-spaces as special cases.
Remark 3.
Here, $(\mu_{i})_{i\in[n]}$ can be seen as a conditional probability measure $\mu_{X|W}$ with $\mu_{X|W=i}=\mu_{i}$, $\frac{1}{n}\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})$ is known as the conditional information density of $\mu_{X|W}$ w.r.t. $\nu$ given $\mu_{W}\sim\mathrm{Unif}[n]$, and its expectation $D(\mu_{X|W}\|\nu|\mu_{W}):=\frac{1}{n}\sum_{i=1}^{n}D(\mu_{i}\|\nu)$ is known as the conditional relative entropy. That is why it is called the conditional typical set. In addition, the standard definition of the conditional relative entropy typical set is $\big\{\mathbf{x}\in\mathcal{M}^{n}:\big|\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})-\sum_{i=1}^{n}D(\mu_{i}\|\nu)\big|\leq n\epsilon\big\}$ [27, (11.206)]. Our definition here is slightly different from the standard one, and has benefit that it incorporates the isoperimetric minimizers for Gaussian and Lebesgue measures as special cases, which is explained below.
The concept of $\mathbf{f}$-coordinate half-space generalizes the standard half-spaces $\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}w_{i}x_{i}\leq t\}$ in Euclidean spaces, and the former reduces to the latter when $f_{i}(x_{i})=-w_{i}x_{i}$ and $t$ is rechosen properly, or essentially equivalently, set $\nu=\mathcal{N}(0,1),\mu_{i}=\mathcal{N}(w_{i},1)$ in the conditional typical set. It also generalizes the Euclidean balls $\left\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq t\right\}$ with $t>0$, and the former reduces to the latter when $f_{i}(x_{i})=-x_{i}^{2}$ and $t$ is rechosen properly, or essentially equivalently, set $\nu=\mathrm{Lebesgue},\mu=\mathcal{N}(0,c)$ with $c>0$ in the conditional typical set (although as an infinite measure, the Lebesgue measure is not considered in this paper). In both these two cases, $\mu_{i}$ are the extremizers in the optimal nonlinear log-Sobolev inequality for $\nu$ [9, 59]. Furthermore, it is well known that standard half-spaces and Euclidean balls are respectively isoperimetric minimizers for Gaussian measures and Lebesgue measures. This observation indicates that the isoperimetric minimizers perhaps can be identified by the extremizers in the optimal nonlinear log-Sobolev inequality.
We now formalize the above observation and establish the large deviations principle for the isoperimetric problem, a result of particular interest as it reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. We state our results by splitting into cases according to whether the following condition holds.
Condition 2 (Exponential Integrability).
For some (thus all) $x_{0}\in\mathcal{M}$,
$\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\exists\lambda\in(0,\infty).$
This condition can be partitioned into two cases:
(1.9) $\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\forall\lambda\in(0,\infty),$
and
(1.10) $\mathbb{E}_{\nu}[e^{\lambda_{1}d^{2}(X,x_{0})}]<\mathbb{E}_{\nu}[e^{\lambda_{2}d^{2}(X,x_{0})}]=\infty,\;\exists 0<\lambda_{1}<\lambda_{2}<\infty.$
Under Condition 1, Condition 2 is in fact equivalent to that $\nu$ admits a standard log-Sobolev inequality [74]. We need the following assumption when the condition in (1.10) holds.
Condition 3 (Ricci Curvature Bounded Below).
$\mathrm{Ric}\geq K$ for some $K\in\mathbb{R}$.
Theorem 1 (Large Deviations Principle for Isoperimetry).
Suppose the weighted Riemannian manifold $(\mathcal{M},d,\nu)$ satisfies Condition 1.
  1. (1)
    Suppose Condition 2 holds. Suppose Condition 3 (or Conditions 4, 5, or 6 given in Section 3.3) holds when the condition in (1.10) holds. Then, for any $\alpha\geq 0$,
    (1.11) $\Lambda(\alpha)=\sqrt{\breve{\Theta}(\alpha)}$}.
    Moreover, a sequence of conditional typical sets $A_{n,\epsilon}((\mu_{1}^{*},...,\mu_{1}^{*},\mu_{2}^{*},...,\mu_{2}^{*})\|\nu)$ with $\mu_{1}^{*},\mu_{2}^{*}$ respectively appearing $\lfloor n\lambda^{*}\rfloor$ and $n-\lfloor n\lambda^{*}\rfloor$ times asymptotically attains the limit $\Lambda(\alpha)$ as $n\to\infty$ first and $\epsilon\to 0$ then, where $(\lambda^{*},\mu_{1}^{*},\mu_{2}^{*})$ is the optimal solution in (1.7).
  2. (2)
    If Condition 2 does not hold, then, for any $\alpha\geq 0$,
    $\Lambda(\alpha)=\sqrt{\breve{\Theta}(\alpha)}=0$}.
Although the theorem above is valid for spaces that are general enough, it does not cover isoperimetric problems for log-concave probability measures defined on closed convex subsets of Euclidean spaces. To cover them, we extend the theorem above to an $\mathrm{RCD}(0,\infty)$-subspace of the Riemannian manifold $\mathcal{M}$. Here, an $\text{RCD}(0,\infty)$-space (a space satisfying Riemannian Ricci Curvature Dimension bound $\text{RCD}(0,\infty)$) is a generalization of Riemannian manifolds satisfying Condition 1 to possibly non-smooth setting, e.g., compact subsets of Euclidean spaces equipped with log-concave probability measures. From the perspective of the geometry of optimal transport, an $\text{RCD}(0,\infty)$-space is a Polish metric probability measure space such that infinitesimal Hilbertianity is satisfied (i.e., the Cheeger energy is a quadratic form), and meanwhile, the relative entropy functional $D(\cdot\|\nu)$ is not merely $K$-geodesically convex on the Wasserstein space $(\mathscr{P}_{2}(\mathcal{X}),\mathsf{W}_{2})$, but that this convexity is robust: it extends to hold along all interpolated measures induced by weighted optimal geodesic plans (or test plans). See details for the definition of the $\text{RCD}(0,\infty)$-space, or equivalently, the Riemannian Energy measure space satisfying $\text{BE}(K,\infty)$ in Section 3.2 or [7, 8].
Theorem 2.
Let $\nu$ be an absolutely continuous probability measure with density $f$ (w.r.t. Riemannian volume) on a finite-dimensional smooth complete oriented connected Riemannian manifold $\mathcal{M}$ such that $\nu$ together with its support forms a $\mathrm{RCD}(0,\infty)$-subspace of $\mathcal{M}$. Suppose that there exists a sequence of probability measures $\nu_{m}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),m\in\mathbb{N}$ with densities $f_{m}$ (w.r.t. Riemannian volume) satisfying Condition 1 and $f_{m}\geq(1-1/m)f$, as well as Condition 3 if the condition in (1.10) holds. Then, for any $\alpha\geq 0$, $\Lambda(\alpha)=\sqrt{\breve{\Theta}(\alpha)}$}.
Remark 4.
The convergence of $\{\nu_{m}\}$ here is called “convergence from above” by E. Milman [51]. In addition, another notion, “convergence from within”, was also introduced by him. Theorem 2 still holds if convergence of $\{\nu_{m}\}$ from above is replaced by convergence from within.
The equivalence $\Lambda=\sqrt{\breve{\Theta}}$ in the two theorems above can be alternatively formulated as follows: For any convex $\theta:[0,\infty)\to[0,\infty)$,
${\displaystyle\Lambda(\alpha)\geq\sqrt{\theta(\alpha)}},\;\forall\alpha\geq 0\quad\Longleftrightarrow\quad I(\mu\|\nu)\geq\theta(D(\mu\|\nu)),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}).$
It is also equivalent to that for any $\alpha>0$,
(1.12) $\lim_{n\to\infty}\frac{I_{n}(a_{n})}{I_{\mathrm{G}}(a_{n})}=\lim_{n\to\infty}\frac{I_{n}(a_{n})}{a_{n}\sqrt{2\ln\frac{1}{a_{n}}}}=\sqrt{\frac{\breve{\Theta}(\alpha)}{2\alpha}}$},
where $a_{n}=e^{-n\alpha}$ and $I_{\mathrm{G}}$ is the isoperimetric profile for the standard Gaussian measure. That is, as the weighted volume vanishes exponentially, the asymptotic behavior of the isoperimetric profile for any product probability measure resembles the one for the standard Gaussian measure, but with different factors that are determined by the exponential convergence rate of the weighted volume. One prototypical case occurs when $\Theta(\alpha)$ grows linearly in $\alpha$—for instance, $\Theta(\alpha)=2\alpha$ for Gaussian measures—implying that $I_{n}(a)$ behaves as $C\cdot a\sqrt{\ln\frac{1}{a}}$, with standard half-spaces likely serving as approximate extremizers for small weighted volumes. A second prototypical case is when $\Theta(\alpha)$ grows exponentially in $\alpha$, as illustrated in the subsequent example, which implies that $I_{n}(a)$ behaves as $C\cdot\sqrt{n}\,a^{1-1/n}$, with geodesic balls acting as approximate extremizers for small weighted volumes. Theorem 1 provides a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers of small weighted volumes behave fundamentally differently across spaces with distinct geometric structures.
Example 1 (Log-Concave Probability Measures).
All smooth log-concave probability densities on Euclidean spaces satisfy Conditions 1 and 3. Moreover, by standard regularization techniques (e.g., the Moreau-Yosida regularization and convolution with Gaussian kernels), all log-concave probability densities defined on closed convex subsets of Euclidean spaces can be approximated by smooth log-concave probability densities defined on the whole Euclidean spaces in the sense as in Theorem 2. Moreover, closed convex subsets equipped with log-concave probability densities satisfy $\mathrm{RCD}(0,\infty)$. Thus, $\Lambda=\sqrt{\breve{\Theta}}$ holds for all (smooth or non-smooth) log-concave probability measures on Euclidean spaces (supported on the whole spaces or convex subsets). The standard Gaussian measure is one of the most important examples, which admits $\breve{\Theta}(\alpha)=\Theta(\alpha)=2\alpha$ and thus, $\Lambda(\alpha)=\sqrt{2\alpha}$, consistent with the formula in (1.4).
Example 2 (Bounded Manifolds admit Exponential Growth of $\Theta$).
Let $D(t)=D(\mu_{t}\|\nu)$, $I(t)=I(\mu_{t}\|\nu)$, and $N(t)=\exp(-2D(\mu_{t}\|\nu))$ respectively be the relative entropy, Fisher information, and entropy power along the (generalized) heat flow $\partial_{t}\rho=L\rho$ with $L=\Delta-\nabla V\cdot\nabla$ and $\mu_{t}=\rho_{t}\nu$. It is easy to see that $\Theta(\alpha)$ grows exponentially as $\alpha\to\infty$, as long as $N(t)$ is concave in the time $t$ and $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$. This is because, in this case, $D(t)$ continuously decays to zero as $t\to\infty$ and $N^{\prime}(t)=2N(t)I(t)$ is nonincreasing, which implies that for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, we always have $N(0)I(0)\geq N(t_{0})I(t_{0})\geq\Theta(\alpha_{0})e^{-2\alpha_{0}}$ with $t_{0}$ being the solution to $D(t_{0})=\alpha_{0}$. Therefore, $I(0)\geq\Theta(\alpha_{0})e^{2(D(0)-\alpha_{0})}$ for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, i.e., exponential growth of $\Theta$ (when $\Theta(\alpha_{0})>0$): $\Theta(\alpha)\geq\Theta(\alpha_{0})e^{2(\alpha-\alpha_{0})}$ for all $\alpha\geq\alpha_{0}$. The concavity of the entropy power holds under the $\mathrm{CD}(0,m)$-condition with $m\in[n,\infty)$ for smooth Riemannian manifolds, as shown by S. Li and X. Li [47, Theorem 2.2], and also under the $\mathrm{RCD}(0,m)$-condition with $m\in[n,\infty)$ and $\int_{\mathcal{X}}\left[\frac{|\Delta\rho|^{2}}{\rho}+\frac{|\nabla\Delta\rho|^{2}}{\rho}\right]{\text{\rm d}}\nu<\infty$ for Polish metric measure spaces with $\rho$ denoting the heat flow, as shown by X. Li and E. Zhang [48, Theorem 3.1 and Remark 3.3]. The condition that $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$ is guaranteed as long as a standard log-Sobolev inequality (or a Poincaré inequality) is admitted. These two conditions, concavity of $N$ and positivity of $\Theta(\alpha_{0})$, are satisfied by the (normalized) unweighted measure on a bounded manifold with nonnegative Ricci curvature, which hence admits exponential growth of $\Theta$ and also exponential growth of $\Lambda$ (since Conditions 1 and 3 are satisfied in this setting and Theorem 1 is valid). See other sufficient conditions to ensure exponential growth of $\Theta$ in [10, Theorem 6.8.1][47, Theorem 2.5].
We next consider a simple but nontrivial case—the uniform measure on hypercube.
Example 3 (Hypercube).
Consider the hypercube $0,1]^{n}$ (equivalently, $(0,1)^{n}$) with Euclidean metric and uniform measure. The standard log-Sobolev inequality and the Poincaré inequality with constants $K_{LS}=K_{P}=\pi^{2}$ ($K_{LS},K_{P}$ defined later in (1.14) and (1.20)) respectively hold in this setting [32], which implies $\lim_{\alpha\to 0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}=\pi^{2}$. Moreover, the $\mathrm{RCD}(0,m)$-condition holds obviously for $m>n$ and the condition $\int_{[0,1]}\left[\frac{|\Delta\rho|^{2}}{\rho}+\frac{|\nabla\Delta\rho|^{2}}{\rho}\right]{\text{\rm d}}\nu<\infty$ is satisfied for any $t>0$ for heat flow $\rho$ with the Neumann boundary condition [67, Section 4.2] (since $[0,1]$ is compact, $\rho\in C^{\infty}([0,1])$, and $\rho_{t}>0$ for any $t>0$). Thus, as mentioned in Example 2, we have exponential growth of $\Theta$, i.e., $\Theta(\alpha)\geq 2\pi^{2}\alpha_{0}\cdot e^{2(\alpha-\alpha_{0})}\geq\frac{\pi^{2}}{e}\cdot e^{2\alpha}$ (with $\alpha_{0}$ chosen as $1/2$), which further implies exponential growth of $\Lambda$, i.e., $\Lambda(\alpha)\geq\sqrt{\frac{\pi^{2}}{e}}\cdot e^{\alpha}$. This lower bound is sharp in the exponent by verifying balls centered at the vertices of the hypercube, since for these balls with volume $e^{-n\alpha}$, their surface areas are $(1+o(1))\sqrt{\frac{\pi en}{2}}\cdot e^{\alpha(1-n)}$ as $n\to\infty$, giving $\Lambda(\alpha)\leq\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$. In fact, it was already known [63, 35] that given each $n$, for all small volumes, balls centered at the vertices are isoperimetric minimizers. However, it is unclear whether “small volume” here is large enough to include “exponentially small volume”. We conjecture the following precise relation holds: $\Lambda(\alpha)\sim\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$, or equivalently, $\breve{\Theta}(\alpha)\sim\frac{\pi e}{2}\cdot e^{2\alpha}$ as $\alpha\to\infty$, suggesting that half-normal distributions truncated on $(0,1)$ with sufficiently small variances approximately saturate nonlinear log-Sobolev inequalities for large $\alpha$. The factor $\frac{\pi^{2}}{e}$ obtained above is about $0.85\cdot\frac{\pi e}{2}$.
Theorem 1 reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. It is termed the large deviations principle for isoperimetry for two reasons. First, both this theorem and large deviations theory focus on sets of exponentially small measures. Second, isoperimetric minimizers have been proven to be expressed in terms of $\mathbf{f}$-coordinate half-spaces (or conditional typical sets) which concern the remote tails of a sum. These remote tails themselves are the core object of investigation in large deviations theory.
Showing equivalence among inequalities from various fields is an interesting topic, since it will unveil the intrinsic relationships among diverse branches of mathematics. However, most of equivalences in the literature on this topic are known only in a weak sense. In particular, it was proved by Ledoux [41, 42] on Gaussian spaces or compact Riemannian manifolds and by Bobkov [20] on Euclidean spaces with log-concave probability measures that for a fixed $n$, a Gaussian-type isoperimetric inequality
$I_{n}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a$
holds for some $K>0$, if and only if the standard log-Sobolev inequality holds with some constant $K^{\prime}>0$. Theorem 1 provides an equivalence in a stronger sense, which showing a precise equivalence between the isoperimetric inequality and (nonlinear) log-Sobolev inequality. As demonstrated in the case of bounded manifolds, Ledoux and Bobkov’s weak equivalence exhibits significant looseness for small $a$ (that is, in the context of small sets).
Although our equivalence is proven for product spaces, it in fact implies Ledoux and Bobkov’s weak equivalence for non-product spaces. In particular, by the facts that $I_{1}\geq I_{n}$ and $\mathcal{M}$ is an arbitrary (non-product) space, if the standard log-Sobolev inequality holds with some constant $K^{\prime}>0$, then our equivalence implies that $\liminf_{a\to 0}\frac{I_{1}(a)}{I_{\mathrm{G}}(a)}\geq\liminf_{n\to\infty}\frac{I_{n}(a_{n})}{I_{\mathrm{G}}(a_{n})}=\sqrt{K^{\prime}},$ where $a_{n}=e^{-n\alpha}$ for any $\alpha>0$. (More details on the inequality here is given in the following subsubsection.) That is, $I_{1}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a$ for some $K>0$, yielding a Gaussian-type isoperimetric inequality for the space $\mathcal{M}$. Conversely, if a Gaussian-type isoperimetric inequality for $\mathcal{M}$ holds with constant $K>0$, then by the well known fact that $I_{1}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a\Longleftrightarrow I_{n}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a,n$, our equivalence implies the standard log-Sobolev inequality with the same constant $K$.
Our equivalence will be applied in the following to compare the optimal constants in various inequalities.
1.1.2. Companion of Various Constants
While Theorem 1 is primarily tailored to subsets of weighted volumes decaying exponentially in $n$, it admits application to bounding the isoperimetric profile for weighted volumes decaying arbitrarily slowly or even constant weighted volumes. In other words, Theorem 1 can be used to quantitatively compare the log-Sobolev constant, the isoperimetric constant, and other constants for various inequalities. To this end, we define the following constants.
  • $K_{CD}$ denotes the optimal constant $K$ in the Bakry–Émery–Ricci curvature-dimension condition
    (1.13) $\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq Kg.$
  • $K_{LS}$ denotes the optimal constant $K$ in the log-Sobolev inequality
    (1.14) $I(\mu\|\nu)\geq 2K\cdot D(\mu\|\nu),\;\forall\mu,$
    or equivalently,
    $K_{LS}=\frac{\breve{\Theta}^{\prime}(0)}{2}=\inf_{\alpha>0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\inf_{\alpha>0}\frac{\Theta(\alpha)}{2\alpha},$
    where $\breve{\Theta}^{\prime}(0)$ is the right derivative of $\breve{\Theta}$ at 0.
  • $K_{LS}^{+}$ denotes
    (1.15) $K_{LS}^{+}:=\sup_{\alpha>0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to\infty}\frac{\breve{\Theta}(\alpha)}{2\alpha}.$
  • $K_{HC}$ denotes the optimal constant $K$ in the hypercontractivity inequality for the semigroup $T_{t}=e^{-tL}$ with $L=\Delta-\nabla V\cdot\nabla$,
    $\|T_{t}f\|_{q}\leq\|f\|_{p},\;\forall f,$
    for all $t\geq\frac{1}{2K}\ln\frac{q-1}{p-1}$, where the norms are taken w.r.t. $\nu$.
  • $K_{T}$ denotes the optimal constant $K$ in the transport inequality
    $\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),\quad\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},$
    where
    $\mathsf{W}_{2}(\mu,\nu):=\left(\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{M}\times\mathcal{M}}d^{2}(x,y){\text{\rm d}}\pi(x,y)\right)^{1/2}$
    is the Wasserstein distance of order $2$ and $\Pi(\mu,\nu)$ is the set of couplings of $\mu,\nu$, i.e., the set of joint distributions with marginals equal to $\mu,\nu$.
  • $K_{T}^{+}$ denotes the optimal constant $K$ in the transport inequality that there is some constant $D_{0}\geq 0$ such that for all
    $\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},\quad\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})\textrm{ s.t. }D(\mu\|\nu)\geq D_{0}.$
  • $K_{C}$ denotes the optimal constant $K$ in the dimension-free Gaussian concentration inequality that for all $n\geq 1$ and $r\geq 0$, $\nu_{n}((A^{r})^{c})\leq e^{-\frac{Kr^{2}}{2}}$ for all $A$ such that $\nu_{n}(A)\geq 1/2$.
  • $K_{C}^{+}$ denotes the optimal constant $K$ in the asymptotic Gaussian concentration inequality that there is some constant $r_{0}\geq 0$ such that for all $n\geq 1$ and $r\geq\sqrt{n}r_{0}$, $\nu_{n}((A^{r})^{c})\leq e^{-\frac{Kr^{2}}{2}}$ for all $A$ such that $\nu_{n}(A)\geq 1/2$.
  • $K_{IS}$ denotes the optimal constant $K$ in the large deviations asymptotics of the isoperimetric inequalities
    $\liminf_{n\to\infty}\frac{I_{n}(a_{n})}{a_{n}\sqrt{2\ln\frac{1}{a_{n}}}}\geq\sqrt{K},\;\forall\alpha>0,$
    where $a_{n}=e^{-n\alpha}$.
  • $K_{IS}^{+}$ denotes the optimal constant $K$ in the following asymptotic isoperimetric inequality
    $\liminf_{a\to 0}\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq\sqrt{K},\;\forall n\geq 1.$
  • $K_{IS}^{-}$ denotes the constant in the dimension-free (also infinite-dimensional) isoperimetric inequality
    (1.16) $\liminf_{a\to 0}\frac{I_{\inf}(a)}{a\sqrt{2\ln\frac{1}{a}}}=\sqrt{K_{IS}^{-}},$
    where
    $I_{\inf}(a):=\lim_{n\to\infty}I_{n}(a)=\inf_{n\geq 1}I_{n}(a).$
From the definitions, the constants $K_{IS}$, $K_{IS}^{+}$, and $K_{IS}^{-}$ are the asymptotic isoperimetric ratios $\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}$ respectively for exponentially, superexponentially, and subexponentially small sequences of volumes. We now compare all optimal constants defined above.
Theorem 3.
Suppose the weighted Riemannian manifold $(\mathcal{M},d,\nu)$ satisfies Condition 1 (i.e., $K_{CD}\geq 0$). Then, it holds that
(1.17) $K_{CD}\overset{\mathrm{(a)}}{\leq}\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\overset{\mathrm{(b)}}{\leq}$ $K_{IS}^{-}\overset{\mathrm{(c)}}{\leq}K_{IS}\overset{\mathrm{(d)}}{\leq}K_{LS}\overset{\mathrm{(e)}}{=}K_{HC}$
$\overset{\mathrm{(f)}}{\leq}K_{T}\overset{\mathrm{(g)}}{=}K_{C}\overset{\mathrm{(h)}}{\leq}K_{C}^{+}\overset{\mathrm{(i)}}{=}K_{T}^{+},$
and $K_{LS}\overset{\mathrm{(j)}}{\leq}K_{LS}^{+}.$ If additionally, Condition 3 holds when the condition in (1.10) holds, then the equality in (d) holds and moreover, $K_{LS}^{+}\overset{\mathrm{(k)}}{\leq}K_{IS}^{+}.$ In addition, if $K_{CD}$ is equal to any one of $K_{IS}^{-},K_{IS},K_{LS},K_{HC},K_{T},K_{C}$, then all the inequalities in (a)–(g) of (1.17) become equalities.
Inequality (a) is trivial, by noting that $0\leq K_{CD}/K_{C}\leq 1$. Inequality (c) is due to the fact that $n\mapsto I_{n}(a)$ is nonincreasing. Equality (d) is implied by Theorem 1. Equality (e) is a classic result of Gross [38]. The inequality $K_{LS}\leq K_{T}$ (or equivalently, Inequality (f)) was proven by Otto and Villani [58]. Equality (g) is also well known; see e.g. [73, Remark 22.23]. Inequalities (h) and (j) follow by definition. Equality (i) follows similarly to (g). Inequality (k) follows by Theorem 1, with a detailed proof given in Appendix C. So, it only remains to prove Inequality (b). This is done in Appendix C. In addition, a weaker version of (b), $\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\leq K_{LS}$, was already known; see Theorem 22.21 of [73]. However, we do not know if this is equivalent to the inequality (b) since we do not know if $K_{IS}^{-}=K_{IS}$ holds.
Example 4 (Hypercube).
We continue considering the hypercube described in Example 3. Statements in Example 3 imply that for this case,
$K_{CD}=0,\;K_{IS}=K_{LS}=K_{HC}=K_{T}=K_{C}=\pi^{2},\;K_{LS}^{+}=K_{C}^{+}=K_{T}^{+}=K_{IS}^{+}=\infty,$
where the density $1+\epsilon\cos\pi x$ asymptotically attains the constants $K_{LS}$ and $K_{T}$ as $\epsilon\to 0$. Moreover, it holds that $\sqrt{2\pi}\leq K_{IS}^{-}\leq\pi^{2}$; the lower bound will be explained in Example 5. However, the exact value of $K_{IS}^{-}$ is still unknown.
An open problem on this topic is the following one.
Problem 1.
Under Condition 1, is it true that $K_{IS}^{-}=K_{LS}$?
1.1.3. Central Limit Regime
We now consider the asymptotics of the isoperimetric profile for fixed weighted volume as $n\to\infty$. Let $I_{\infty}$ be the isoperimetric profile defined for the infinite dimensional product space $\mathcal{M}^{\infty}$ equipped with the product topology and the product measure $\nu_{\infty}:=\nu^{\otimes\infty}$, but the enlargement of a set is defined under $d_{\infty}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}}$ (although it might diverse to infinity for two arbitrary points). Then, it is obvious that $I_{\inf}(a)\geq I_{\infty}(a)$. In the following, we show that this is actually an equality. That is, the isoperimetric profile $I_{\infty}$ on the infinite dimensional space is the pointwise limit of the isoperimetric profiles on finite dimensional spaces. The proof is provided in Appendix D.
Theorem 4.
Assume that $I_{\inf}$ is continuous (e.g., Condition 1 holds). Then, $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in[0,1]$.
A well-known lower bound on $I_{\inf}$ is that
(1.18) $I_{\inf}(a)\geq K_{0}\cdot I_{\mathrm{G}}(a),\;\forall a\in[0,1],$
where $I_{\mathrm{G}}$ is the Gaussian isoperimetric profile and $K_{0}:=\inf_{a\in(0,1)}\frac{I_{\nu}(a)}{I_{\mathrm{G}}(a)}$ [19, 11]. An upper bound can be obtained by considering the $\mathbf{f}$-coordinate half-space given in (1.8), as shown in the following theorem. The proof is based on the coarea formula and the central limit theorem, which is given in Appendix E.
Theorem 5 (Upper Bound for Isoperimetry in Central Limit Regime).
Assume that $I_{\inf}$ is continuous (e.g., Condition 1 holds). Then, for any $a\in[0,1]$,
(1.19) $I_{\inf}(a)\leq\sqrt{K_{P}}\cdot I_{\mathrm{G}}(a),$
where $K_{P}$ is the spectral gap (the reciprocal of the Poincaré constant) for $\nu$, i.e., the optimal constant in
(1.20) $K_{P}\cdot\mathrm{Var}_{\nu}(f)\leq\mathbb{E}_{\nu}\left[\|\nabla f\|^{2}\right]$
for all $f$ in the Sobolev space $W^{1,2}(\mathcal{M})$.
As a related result, the Poincaré constant and the Cheeger constant are known to be closely related; see e.g., [40]. The Cheeger constant characterizes the isoperimetric behavior exclusively for large subsets, whereas the inequality in (1.19) captures isoperimetry for all subsets without restriction.
The upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ in (1.19) can be used to bound $K_{IS}^{-}$, namely, obtaining that $K_{IS}^{-}\leq K_{P}$. However, this bound is worse than the upper bound $K_{LS}$, since in general $K_{LS}\leq K_{P}$. Note that by perturbation analysis, one can see that
$K_{P}=\liminf_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}.$
Thus, $K_{LS}=K_{P}$ if $\Theta(\alpha)\geq 2K_{P}\alpha$ (e.g., if $\Theta$ is convex), and otherwise, $K_{LS}<K_{P}$ (e.g., for exponential measures [73, Example 21.19]). For the latter case, the upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ is not tight for sufficiently small $a$. Combining the best known lower bound in (1.18), the upper bound in (1.19), and the trivial upper bound $I_{1}$ yields
$K_{0}I_{\mathrm{G}}(a)\leq I_{\inf}(a)\leq I_{1}(a)\wedge\sqrt{K_{P}}I_{\mathrm{G}}(a).$
Example 5 (Hypercube).
We continue considering the hypercube described in Example 3. Applying the inequalities above to the hypercube $(0,1)^{n}$ yields $\sqrt{2\pi}I_{\mathrm{G}}(a)\leq I_{\inf}(a)\leq 1\wedge\pi I_{\mathrm{G}}(a)$. As shown in [35], the lower bound is not tight.
A challenging problem is to exactly determine the function $I_{\inf}$ (or equivalently, $I_{\infty}$). This problem is a commonly recognized open one.
Problem 2.
Under Condition 1, what is the explicit expression of $I_{\inf}$?
1.2. Related Works
  • Isoperimetry
In the one-dimensional setting, Bobkov [17] established that half-lines are extremal sets in the isoperimetric inequality for log-concave probability measures. Extending to higher-dimensional structures, Bobkov and Houdré [21] further explored the isoperimetric constant and log-Sobolev inequalities in product spaces, laying a foundation for multi-dimensional isoperimetric analysis. The optimality of standard half-spaces— a core topic in isoperimetry— has been systematically studied by multiple scholars: Sudakov and Tsirel’son [68] initiated early investigations, Bobkov [17] deepened this topic by addressing the isoperimetric problem for log-concave product measures on $\mathbb{R}^{n}$ equipped with the uniform distance, and explicitly provided necessary and sufficient conditions for standard half-spaces to be extremal, and Barthe [12] investigated the extremal properties of central half-spaces (with measure $1/2$) for product probability measures on Riemannian manifolds equipped with the geodesic distance. It should be noting that the isoperimetric problem under the geodesic distance and the one under the uniform distance are significantly different. When equipped with the uniform distance, the asymptotics of the isoperimetric profile has been characterized by Bobkov in [18]. However, for the geodesic distance as in our setup, the asymptotics is still widely unknown except for special cases, e.g., the Gaussian setting. Roberto [64] offers a critical survey for a comprehensive overview of isoperimetry in product spaces. See also other related works or survey in [14, 13, 65, 51]. Furthermore, the isoperimery for small sets was already investigated in the literature; see, e.g., [46] for the continuous setting and [61, 57] for the discrete setting. In the discrete setting, the isoperimery for small sets is known as the small set expansion and has important applications in complexity theory [61]. It should be noted that exact characterization of the isoperimetric profile remains widely open, and previously so does its asymptotics.
  • Concentration of Measure
The concentration of measure in a probability metric space describes a key phenomenon: a slight enlargement of any measurable set with non-negligible probability will always result in a set with large probability. In functional analysis terms, this is equivalent to the fact that the value of any Lipschitz function is concentrated around its median. The modern study of this phenomenon began in the early 1970s, when V. Milman pioneered its application to the asymptotic geometry of Banach spaces. Subsequent decades saw in-depth exploration by scholars including Gromov, Maurey, Pisier, Schechtman, Talagrand, and Ledoux, solidifying its role in analysis and probability. A critical advancement came from Talagrand [69], who studied concentration of measure in product spaces equipped with product probability measures and derived a suite of sharp concentration inequalities. In information theory, this phenomenon is known as the blowing-up lemma [1, 50], which Gács, Ahlswede, and Körner leveraged to prove strong converses for two classic coding problems.
Concentration of measure has natural connections to isoperimetric inequalities. In fact, it can be seen as a variant version of isoperimetric problem concerning on sets with probability measure $1/2$ in which the boundary of a set is set to the $r$-neighbourhood of the set with $r>0$. A landmark methodological contribution in concentration of measure was made by Marton [50], who was the first to introduce information-theoretic techniques (especially transport-entropy inequalities) to the study of concentration of measure, yielding an elegant and concise proof for a dimension-free bound on concentration of measure. Talagrand extended this idea by developing a new transport-entropy inequality, adapting it to Gaussian measures and the Euclidean metric [70]; this argument has since become a standard reference, featured in prominent textbooks [43, 62, 72]. Gozlan and Léonard further refined this approach by replacing the “linear” transport-entropy inequality in Marton’s argument with a “nonlinear” version, obtaining a dimension-free sharp bound on concentration of measure—exponentially tight, as the exponent of the bound is asymptotically attained [37]. Prior to Gozlan and Léonard’s work, Alon et al. [4] are the first to characterize the sharp convergence exponent for concentration of measure, but limited to finite spaces. Gozlan [36] also used Marton’s framework to prove the equivalence between the Gaussian bound and Talagrand’s transport-entropy inequality. Additionally, Dembo [28] proposed a new class of transport-entropy inequalities, which he used them to recover and generalize several of Talagrand’s key results in [69]. Furthermore, E. Milman [52] established an equivalence between isoperimetry and concentration of measure. There are many existing works investigating isoperimetric inequalities or concentration of measure in the discrete setting, e.g., on various graphs, e.g., [49, 1, 23, 31].
Ahlswede and Zhang [3] and Yu [77] contributed to the study of isoperimetry with thick boundaries for exponentially small sets. This isoperimetric variant refines the concentration of measure phenomenon by relaxing the set size requirement to allow arbitrarily exponentially small measures, yet it behaves fundamentally differently from—and is far more intricate than—concentration of measure. This is because, in contrast to the latter, the isoperimetric variant admits no dimension-independent bounds. Ahlswede and Zhang [3] focused on finite spaces for which a key tool called the inherently typical subset lemma was developed by them. Yu [77] generalized their result to Polish metric spaces (complete, separable metric spaces), expanding the problem’s applicability, which plays a key role in our proof in the present paper.
  • Log-Sobolev Inequalities
Ledoux [42, 44] proposed a semigroup approach to derive isoperimetric inequalities from hypercontractivity (which is equivalent to log-Sobolev inequalities). The semigroup approach can be used to establish isoperimetric inequalities for the whole range of volumes, but the constant obtained via this method is in general suboptimal. In contrast, the constant in our work is asymptotically optimal. Otto and Villani [58] established that log-Sobolev inequalities imply transport-entropy inequalities, bridging log-Sobolev theory with concentration of measure. Building on this, Ledoux [45] (alongside Milman [52]) used a semigroup approach to derive isoperimetric inequalities from concentration inequalities (equivalently, transport-entropy inequalities). As a nonlinear log-Sobolev inequality for the standard heat semigroups (with Lebesgue measures as stationary measures), the entropic isoperimetric inequality was investigated in [26, 29] which is in fact equivalent to the standard log-Sobolev inequality for Ornstein–Uhlenbeck semigroups (with Gaussian measures as stationary measures). Nonlinear log-Sobolev inequalities for general heat (diffusion) semigroups were investigated by Bakry [9] (called the entropy-energy inequality therein), F. Y. Wang [75], Bakry, Gentil, and Ledoux [10], as well as Polyanskiy and Samorodnitsky [59]. Specifically, F. Y. Wang [75] showed that nonlinear log-Sobolev inequalities imply nonlinear transport-entropy inequalities, extending Otto and Villani’s linear version to the nonlinear setting. Bakry [9], Bakry, Gentil, and Ledoux [10], and Polyanskiy and Samorodnitsky [59] applied nonlinear log-Sobolev inequalities to strengthen hypercontractive inequalities. Though nonlinear Sobolev inequalities have been studied in the literature not directly to address asymptotic isoperimetry, an exact equivalence between these two concepts does exist—one we establish in the present work. Furthermore, equivalence between isoperimetry and Sobolev-type inequalities was investigated in [22].
1.3. Proof Idea and Paper Organization
Log-Sobolev Variant Isoperimetry TransportIsoperimetry OT IT GMT IT + OT or GMT
Figure $1.$ Proof strategy of Theorem 1, where OT, IT, and GMT respectively refer to optimal transport, information theory, and geometric measure theory.
Ledoux’s semigroup method, as developed in [42, 44], seems insufficient for establishing Theorem 1. This limitation originates from the core strategy of the method: it transforms the isoperimetric problem into a “soft boundary” variant known as the noise stability problem. However, obtaining a precise solution to the former demands an exact (or sufficiently precise, not merely asymptotic) solution to the latter—an outcome that remains elusive in general. While asymptotically optimal solutions for the noise stability problem are well-established [57, 39, 78, 76], an exact resolution has yet to be achieved. Accordingly, a novel approach is required to advance the proof of Theorem 1.
The heuristic reasoning behind our proof idea is as follows. Given an arbitrary set in a high-dimensional space, due to lack of structures, its (weighted) surface area does not admit any simple formulas and thus, is difficult to analyze. To resolve this difficulty, a possible way is to tailor the set to a new one that is “regular” in a certain sense, so that the surface area of the resulting set admits a simple formula. However, this is difficult to realize in a direct way, since the surface area is very sensitive to the shape of the set, and thus any subtle change of the set could lead to drastic change in surface area.
Our proof idea of Statement 1 of Theorem 1 is as follows, which is new and totally different from the semigroup approach. Firstly, using a generalized Heintze-Karcher theorem [55] from geometric measure theory, we convert the standard isoperimetric problem into a variant version in which the “perimeter” is measured by the size of a thick shell of the set. Secondly, using information-theoretic tools—the inherently typical subset lemma [2] and its consequence [77]—as well as covering arguments, we tailor the set into a “regular” one the measure of whose thick shell admits a simple formula, which is expressed in terms of nonlinear transport-entropy inequality. The feasibility of this step is due to the fact that the size of thick shell is robust to any slight change of the set. Lastly, by establishing a nonlinear transport inequality analogous to [75] using gradient flow (heat flow) from optimal transport [73], we simplify the resulting formula and express it in terms of nonlinear log-Sobolev inequality, leading to the equivalence between the standard isoperimetric problem and the nonlinear log-Sobolev inequality.
Our specific strategy is illustrated in Fig. 1.1, which is described in the order opposite to the one given above. Specifically, using tools from optimal transport, in Section 2 we first deduce a specific form of nonlinear transport inequality from the nonlinear log-Sobolev inequality. By tools from information theory as well as covering and heat flow arguments, we show in Section 3 that the nonlinear transport inequality in turn implies a variant isoperimetric inequality—the isoperimetric inequality with thick boundaries. Lastly, by tools from geometric measure theory, we show in Section 4 that the variant isoperimetric inequality implies the original isoperimetric inequality. Putting everything together, in Section 5 we deduce that the nonlinear log-Sobolev inequality implies the isoperimetric inequality. As for the other direction, we provide two proofs in Section 6: one is based on information theory and optimal transport, and the other is based on geometric measure theory.
We prove Statement 2 of Theorem 1, Theorem 2, Theorem 3, Theorem 4, and Theorem 5 respectively in Appendices A–E.
2. From Log-Sobolev to Transport
In this section, we connect log-Sobolev inequality and transport-entropy inequality. Let $(\mathcal{X},d,\nu)$ be a Polish (complete and separable) metric probability measure space, equipped with the $\nu$-completion of the Borel $\sigma$-algebra.
We now need some metric notions; refer to [6, 33] for details. A curve $\gamma:[0,\infty)\to\mathcal{X}$ is said to be absolutely continuous if there exists $g\in L^{1}(0,\infty)$ such that
(2.1) $d(\gamma_{s},\gamma_{t})\leq\int_{s}^{t}g(r)\,{\text{\rm d}}r,\quad\forall 0\leq s<t<\infty.$
In this case, the metric speed of $\gamma$ is well-defined for a.e. $t$ by
$|\dot{\gamma}_{t}|:=\lim_{h\to 0}\frac{d(\gamma_{t+h},\gamma_{t})}{|h|}.$
It turns out that $|\dot{\gamma}|\in L^{1}(0,1)$ and this is the minimal $L^{1}$ function for which (2.1) holds.
Given a function $f:X\to\mathbb{R}$, the local Lipschitz constant $|\nabla f|:\mathcal{X}\to[0,\infty]$ is defined by
$|\nabla f|(x):=\limsup_{y\to x}\frac{|f(x)-f(y)|}{d(x,y)}.$
Define the Cheeger energy functional $\operatorname{Ch}:L^{2}(\mathcal{X},\nu)\to[0,\infty]$ by
(2.2) $\operatorname{Ch}(f):=\inf\liminf_{n\to\infty}\frac{1}{2}\int|\nabla f_{n}|^{2}\,{\text{\rm d}}\nu,$
where the infimum is taken among all sequences $(f_{n})$ of Lipschitz functions converging to $f$ in $L^{2}(\mathcal{X},\nu)$. Then the Sobolev space $W^{1,2}(\mathcal{X},d,\nu)$ is defined as the set of $f\in L^{2}(\mathcal{X},\nu)$ such that $\operatorname{Ch}(f)<\infty$ endowed with the norm
(2.3) $\|f\|_{W^{1,2}}:=\|f\|_{L^{2}}^{2}+2\operatorname{Ch}(f).$
A function $G\in L^{2}(\mathcal{X},\nu)$ is said to be a relaxed gradient of $f\in L^{2}(\mathcal{X},\nu)$ if there exist Lipschitz functions $f_{n}\in L^{2}(\mathcal{X},\nu)$ such that:
  • $f_{n}\to f$ in $L^{2}(\mathcal{X},\nu)$ and $|\nabla f_{n}|$ weakly converge to $\tilde{G}$ in $L^{2}(\mathcal{X},\nu)$;
  • $\tilde{G}\leq G$ $\nu$-a.e. in $\mathcal{X}$.
A function $G$ is said to be the minimal relaxed gradient of $f$ if its $L^{2}(\mathcal{X},\nu)$ norm is minimal among relaxed gradients. We shall denote by $|\nabla f|_{*}$ the minimal relaxed gradient. In terms of the minimal relaxed gradient, the Cheeger energy can be equivalently expressed as
$\operatorname{Ch}(f)=\frac{1}{2}\int|\nabla f|_{*}^{2}\,{\text{\rm d}}\nu.$
For a probability measure $\mu$ defined on $\mathcal{X}$ with density $f$ w.r.t. $\nu$, the Fisher information is defined as
$I(\mu\|\nu):=\int_{\{f>0\}}\frac{|\nabla f|_{*}^{2}}{f}{\text{\rm d}}\nu,$
which reduces to the definition in when the space is a smooth Riemannian manifold. Recall the definition of the relative entropy:
$D(\mu\|\nu)$ $:=\int f\ln f{\text{\rm d}}\nu.$
We first the following equivalent formulations of the nonlinear log-Sobolev inequality, which is a nonlinear analogue of the linear version given in [33, Proposition 1]. The proof is almost same as the linear version, and omitted here.
Proposition 2 (Equivalent formulations of the nonlinear log Sobolev inequality [33]).
Let $(\mathcal{X},d,\nu)$ be a Polish metric measure space and let $\theta$ be a continuous function. Then the following are equivalent:
  • For any Lipschitz function $f:\mathcal{X}\to[0,\infty)$ with $\int f\,{\text{\rm d}}\nu=1$, it holds
    $\theta(\int f\ln f\,{\text{\rm d}}\nu)\leq\int_{\{f>0\}}\frac{|\nabla f|^{2}}{f}\,{\text{\rm d}}\nu.$
  • For any non-negative $f\in W^{1,2}(\mathcal{X},d,\nu)$ with $\int f\,{\text{\rm d}}\nu=1$, it holds
    $\theta(\int f\ln f\,{\text{\rm d}}\nu)\leq\int_{\{f>0\}}\frac{|\nabla f|_{*}^{2}}{f}\,{\text{\rm d}}\nu.$
Given $f\in W^{1,2}(\mathcal{X},d,\nu)$, we write $\partial^{-}\mathrm{Ch}(f)\subset L^{2}(\mathcal{X},\nu)$ for the subdifferential at $f$ of the restriction to $L^{2}(\mathcal{X},\nu)$ of Cheeger’s energy, namely, $\xi\in\partial^{-}\mathrm{Ch}(f)$ if and only if
$\mathrm{Ch}(g)\geq\mathrm{Ch}(f)+\int\xi(g-f)\,\mathrm{d}\nu\quad\forall g\in L^{2}(\mathcal{X},\nu).$
Notice that the functional $\operatorname{Ch}:L^{2}(\mathcal{X},\nu)\to[0,\infty]$ is convex, lower semicontinuous and with dense domain, so that the standard theory of gradient flows on Hilbert space applies. This means that [7, Theorem 2.13] for any $f\in L^{2}(\mathcal{X},\nu)$ there exists a unique locally absolutely continuous curve $(f_{t})\subset L^{2}(\mathcal{X},\nu)$ on $0,\infty)$ such that $f_{0}=f$ and it holds
(2.4) $\frac{\mathrm{d}}{\mathrm{d}t}f_{t}\in-\partial^{-}\operatorname{Ch}(f_{t}),\quad\text{a.e. }t>0,$
which is the gradient flow of the Cheeger energy. The operator $P_{t}$ given by $f_{0}\mapsto f_{t}$ is known as the heat semigroup in $L^{2}(\mathcal{X},\nu)$.
In [6] properties of this gradient flow have been studied extensively, especially in connection with concepts coming from optimal transport theory like Wasserstein distance and relative entropy. In particular, the main results of [6] are critical in the identification of the gradient flows for the Cheeger energy in $L^{2}(\mathcal{X},\nu)$ and for the entropy functional in $(\mathscr{P}_{2}(\mathcal{X}),\mathsf{W}_{2})$.
It is well known that Sobolev-type inequalities (including log-Sobolev inequalities) imply Talagrand’s transport inequalities, as shown by Otto and Villani [58, 73] on Euclidean spaces and Riemannian manifolds and extended by Gigli and Ledoux [33] to Polish metric measure spaces. A nonlinear analogue of this result was given by F. Y. Wang [75, Theorem 2.1] which is proven by extending Otto and Villani’s idea. Specifically, F. Y. Wang showed that if a weighted manifold $(\mathcal{M},\nu)$ admits a $\theta$-nonlinear log-Sobolev inequality, then the following nonlinear transport inequality holds:
(2.5) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\Upsilon(0,D(\mu\|\nu)),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),$
where
$\Upsilon(s,\alpha):=\int_{s}^{\alpha}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
In the present paper, we require a variant of F. Y. Wang’s result. Given $\mu_{W}$, we consider the conditional Wasserstein metric $\mathsf{W}_{2}\left(\mu_{X|W},\mu_{Y|W}|\mu_{W}\right)$ as a distance between conditional probabilities $\mu_{X|W}$ and $\mu_{Y|W}$.
Theorem 6 (From Nonlinear Log-Sobolev to Nonlinear Transport).
Let $(\mathcal{X},d)$ be a Polish (complete and separable) metric space and $\nu$ is a Borel probability measure on $\mathcal{X}$ which admits a $\theta$-nonlinear log-Sobolev inequality for some function $\theta$. Let $\mu_{X_{0}W}$ be a joint probability measure such that $\mu_{X_{0}}=f_{0}\nu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{X})$ and $D(\mu_{X_{0}|W}\|\nu|\mu_{W})<\infty$. Let $\mu_{X_{t}}=f_{t}\nu$ be the heat flow with initial data $\mu_{X_{0}}$ and denote the induced transition probability $\mu_{X_{t}|X_{0}}$. Let $\mu_{X_{t}X_{0}W}=\mu_{X_{0}W}\mu_{X_{t}|X_{0}}$. Then, the curve $t\mapsto\mu_{X_{t}|W}$ is absolutely continuous on $0,\infty)$ with respect to the conditional Wasserstein metric $\mathsf{W}_{2}$ and for its metric speed $|\dot{\mu}_{X_{t}|W}|$, it holds that $|\dot{\mu}_{X_{t}|W}|^{2}\leq I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$. Moreover, the map $t\mapsto D(\mu_{X_{t}|W}\|\nu|\mu_{W})$ is absolutely continuous and non-increasing on $0,\infty)$, and
(2.6) $\mathsf{W}_{2}\left(\mu_{X_{0}|W},\mu_{X_{t}|W}|\mu_{W}\right)\leq\Upsilon(D(\mu_{X_{t}|W}\|\nu|\mu_{W}),D(\mu_{X_{0}|W}\|\nu|\mu_{W})).$
Remark 5.
If $W$ is constant, then the unconditional version is obtained. This theorem can be generalized to the $\Phi$-entropy version by incorporating some ideas from [73] to the proof.
Proof.
Since $\mu_{X_{0}}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{X})$, we have $\mu_{X_{0}|W=w}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{X})$ for $\mu_{W}$-a.e. $w$. Observe that for $0\leq s<t\leq 1$,
$\mathsf{W}_{2}^{2}\left(\mu_{X_{s}|W},\mu_{X_{t}|W}|\mu_{W}\right)$ $=\mathbb{E}_{\mu_{W}}[\mathsf{W}_{2}^{2}\left(\mu_{X_{s}|W},\mu_{X_{t}|W}\right)]$
$\leq(t-s)\mathbb{E}_{\mu_{W}}[\int_{s}^{t}I(\mu_{X_{r}|W}\|\nu){\text{\rm d}}r]$
(2.7) $=(t-s)\int_{s}^{t}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r,$
where the inequality follows since $\mathsf{W}_{2}\left(\mu_{X_{s}|W=w},\mu_{X_{t}|W=w}\right)\leq(t-s)\int_{s}^{t}I(\mu_{X_{r}|W=w}\|\nu){\text{\rm d}}r$ for $\mu_{W}$-a.e. $w$; see [6, p. 359]. Hence, for $0\leq s<t\leq 1$,
$\mathsf{W}_{2}\left(\mu_{X_{s}|W},\mu_{X_{t}|W}|\mu_{W}\right)\leq\frac{1}{2}(t-s+\int_{s}^{t}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r).$
By the triangle inequality for $\mathsf{W}_{2},$ this inequality extends to the case of $0\leq s<t<\infty$. The integral here is finite for any finite $t$ since
$\int_{0}^{t}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r$ $=\mathbb{E}_{\mu_{W}}[\int_{0}^{t}I(\mu_{X_{r}|W}\|\nu){\text{\rm d}}r]$
(2.8) $=\mathbb{E}_{\mu_{W}}[D(\mu_{X_{0}|W}\|\nu)-D(\mu_{X_{t}|W}\|\nu)]$
(2.9) $=D(\mu_{X_{0}|W}\|\nu|\mu_{W})-D(\mu_{X_{t}|W}\|\nu|\mu_{W}),$
and hence
(2.10) $\int_{0}^{\infty}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r\leq D(\mu_{X_{0}|W}\|\nu|\mu_{W})<\infty$
where (2.8) follows by [6, Proposition 4.22]. Hence, $t\mapsto\mu_{X_{t}|W}$ is absolutely continuous on $0,\infty)$. From (2.7), we estimate $|\dot{\mu}_{X_{t}|W}|^{2}\leq I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$.
Equations (2.9) and (2.10) imply that $t\mapsto D(\mu_{X_{t}|W}\|\nu|\mu_{W})$ is absolutely continuous and nonincreasing on $0,\infty)$. In fact, (2.9) further implies that $\frac{{\text{\rm d}}}{{\text{\rm d}}t}D(\mu_{X_{t}|W}\|\nu|\mu_{W})=-I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$.
We next prove (2.6) by following Otto and Villani’s idea in [58] and Gigli and Ledoux’s extension [33]. For $\mu_{W}$-a.e. $w$, $\theta(D(\mu_{X|W=w}\|\nu))\leq I(\mu_{X|W=w}\|\nu).$ Taking expectation and by Jensen’s inequality,
$\theta(D(\mu_{X|W}\|\nu|\mu_{W}))\leq\mathbb{E}_{\mu_{W}}[\theta(D(\mu_{X|W}\|\nu))]\leq I(\mu_{X|W}\|\nu|\mu_{W}).$
Define
$f(t)=\int_{0}^{t}\sqrt{I(\mu_{X_{r}|W}\|\nu|\mu_{W})}{\text{\rm d}}r+\Upsilon(0,D(\mu_{X_{t}|W}\|\nu|\mu_{W})).$
Then, it holds that for a.e. $t$,
$\frac{{\text{\rm d}}^{+}}{{\text{\rm d}}t}f(t)$ $\leq\sqrt{I(\mu_{X_{t}|W}\|\nu|\mu_{W})}+\frac{\frac{{\text{\rm d}}}{{\text{\rm d}}t}D(\mu_{X_{t}|W}\|\nu|\mu_{W})}{\sqrt{\theta(D(\mu_{X_{t}|W}\|\nu|\mu_{W}))}}$
$=\sqrt{I(\mu_{X_{t}|W}\|\nu|\mu_{W})}-\frac{I(\mu_{X_{t}|W}\|\nu|\mu_{W})}{\sqrt{\theta(D(\mu_{X_{t}|W}\|\nu|\mu_{W}))}}\;\leq 0,$
where $\frac{{\text{\rm d}}^{+}}{{\text{\rm d}}t}$ is the right derivative.
So, $f$ is non-increasing as a function of $t$, which implies $f(t)\leq f(0)=\Upsilon(0,D(\mu_{X_{0}|W}\|\nu|\mu_{W}))$ for any $t\geq 0$. Thus,
$\int_{0}^{t}\sqrt{I(\mu_{X_{r}|W}\|\nu|\mu_{W})}{\text{\rm d}}r\leq\Upsilon(D(\mu_{X_{t}|W}\|\nu|\mu_{W}),D(\mu_{X_{0}|W}\|\nu|\mu_{W})).$
Combining it with $|\dot{\mu}_{X_{t}|W}|^{2}\leq I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$. yields (2.6). ∎
We immediately obtain a corollary for Gaussian measure.
Corollary 1 (From Nonlinear Log-Sobolev to Nonlinear Transport for Gaussians).
Let $\gamma$ be the standard Gaussian measure on $\mathbb{R}^{k}$. Let $X_{t}=e^{-t}X+\sqrt{1-e^{-2t}}Z$ be an Ornstein–Uhlenbeck process on $\mathbb{R}^{k}$ with distribution $\mu_{t}$, where $X\sim\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathbb{R}^{k})$ and $Z\sim\gamma$ are independent. Then, for any $t\geq 0$,
$\sqrt{2D(\mu_{t}\|\gamma)}+\mathsf{W}_{2}(\mu_{t},\mu)\leq\sqrt{2D(\mu\|\gamma)}.$
The importance of Theorem 6 is as follows. Given $\tau\geq 0,\mu,\nu$, consider the following optimization problem:
(2.11) $\omega(\beta):=\inf_{\pi:\mathsf{W}_{2}(\pi,\mu)\leq\beta}D(\pi\|\nu),$
which characterizes the best tradeoff between $D(\pi\|\nu)$ and $\mathsf{W}_{2}(\pi,\mu)$ when we vary $\beta\geq 0$. In other words, within the Wasserstein ball of center $\mu$ and radius $\beta$, we aim at finding a distribution $\pi$ which minimizes $D(\pi\|\nu)$. By Sanov’s theorem in large deviations theory [30], $\omega(\beta)$ is the convergence rate of the probability that the empirical measure of i.i.d. random variables $X_{1},...,X_{n}$ with each $\sim\nu$ belongs to the Wasserstein ball of center $\mu$ and radius $\beta$. This probability naturally appears in concentration of measure and isoperimetric inequalities.
The Wasserstein gradient flow of the relative entropy starting from $\mu$ forms a feasible solution to (2.11). Specifically, the Wasserstein gradient flow is a curve that evolves to decrease $D(\pi\|\nu)$ most steeply with respect to the Wasserstein metric. If it stops until touching the Wasserstein sphere of center $\mu$ and radius $\beta$, we then obtain a feasible solution to (2.11) that is locally greedy. In other words, the Wasserstein gradient flow $\mu_{t}$ yields an upper bound on $\omega(\beta)$, although in general, $\mu_{t}$ is not the globally optimal. That is,
$\omega(\beta)$ $\leq D(\mu_{T(\beta)}\|\nu),$
where $T(\beta)$ is the time $t$ such that $\mathsf{W}_{2}(\mu_{t},\mu)=\beta$, corresponding to the escape time from the Wasserstein ball of center $\mu$ and radius $\beta$. Optimization problems like the one in (2.11) will be investigated in details in the next section.
3. From Transport to Variant Isoperimetry
3.1. Problem Formulation and Preliminaries
We now consider a variant of the isoperimetric problem in which the perimeter of a set is identified by the size of an enlargement shell of the set. We then provide a solution to this variant isoperimetric problem by using the nonlinear transport inequality obtained in the last section.
Let $(\mathcal{X},d,\nu)$ be a Polish metric measure space with $\nu$ being a Borel probability measure. Denote $(\mathcal{X}^{n},d_{n},\nu_{n})$ as the $n$-fold product space of $\mathcal{X}$ equipped with the product metric $d_{n}$ and product measure $\nu_{n}=\nu^{\otimes n}$. Define the variant isoperimetric profile as for $a\in[0,1],t\geq 0$,
(3.1) $\Gamma_{n}(a,t):=\inf_{A:\nu^{\otimes n}(A)\geq a}\nu^{\otimes n}(A^{t}).$
We set $a=e^{-n\alpha}$ and $t=\sqrt{n\tau}$. Define the isoperimetric exponent as for $\alpha,\tau\geq 0$,
(3.2) $E_{n}(\alpha,\tau)$ $:=-\frac{1}{n}\ln\Gamma_{n}(e^{-n\alpha},\sqrt{n\tau}).$
Recall that the Wasserstein metric $\mathsf{W}_{2}(\pi_{X},\pi_{Y})$ (of order $2$) between two probability measures $\pi_{X}$ and $\pi_{Y}$ is defined by
(3.3) $\mathsf{W}_{2}^{2}(\pi_{X},\pi_{Y}):=\min_{\pi\in\Pi(\pi_{X},\pi_{Y})}\int d^{2}(x,y){\text{\rm d}}\pi(x,y),$
where $\Pi(\pi_{X},\pi_{Y})$ is the set of joint probability measures on $\mathcal{X}\times\mathcal{Y}$ with marginals equal to $\pi_{X},\pi_{Y}$ respectively. The existence of the minimizers are well-known; see, e.g., [72, Theorem 1.3]. For a probability measure $\pi_{W}$ and two transition probabilities $\pi_{X|W}$ and $\nu_{X|W}$, the conditional Wasserstein metric between $\pi_{X|W}$ and $\pi_{Y|W}$ given $\pi_{W}$ is defined by
$\mathsf{W}_{2}^{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W}):=\int\mathsf{W}_{2}^{2}(\pi_{X|W=w},\pi_{Y|W=w}){\text{\rm d}}\pi_{W}(w).$
The conditional relative entropy of $\pi_{X|W}$ w.r.t. $\nu_{X|W}$ given $\pi_{W}$ is defined as
$D(\pi_{X|W}\|\nu_{X|W}|\pi_{W}):=\int D(\pi_{X|W=w}\|\nu_{X|W=w}){\text{\rm d}}\pi_{W}(w).$
Let $\nu_{X}$ and $\nu_{Y}$ be two probability measures on the space $(\mathcal{X},d)$. Define $E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})$ similarly as the exponent in (3.2) but with $\Gamma_{n}$ replaced by
$\Gamma_{n}(a,t|\nu_{X},\nu_{Y}):=\inf_{A:\nu_{X}^{\otimes n}(A)\geq a}\nu_{Y}^{\otimes n}(A^{t}).$
Define
(3.4) $\psi_{\kappa}(\alpha,\tau|\nu_{X},\nu_{Y})$ $:=\sup_{\begin{subarray}{c}\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\{0,1\}):\\ D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha\end{subarray}}\eta_{\tau,\kappa}(\pi_{XW}|\nu_{Y}),$
where
(3.5) $\eta_{\tau,\kappa}(\pi_{XW}|\nu_{Y})$ $:=\inf_{\begin{subarray}{c}\pi_{Y|XW}:\mathbb{E}[d^{2}(X,Y)]\leq\tau,\\ \mathbb{E}[\mathrm{Var}(d^{2}(X,Y)|X,W)]\leq\kappa\end{subarray}}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}).$
Denote $\psi(\alpha,\tau|\nu_{X},\nu_{Y})=\psi_{\infty}(\alpha,\tau|\nu_{X},\nu_{Y})$, equivalently in which case, the variance constraint in (3.5) is removed. Without changing the optimal value, the range of $W$ in (3.4) can be changed to any larger finite set. This point can be proven by following proof steps similar to that of [77, Theorem 20].
Based on $\psi$ and $\psi_{\kappa}$, the asymptotic expression of $E_{n}$ for compactly supported $\nu_{X}$ is characterized in our previous work, which is a generalization of Ahlswede and Zhang’s work [3].
Theorem 7 (Asymptotic Exponent for Compactly-Supported Measures [77, Theorem 12]).
Let $(\mathcal{X},d)$ be a Polish metric space equipped with two probability measures $\nu_{X}$ and $\nu_{Y}$. The following hold.
  1. (1)
    Suppose $\nu_{X}$ is compactly supported. Then, for any $\alpha\geq 0,\tau\geq 0$,
    (3.6) $\limsup_{n\to\infty}E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})\leq\lim_{\kappa\uparrow\infty}\psi_{\kappa}(\alpha,\tau|\nu_{X},\nu_{Y}).$
    In fact, a stronger result holds: Given $\delta,\kappa,\tau_{0}>0$, for sufficiently large $n$ and any set $A\subseteq\mathcal{X}^{n}$ such that $\nu_{X}^{\otimes n}(A)\geq e^{-n\alpha}$, there is a probability measure $\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\mathbb{N})$ induced by $A$ such that $D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha+\delta$ and for any $\tau\in(\delta,\tau_{0}]$,
    $\frac{1}{n}\ln\nu_{Y}^{\otimes n}(A^{\sqrt{n\tau}})\leq(1+\delta)\eta_{\tau-\delta,\kappa}(\pi_{XW}|\nu_{Y})+\delta.$
  2. (2)
    For any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi:=\left\{(\alpha,\tau):\psi(\alpha,\tau|\nu_{X},\nu_{Y})<\infty\right\}$, it holds that
    (3.7) $\liminf_{n\to\infty}E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})\geq\psi(\alpha,\tau|\nu_{X},\nu_{Y}).$
    In particular, if further assume $\nu_{X}=\nu_{Y}$, then (3.7) holds for any $\alpha\geq 0,\tau\geq 0$.
Remark 6.
Theorem 7 still holds if the constraint $\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\{0,1\})$ in (3.4) is replaced by $\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$.
Why is the asymptotic exponent determined by the relative entropy and the Wasserstein metric? The intuition is as follows. For a distribution $\pi$, denote $B_{\epsilon]}(\pi)$ as the closed ball of center $\nu$ and radius $\epsilon$ under the Lévy–Prokhorov metric. Let $\mathsf{L}_{n}:\mathbf{x}\in\mathcal{X}^{n}\mapsto\mathsf{L}_{\mathbf{x}}\in\mathscr{P}(\mathcal{X})$ be the empirical measure map, where $\mathsf{L}_{\mathbf{x}}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ is the empirical measure of $x$, and $\mathscr{P}(\mathcal{X})$ is the set of probability measures on $\mathcal{X}$. For a Polish space $\mathcal{X}$ and $\epsilon>0$, the empirically $\epsilon$-typical set of $\pi$ [53] is defined as $\mathcal{T}_{\epsilon}^{(n)}(\pi):=\mathsf{L}_{n}^{-1}(B_{\epsilon]}(\pi)).$
For an empirically $\epsilon$-typical set $A=\mathsf{L}_{n}^{-1}(B_{\epsilon}(\pi_{X}))$ (i.e., $A=\{\mathbf{x}\in\mathcal{X}^{n}:\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$), roughly speaking, by Sanov’s theorem it holds that $\nu_{X}^{\otimes n}(A)\approx e^{-nD(\pi_{X}\|\nu_{X})}$ for $\epsilon$ chosen sufficiently small, which is no smaller than $e^{-n\alpha}$ if $D(\pi_{X}\|\nu_{X})\leq\alpha$. The $\sqrt{n\tau}$-enlargement of this set can be written as
$A^{\sqrt{n\tau}}$ $=\{\mathbf{y}:\frac{1}{n}d_{n}^{2}(\mathbf{x},\mathbf{y})\leq\tau,\exists\mathbf{x}\in A\}$
$=\{\mathbf{y}:\mathbb{E}_{\mathsf{L}_{\mathbf{x},\mathbf{y}}}[d^{2}(X,Y)]\leq\tau,\mathsf{L}_{\mathbf{x}}\approx\pi_{X},\exists\mathbf{x}\}$
(3.8) $=\{\mathbf{y}:\mathsf{W}_{2}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})\leq\sqrt{\tau},\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$
$\approx\mathsf{L}^{-1}\{\pi_{Y}:\mathsf{W}_{2}(\pi_{X},\pi_{Y})\leq\sqrt{\tau}\},$
where (3.8) is due to Birkhoff’s theorem, $\min_{\sigma}\mathbb{E}_{\mathsf{L}_{\sigma(\mathbf{x}),\mathbf{y}}}[d^{2}(X,Y)]=\mathsf{W}_{2}^{2}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})$, with $\sigma$ denoting a permutation operator [72, Page 5]. By Sanov’s theorem again,
$\nu_{Y}^{\otimes n}(A^{\sqrt{n\tau}})\approx\exp\{-n\inf_{\pi_{Y}:\mathsf{W}_{2}(\pi_{X},\pi_{Y})\leq\sqrt{\tau}}D(\pi_{Y}\|\nu_{Y})\}.$
Optimizing over $\pi_{X}$ such that $D(\pi_{X}\|\nu_{X})\leq\alpha$, we obtain the isoperimetric exponent for $A$,
$\sup_{\pi_{X}:D(\pi_{X}\|\nu_{X})\leq\alpha}\inf_{\pi_{Y}:\mathsf{W}_{2}(\pi_{X},\pi_{Y})\leq\sqrt{\tau}}D(\pi_{Y}\|\nu_{Y}).$
By product construction of such $A$’s (forming a conditional empirically typical set), we obtain the exponent $\psi(\alpha,\tau|\nu_{X},\nu_{Y})$.
The intuition above provides a proof idea for Statement 2 of Theorem 7. However, Statement 1 is highly nontrivial and its proof relies on the so-called the inherently typical subset lemma [2]. This lemma enables us to tailor a set into a new one that is “regular” in some sense, so that we can analyze the asymptotic expression of the boundary size of the set.
3.2. Riemannian Energy Measure Spaces Satisfying $\text{BE}(K,\infty)$ and $\operatorname{RCD}(K,\infty)$-spaces
Our goal in the whole section is to generalize Statement 1 of Theorem 7 to probability measures satisfying Condition 2. However, since the class of Polish metric measure spaces is too broad to allow for straightforward derivation of bounds, we restrict our analysis to a subclass of Polish spaces with favorable properties—namely, Riemannian energy measure spaces satisfying the $\text{BE}(K,\infty)$ condition, or equivalently, $\text{RCD}(0,\infty)$-spaces. This choice enables us to use the well-defined heat flow to establish bounds for variant isoperimetry. To this end, we now introduce these two concepts; for further details, we refer the reader to [7, 8].
We consider an Energy measure space $(\mathcal{X},d,\nu)$, where $(\mathcal{X},d)$ is a Polish metric space, $\nu$ is a Borel probability measure with full support, and the underlying $\sigma$-algebra is the $\nu$-completion of the Borel $\sigma$-algebra. The structure is governed by the Cheeger energy $\text{Ch}:L^{2}(\mathcal{X},\nu)\to[0,\infty]$ given in (2.2), defined as the $L^{2}$-relaxation of the local Lipschitz constant. The space is assumed to be infinitesimally Hilbertian in the following sense.
  • $\mathcal{E}=2\text{Ch}$ is a quadratic form that constitutes a strongly local, symmetric Dirichlet form $(\mathcal{E},\mathbb{V})$ with domain $\mathbb{V}:=D(\mathcal{E})=\bigl\{f\in L^{2}(\mathcal{X},\nu):\mathcal{E}(f)<\infty\bigr\}$.
Define the set $\mathbb{G}\subseteq\mathbb{V}$ such that for $f\in\mathbb{G}$, the linear form
$\boldsymbol{\Gamma}[f;\varphi]:=\mathcal{E}(f,f\varphi)-\tfrac{1}{2}\mathcal{E}(f^{2},\varphi),\;\varphi\in\mathbb{V}_{\infty}:=\mathbb{V}\cap L^{\infty}(\mathcal{X},\nu)$
is represented by the $L_{+}^{1}$ density $\Gamma(f)=|\nabla f|_{*}^{2}$, satisfying $\mathcal{E}(f,f)=\int\Gamma(f)\,{\text{\rm d}}\nu$. Define $\Gamma(f,g)=\frac{1}{4}\Gamma(f+g)-\frac{1}{4}\Gamma(f-g)$ which is known as the carré du champ operator. By the uniform continuity property, we extend $\boldsymbol{\Gamma}$ to a real multilinear map defined in $\mathbb{V}\times\mathbb{V}\times\mathbb{V}_{\infty}$, which satisfies
$\mathbf{\Gamma}[f,g;\varphi]=\int\Gamma(f,g)\varphi\,{\text{\rm d}}\nu\quad\text{if }f,g\in\mathbb{G}.$
Define $\mathcal{L}:=\{\psi\in\mathbb{G}:\Gamma(\psi)\leq 1\quad\nu\text{-a.e. in }\mathcal{X}\}$ and $\mathcal{L}_{C}:=\mathcal{L}\cap C(\mathcal{X})$. The compatibility between the metric $d$ and the energy $\mathcal{E}$ is enforced by the following structural assumptions:
  • Identification: $\mathcal{L}=\mathcal{L}_{C}$. That is, every function with a gradient density bounded by one $\nu$-a.e. admits a continuous representative.
  • Lipschitz Consistency: Every function in $\mathcal{L}_{C}$ is $1$-Lipschitz with respect to $d$.
The above three assumptions turns the space $(\mathcal{X},d,\nu)$ described above into a Riemannian Energy measure space in which the intrinsic distance $d_{\mathcal{E}}$ induced by the Dirichlet form coincides with the distance $d$, generalizing smooth Riemannian manifolds to the possibly non-smooth setting.
The Dirichlet form $\mathcal{E}$ induces a densely defined self-adjoint operator $\Delta_{\mathcal{E}}:D(\Delta_{\mathcal{E}})\subset\mathbb{V}\to L^{2}(\mathcal{X},\nu)$ defined by the integration by parts formula $\mathcal{E}(f,g)=-\int g\Delta_{\mathcal{E}}f\,{\text{\rm d}}\nu$ for all $g\in\mathbb{V}$. This self-adjoint operator generates the heat flow $(P_{t})_{t\geq 0}$, an analytic Markov semigroup characterized by $\frac{\mathrm{d}}{\mathrm{d}t}P_{t}f=\Delta_{\mathcal{E}}P_{t}f$, a specialization of the heat equation in (2.4) in the Riemannian Energy measure space.
Bakry–Émery conditions are formulated via the iterated gradient form $\boldsymbol{\Gamma}_{2}$. For $(f,\varphi)\in D(\boldsymbol{\Gamma}_{2})$, define:
$\boldsymbol{\Gamma}_{2}[f;\varphi]:=\tfrac{1}{2}\boldsymbol{\Gamma}[f;\Delta_{\mathcal{E}}\varphi]-\boldsymbol{\Gamma}[f,\Delta_{\mathcal{E}}f;\varphi],$
where
$D(\Gamma_{2}):=\left\{(f,\varphi)\in D(\Delta_{\mathcal{E}})\times D(\Delta_{\mathcal{E}}):\Delta_{\mathcal{E}}f\in\mathbb{V},\,\varphi,\Delta_{\mathcal{E}}\varphi\in L^{\infty}(\mathcal{X},\nu)\right\}.$
We can write
$\mathbf{\Gamma}_{2}[f;\varphi]=\int\left(\frac{1}{2}\Gamma(f)\Delta_{\mathcal{E}}\varphi-\Gamma(f,\Delta_{\mathcal{E}}f)\varphi\right){\text{\rm d}}\nu\quad\text{if }(f,\varphi)\in D(\mathbf{\Gamma}_{2}),\ f,\Delta_{\mathcal{E}}f\in\mathbb{G}.$
The space satisfies the Bakry–Émery condition $\text{BE}(K,\infty)$ if for every $\varphi\in D(\Delta_{\mathcal{E}})$ with $\varphi\geq 0$:
$\boldsymbol{\Gamma}_{2}[f;\varphi]\geq K\boldsymbol{\Gamma}[f;\varphi].$
In the setting of smooth weighted Riemannian manifolds, this condition can be written as $\frac{1}{2}\Gamma(f)\Delta_{\mathcal{E}}-\Gamma(f,\Delta_{\mathcal{E}}f)\geq K\Gamma(f)$, or equivalently, $\mathrm{Ric}_{V}\geq Kg$ for probability measure $\nu=e^{-V}\mathrm{vol}$. Under $\text{BE}(K,\infty)$, it is known that $\mathbb{G}=\mathbb{V}$.
The Riemannian Energy measure space $(\mathcal{X},d,\nu)$ (with $d=d_{\mathcal{E}}$) satisfies $\text{BE}(K,\infty)$ if and only if it is an $\text{RCD}(K,\infty)$ space [8, Theorem 4.17]. From the perspective of the geometry of optimal transport, this equivalence implies that the relative entropy functional $D(\cdot\|\nu)$ is not merely $K$-geodesically convex on the Wasserstein space $(\mathscr{P}_{2}(\mathcal{X}),\mathsf{W}_{2})$, but that this convexity is robust: it extends to hold along all interpolated measures induced by weighted optimal geodesic plans (or test plans).
A fundamental consequence of the $\operatorname{RCD}(K,\infty)$ property (equivalently, the Riemannian Energy measure space with $\text{BE}(K,\infty)$ property) is the identification of the metric $\mathsf{W}_{2}$-gradient flow $\mathcal{H}_{t}$ with the analytic $L^{2}$-heat flow $P_{t}$. For any $f\in L^{2}(\mathcal{X},\nu)$ such that $f\nu\in\mathscr{P}_{2}(\mathcal{X})$, the identification $\mathcal{H}_{t}(f\nu)=(P_{t}f)\nu$ holds. Furthermore, the heat flow admits the representation:
$(P_{t}f)\nu=\int f(x)\mathcal{H}_{t}(\delta_{x})\,{\text{\rm d}}\nu(x),$
characterizing the evolved measure as a weighted superposition of the evolved Dirac masses.
3.3. Variant Isoperimetry for Noncompactly-Supported Measures
We next extend Statement 1 of Theorem 7 to the Riemannian Energy measure space satisfying $\text{BE}(K,\infty)$ and Condition 2 (equivalently, a $\mathrm{RCD}(0,\infty)$-space satisfying Condition 2). To this end, we need the following conditions on the covering bound, volume ratio bound, and curvature-dimension bound.
Condition 4 (Covering Bound).
Let $x_{0}\in\mathcal{X}$ be a fixed point. Then, for any $n\geq 1$, $\lambda\geq 1$, and $r>0$, there is cover of the large ball $B_{\sqrt{n}\lambda r}(\mathbf{x}_{0})$ with $\mathbf{x}_{0}=(x_{0},...,x_{0})\in\mathcal{X}^{n}$ that consists of small balls $B_{\sqrt{n}r}(\mathbf{x}_{i})$ with $\mathbf{x}_{i}\in\mathcal{X}^{n}$ and $1\leq i\leq N$, i.e., $B_{\sqrt{n}\lambda r}(\mathbf{x}_{0})\subseteq\bigcup_{1\leq i\leq N}B_{\sqrt{n}r}(\mathbf{x}_{i})$, where $N=N_{n,\lambda,r}$ satisfies
$\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N_{n,\lambda,r}=0,\;\forall\lambda\geq 1.$
Condition 5 (Volume Ratio Bound).
There is a nonnegative (not necessarily finite or probability) measure $\mathfrak{m}$ on $\mathcal{X}^{n}$ such that for all $\lambda\geq 1$,
(3.9) $\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N_{n}(\sqrt{n}\lambda r,\sqrt{n}r)=0,$
where
(3.10) $N_{n}(R,r)=\sup_{\mathbf{x}\in\mathcal{X}^{n}}\frac{\mathfrak{m}(B_{R}(\mathbf{x}))}{\mathfrak{m}(B_{r}(\mathbf{x}))}.$
Condition 6 (Curvature-Dimension Bound).
Suppose $\mathcal{X}$ is a subset of a Riemannian manifold $(\mathcal{M},g)$ or the interior of $\mathcal{X}$ is a Riemannian manifold $(\mathcal{M},g)$. There is a nonnegative (not necessarily finite or probability) measure $\mathfrak{m}=e^{-U}\mathrm{vol}$ on manifold $\mathcal{M}$ with $U\in C^{2}(\mathcal{M})$ satisfying $\mathrm{CD}(K,N)$ for some $N\in(k,\infty)$ and $K\in\mathbb{R}$ with $k=\mathrm{dim}(\mathcal{M})$:
$\mathrm{Ric}_{U}^{N}:=\mathrm{Ric}_{g}+\mathrm{Hess}_{g}U-\frac{\nabla U\otimes\nabla U}{N-k}\geq Kg.$
The conditions specified above and Condition 3 satisfy the following implication relation.
Proposition 3.
It holds that Condition 5 $\Longrightarrow$ Condition 4. Moreover, if $\mathcal{X}$ is a subset of a Riemannian manifold $(\mathcal{M},g)$ or the interior of $\mathcal{X}$ is a Riemannian manifold $(\mathcal{M},g)$, then Condition 3 $\Longrightarrow$ Condition 6 $\Longrightarrow$ Condition 5 $\Longrightarrow$ Condition 4.
Define
$\Upsilon_{\alpha}(s)=\int_{s}^{\alpha}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Define
(3.11) $\bar{\psi}(\alpha,\tau)$ $=\begin{cases}\Upsilon_{\alpha}^{-1}(\sqrt{\tau}),&\Upsilon_{\alpha}(0)\geq\sqrt{\tau}\\ 0,&\Upsilon_{\alpha}(0)<\sqrt{\tau}\end{cases},$
where for each $\alpha$, denote $\Upsilon_{\alpha}^{-1}$ as the inverse of $\Upsilon_{\alpha}$. We are ready to extend Statement 1 of Theorem 7 to all probability measures satisfying Condition 2 or equivalently admitting a standard log-Sobolev inequality.
Theorem 8 (Asymptotic Exponent for Noncompactly-Supported Measures).
Let $(\mathcal{X},d,\nu)$ be a Riemannian Energy measure space satisfying $\text{BE}(K,\infty)$ (equivalently, it is a $\mathrm{RCD}(0,\infty)$-space). Suppose that this space satisfies Condition 2 and it admits a $\theta$-nonlinear log-Sobolev inequality. If the condition in (1.10) holds, we further assume Condition 4 (or Conditions 3, 6, or 5) holds. Then, for any $\alpha\geq 0,\tau\geq 0$,
$\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\bar{\psi}(\alpha,\tau).$
Remark 7.
Under $\mathrm{RCD}(0,\infty)$ (equivalently, the Riemannian Energy setting with $\mathrm{BE}(0,\infty)$) for a probability measure $\nu\in\mathscr{P}_{2}(\mathcal{X})$, Condition 2 holds if and only if it admits a standard log-Sobolev inequality [9, Theorem 3.2].
Remark 8.
Define
$\psi(\alpha,\tau)$ $:=\psi(\alpha,\tau|\nu,\nu)$
$=\sup_{\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}$
(3.12) $\qquad\inf_{\pi_{Y|W}:\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\sqrt{\tau}}D(\pi_{Y|W}\|\nu|\pi_{W}),$
and
$\tilde{\psi}(\alpha,\tau)=\frac{1}{2}\left[\sqrt{2\alpha}-\sqrt{K_{LS}^{+}\tau}\right]_{+}^{2},$
where recall the definition of $K_{LS}^{+}$ in (1.15). By checking our proof, the following bound can be obtained:
$\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\sup_{\alpha_{0}+\alpha_{1}\leq\alpha}\inf_{\tau_{0}+\tau_{1}\leq\tau}\psi(\alpha_{0},\tau_{0})+\tilde{\psi}(\alpha_{1},\tau_{1}).$
This bound is more consistent with the lower bound in Statement 2 of Theorem 7, and it coincides with the latter if $\psi\leq\tilde{\psi}$.
3.4. Proof of Proposition 3
By choosing $\mathfrak{m}$ as the Riemannian volume measure, Condition 3 $\Longrightarrow$ Condition 6 is obvious. We next show that Condition 5 $\Longrightarrow$ Condition 4.
For a subset $A\subseteq\mathcal{M}^{n}$, denote by $N(A,r)$ the minimal covering number (i.e., the smallest number of balls of radius $r$ needed to cover $A$) and by $M(A,r)$ the maximal packing number (i.e., the largest number of disjoint balls of radius $r$ whose centers lie in $A$). By the standard lemma for metric spaces, $N(A,r)\leq M(A,r/2).$ Setting $A=B_{R}(\mathbf{x})$ yields that $N(B_{R}(\mathbf{x}),r)\leq M(B_{R}(\mathbf{x}),r/2).$
For any ball $B(\mathbf{x}_{i},r/2)$ in the packing for $M(B_{R}(\mathbf{x}),r/2)$, since $p_{i}\in B_{R}(\mathbf{x})$, all such balls are contained in $B_{R+r/2}(\mathbf{x})$. Let $i^{*}$ minimize $\mathfrak{m}^{\otimes n}(B_{r/2}(\mathbf{x}_{i}))$ over all $i$. Combining this with the fact that $B_{R+r/2}(\mathbf{x})\subseteq B_{2R+r/2}(\mathbf{x}_{i^{*}})$, we obtain:
$M(B_{R}(\mathbf{x}),r/2)\cdot\mathfrak{m}^{\otimes n}(B_{r/2}(\mathbf{x}_{i^{*}}))\leq\mathfrak{m}^{\otimes n}(B_{R+r/2}(\mathbf{x}))\leq\mathfrak{m}^{\otimes n}(B_{2R+r/2}(\mathbf{x}_{i^{*}})).$
Therefore, for any $\mathbf{x}$ and any $R>r>0$,
$N(B_{R}(\mathbf{x}),r)\leq M(B_{R}(\mathbf{x}),r/2)\leq\frac{\mathfrak{m}^{\otimes n}(B_{2R+r/2}(\mathbf{x}_{i^{*}}))}{\mathfrak{m}^{\otimes n}(B_{r/2}(\mathbf{x}_{i^{*}}))}\leq N_{n}(2R+r/2,r/2).$
Substituting $(R,r)\leftarrow(\sqrt{n}\lambda r,\sqrt{n}r)$ into the inequality above and by the condition (3.9), we obtain that
$\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N(B_{\sqrt{n}\lambda r}(\mathbf{x}_{0}),\sqrt{n}r)=0,\;\forall\lambda\geq 1.$
Therefore, Condition 4 holds.
We now prove that Condition 6 $\Longrightarrow$ Condition 5. Let $\mathfrak{m}=e^{-U}\mathrm{vol}$ be the nonnegative measure on manifold $\mathcal{M}$ given in Condition 6. Then, $\mathfrak{m}^{\otimes n}=e^{-U^{\oplus n}}\mathrm{vol}_{n}$ is a nonnegative measure on manifold $\mathcal{M}^{n}$ with $U^{\oplus n}\in C^{2}(\mathcal{M}^{n})$ such that $\mathrm{Ric}_{U^{\oplus n}}^{nN}\geq nK$.
For a point $p$ in $\mathcal{M}$ and $\xi\in T_{p}\mathcal{M}$ such that $|\xi|=1$, let $\gamma_{\xi}(t)=\exp_{p}(t\xi)$ be the geodesic starting at $p$ with direction $\xi$, and let $\lambda_{p}(t,\xi)=\lambda_{p}^{U}(t,\xi)$ be the density of the weighted measure $\mathfrak{m}$ at the point $\gamma_{\xi}(t)$. The density $\lambda_{p}(t,\xi)$ is defined to be zero whenever $t$ exceeds the distance to the cut point in the direction $\xi$. This density incorporates both the Riemannian volume distortion (Jacobian of the exponential map) and the weight function $e^{-U}$. Let $\lambda^{(K)}(t)$ be the Riemannian volume density in the $(\mathrm{dim}(\mathcal{M})+N)$-dimensional model space $\mathcal{M}^{(K)}$ with constant Ricci curvature $K$. Following [60, Theorem 7], under the condition $\mathrm{Ric}_{U}^{N}\geq Kg$, it holds that for any $\xi$, $\frac{\lambda_{p}(t,\xi)}{\lambda^{(K)}(t)}$ is non-increasing in $t$ for $\gamma_{\xi}(t)\notin\mathrm{cut}(p)$, where $\mathrm{cut}(p)$ is the cut locus of the Riemannian manifold $M$ with respect to the point $p$.
Let $A(p,t)$ be the $(\mathrm{dim}(\mathcal{M})-1)$-dimensional $\mathfrak{m}$-weighted area of a geodesic sphere of radius $t$ at center $p$ in $M$. Then, for any $t>0$,
$A(p,t)$ $=\int_{S_{p}\mathcal{M}}\lambda_{p}(t,\xi)\,{\text{\rm d}}\mathrm{area}(\xi)$
where $S_{p}\mathcal{M}$ is the unit sphere in the tangent space $T_{p}\mathcal{M}$ at point $p$, and ${\text{\rm d}}\mathrm{area}(\xi)$ is the standard area element of the unit (n1)(n-1)-sphere in the tangent space.
Let $A^{(K)}(t)$ be the $(\mathrm{dim}(\mathcal{M})-1)$-dimensional Riemannian area of a geodesic sphere of radius $t$ at center $p$ in the $\mathrm{dim}(\mathcal{M})$-dimensional model space $\mathcal{M}^{(K)}$ with constant Ricci curvature $K$. Then,
$A^{(K)}(t)$ $=\lambda^{(K)}(t)\cdot|S_{p}\mathcal{M}^{(K)}|,$
where $S_{p}\mathcal{M}^{(K)}$ is the unit sphere in the tangent space $T_{p}\mathcal{M}^{(K)}$ at point $p$, and $|S_{p}\mathcal{M}^{(K)}|$ is the standard area of the unit (n1)(n-1)-sphere in the tangent space. Note that the value of $|S_{p}\mathcal{M}^{(K)}|$ is independent of $p$.
Observe that
$\frac{A(p,t)}{A^{(K)}(t)}=\frac{1}{|S_{p}\mathcal{M}^{(K)}|}\int_{S_{p}\mathcal{M}}\frac{\lambda_{p}(t,\xi)}{\lambda^{(K)}(t)}\,{\text{\rm d}}\mathrm{area}(\xi),$
which is non-increasing in $t$. The volume of the ball with center $\mathbf{p}=(p_{1},...,p_{n})$ and radius $R$ in the space $\mathcal{M}^{n}$ is
$V_{n}(\mathbf{p},R)=\int_{\sum_{i=1}^{n}\rho_{i}^{2}\leq R^{2}}\prod_{i=1}^{n}A(p_{i},\rho_{i}){\text{\rm d}}\rho_{i}.$
Similarly, the volume of the ball with radius $R$ (at any center) in the space $(\mathcal{M}^{(K)})^{n}$ is
$V_{n}^{(K)}(R)$ $=\int_{\sum_{i=1}^{n}\rho_{i}^{2}\leq R^{2}}\prod_{i=1}^{n}A^{(K)}(\rho_{i}){\text{\rm d}}\rho_{i}.$
We have the following key lemma.
Lemma 1.
Given any $\mathbf{p}$, $f(R)=\frac{V_{n}(\mathbf{p},R)}{V_{n}^{(K)}(R)}$ is non-increasing.
Proof.
To show that the ratio $f(R)=\frac{V_{n}(\mathbf{p},R)}{V_{n}^{(K)}(R)}$ is non-increasing, we utilize the properties of the Bishop-Gromov comparison theorem and the behavior of ratios of integrals.
Step 1: As shown above, the ratio of the areas of geodesic spheres $\frac{A(p,r)}{A^{(K)}(r)}$ is a non-increasing function of $r$. Let $\rho=(\rho_{1},\dots,\rho_{n})$ be a vector in the integration domain. For any scalar $r>0$ and a unit vector $\omega\in S_{+}^{n-1}=\{\omega\in S^{n-1}:\omega_{i}\geq 0\text{\ for\ all\ }i\}$, let $\rho=r\omega$. The product ratio is: $\frac{\prod_{i=1}^{n}A(p_{i},r\omega_{i})}{\prod_{i=1}^{n}A^{(K)}(r\omega_{i})}=\prod_{i=1}^{n}\frac{A(p_{i},r\omega_{i})}{A^{(K)}(r\omega_{i})}$. Since each term $\frac{A(p_{i},r\omega_{i})}{A^{(K)}(r\omega_{i})}$ is non-increasing in $r$ (because $r\omega_{i}$ increases with $r$ for $\omega_{i}\geq 0$), the product of these non-negative non-increasing functions is also non-increasing in $r$.
Step 2: We express the volumes $V_{n}(R)$ and $V_{n}^{(K)}(R)$ using polar coordinates in the $\rho$-space $\mathbb{R}^{n}$$V_{n}(R)=V_{n}(\mathbf{p},R)=\int_{0}^{R}r^{n-1}\left(\int_{S_{+}^{n-1}}\prod A(p_{i},r\omega_{i})d\omega\right)dr$. Let $G(r)=G(\mathbf{p},r)=r^{n-1}\int_{S_{+}^{n-1}}\prod A(p_{i},r\omega_{i})d\omega$ and $H(r)=r^{n-1}\int_{S_{+}^{n-1}}\prod A^{(K)}(r\omega_{i})d\omega$. From Step 1, the ratio of the integrands $\frac{\prod A(p_{i},r\omega_{i})}{\prod A^{(K)}(r\omega_{i})}$ is non-increasing in $r$ for every $\omega$. Consequently, the ratio of the integrals over the angular components, $\frac{G(r)}{H(r)}$, is also non-increasing.
Step 3: A standard lemma in analysis states that if $G(r)$ and $H(r)$ are positive functions such that $\frac{G(r)}{H(r)}$ is non-increasing, then the ratio of their cumulative integrals $F(R)=\frac{\int_{0}^{R}G(r)dr}{\int_{0}^{R}H(r)dr}$ is also non-increasing. To see this, calculate the derivative $F^{\prime}(R)$: $F^{\prime}(R)=\frac{G(R)\int_{0}^{R}H(r)dr-H(R)\int_{0}^{R}G(r)dr}{(\int_{0}^{R}H(r)dr)^{2}}$. Since $\frac{G(r)}{H(r)}\geq\frac{G(R)}{H(R)}$ for all $r\leq R$, we have $H(R)G(r)\geq G(R)H(r)$. Integrating both sides w.r.t $r$ from 0 to $R$ yields $H(R)\int_{0}^{R}G(r)dr\geq G(R)\int_{0}^{R}H(r)dr$. Thus, $F^{\prime}(R)\leq 0$. The function $f(R)=\frac{V_{n}(R)}{V_{n}^{(K)}(R)}$ is non-increasing for $R>0$. ∎
In a model space $M^{(K)}$ of dimension $k=\mathrm{dim}(\mathcal{M})+N$, the area $A^{(K)}(\rho)$ of a geodesic sphere of radius $\rho$ is given by
$A^{(K)}(\rho)=S_{k-1}\cdot\mathrm{sn}_{\kappa}(\rho)^{k-1},$
where $S_{k-1}=\frac{2\pi^{k/2}}{\Gamma(k/2)}$ is the area of a unit (k1)(k-1)-sphere (in Euclidean spaces), and the function $\mathrm{sn}_{\kappa}(\rho)$ is defined as
$\mathrm{sn}_{\kappa}(\rho)=\begin{cases}\frac{1}{\sqrt{\kappa}}\sin(\sqrt{\kappa}\rho),&K>0,\\ \rho,&K=0,\\ \frac{1}{\sqrt{|\kappa|}}\sinh(\sqrt{|\kappa|}\rho),&K<0,\end{cases}$
with $\kappa=\frac{K}{k-1}$.
To prove the desired result, it suffices to only consider the weaker case, $K<0$. Recalling the definition of $N_{n}$ in (3.10) and by Lemma 1,
$N_{n}(\sqrt{n}R,\sqrt{n}r)=\sup_{\mathbf{p}}\frac{V_{n}(\mathbf{p},\sqrt{n}R)}{V_{n}(\mathbf{p},\sqrt{n}r)}\leq\frac{V_{n}^{(K)}(\sqrt{n}R)}{V_{n}^{(K)}(\sqrt{n}r)}.$
By large deviations theory (or Laplace’s Method), one can determine the asymptotics of $V_{n}^{(K)}(\sqrt{n}R)$ as follows
$\lim_{n\rightarrow\infty}\frac{1}{n}\ln V_{n}^{(K)}(\sqrt{n}R)=\sqrt{|K|(k-1)}R.$
Hence, substituting $R\leftarrow\lambda r$, we obtain that
$\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N_{n}(\sqrt{n}\lambda r,\sqrt{n}r)=\lim_{r\to\infty}\frac{1}{r^{2}}\sqrt{|K|(k-1)}(\lambda-1)r=0.$
3.5. Proof of Theorem 8
We now provide the proof of Theorem 7 with the outline given as follows. By Sanov’s theorem, we may restrict most coordinates to compact sets without altering the exponential order of the set’s size, though the remaining coordinates cannot be constrained. Fortunately, in the case that the condition (1.9) holds, the measure $\nu^{\otimes n}$ exhibits stronger concentration than Gaussian measures, which allows us to cover the majority of the mass of the remaining coordinates at only a negligible enlargement cost (i.e., in terms of distance). In other words, the isoperimetric exponent in this case is determined by the one for compact sets and the latter was already characterized in Theorem 7. On the other hand, in the case that the condition (1.10) holds, under the covering bound assumption (Condition 4), the majority of the mass of the remaining coordinates can be covered by subexponentially many small balls with negligible radius. Thus, there is one small ball that dominates others in the sense that it has almost the same mass (up to a subexponential factor) as the whole mass of the remaining coordinates. Covering this small ball will cover the majority of the mass of the remaining coordinates at only a negligible enlargement cost (i.e., in terms of distance). In this case, the isoperimetric exponent is also determined by the one for compact sets characterized in Theorem 7. We next present the detailed proof.
3.5.1. Step 1: Tailoring the Set
Let $A_{n}$ be a sequence of optimal subsets asymptotically attaining the optimal exponent $\limsup_{n\to\infty}E_{n}(\alpha,\tau)$. Specifically, we let $A_{n}$ be a sequence of subsets such that $\nu^{\otimes n}(A_{n})\geq e^{-n\alpha}$ and $\frac{1}{n}\ln\nu^{\otimes n}(A_{n}^{\sqrt{n\tau}})\geq E_{n}(\alpha,\tau)-\frac{1}{n}.$ When there is no ambiguity, for brevity we denote $A=A_{n}$.
Since $\mathcal{X}$ is Polish, any probability measure on it is tight. So, for any $\eta\in(0,1)$, there is a compact set $F=F_{\eta}\subseteq\mathcal{X}$ such that $\nu(F^{c})\leq\eta$. Let $\mathbf{X}=(X_{1},...,X_{n})\sim\nu^{\otimes n}$ and $S_{i}:=\mathbbm{1}_{F^{c}}(X_{i})$, $i\in[n]$. Then, $\mathbf{S}\sim\mathrm{Bern}(\nu(F^{c}))^{\otimes n}$. By Sanov’s theorem, for any $\epsilon\in(\eta,1)$,
$\mathbb{P}\{|\mathbf{S}|>n\epsilon\}\leq e^{-nD(\epsilon\|\nu(F^{c}))}\leq e^{-nD(\epsilon\|\eta)},$
where $|\mathbf{S}|=\sum_{i=1}^{n}S_{i}$ denotes the number of “1” in $\mathbf{S}$, $D(\epsilon\|\eta)=\epsilon\ln\frac{\epsilon}{\eta}+(1-\epsilon)\ln\frac{1-\epsilon}{1-\eta}$ is the binary relative entropy function, and the second inequality follows since $\eta\mapsto D(\epsilon\|\eta)$ is decreasing for $\eta<\epsilon$. Since $\eta\mapsto D(\epsilon\|\eta)$ goes to infinity as $\eta\downarrow 0$, given any $\epsilon>0$, we can choose $\eta=\eta_{\epsilon}$ small enough so that $D(\epsilon\|\eta)>\alpha$. For example, we can choose $\eta=\epsilon e^{-\epsilon^{-2}}$ and choose $\epsilon$ small enough. By basic probabilistic inequalities, we have that
$\mathbb{P}\{\mathbf{X}\in A,\,|\mathbf{S}|\leq n\epsilon\}$ $\geq\mathbb{P}\{\mathbf{X}\in A\}-\mathbb{P}\{|\mathbf{S}|>n\epsilon\}\geq e^{-n\alpha}-e^{-nD(\epsilon\|\eta)},$
which implies that
(3.13) $\mathbb{P}\{\mathbf{X}\in A,\,|\mathbf{S}|\leq n\epsilon\}\geq e^{-n(\alpha+o(1))}$
with $o(1)$ denoting a term vanishing as $n\to\infty$.
Under the condition $|\mathbf{S}|\leq n\epsilon$, the random vector $\mathbf{S}$ takes ${n\choose\leq n\epsilon}:=\sum_{i=1}^{\left\lfloor n\epsilon\right\rfloor}{n\choose i}$ distinct vector values. By Sanov’s theorem, ${n\choose\leq n\epsilon}\leq e^{nH(\epsilon)}$, where $H(\epsilon)=-\epsilon\ln\epsilon-(1-\epsilon)\ln(1-\epsilon)$ is the binary entropy function. Combining this with (3.13) yields that
(3.14) $\max_{|\mathbf{s}|\leq n\epsilon}\mathbb{P}\{\mathbf{X}\in A,\mathbf{S}=\mathbf{s}\}\geq e^{-n(\alpha+H(\epsilon)+o(1))}.$
Let $\mathbf{s}^{*}$ be the optimal $\mathbf{s}$ attaining the maximum in the above equation. By rearrangement of the coordinates of $\mathbf{s}^{*}$, we assume $\mathbf{s}^{*}=(0,...,0,1,...,1)$ such that $|\mathbf{s}^{*}|\leq n_{1}:=n\epsilon$. (For brevity, $n\epsilon$ is assumed to be an integer, otherwise, denote $n_{1}:=\left\lfloor n\epsilon\right\rfloor$ and our proof still works.) Let $n_{0}=n-n_{1}=(1-\epsilon)n$. Then, (3.14) implies that $A\cap(F^{n_{0}}\times\mathcal{X}^{n_{1}})$ has probability at least $e^{-n(\alpha+H(\epsilon)+o(1))}$.
By Chernoff’s inequality, for $\mathbf{X}_{1}\sim\nu^{\otimes n_{1}}$,
$\mathbb{P}\{d^{2}(\mathbf{X}_{1},\mathbf{v}_{0})>n\beta\}\leq e^{-n_{1}\left(\lambda\beta/\epsilon-\Psi(\lambda)\right)}=e^{-n\left(\lambda\beta-\epsilon\Psi(\lambda)\right)},$
where $\mathbf{v}_{0}=(x_{0},...,x_{0})\in\mathcal{X}^{n_{1}}$, $\Psi(\lambda):=\ln\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]$ is the logarithmic moment generating function of $d^{2}(X,x_{0})$, and $\lambda$ is chosen such that $\Psi(\lambda)<\infty$ (by Condition 2, such $\lambda$ exists).
In the case that the condition (1.10) holds, we choose $\beta$ large enough so that $\lambda\beta-\epsilon\Psi(\lambda)>\alpha+H(\epsilon)$ (e.g., $\beta=\beta_{\alpha}:=\frac{\alpha+1+\Psi(\lambda)}{\lambda}$). In the case that the condition (1.9) holds, for any small $\beta>0$, we can choose $\lambda$ large enough and $\epsilon$ small enough so that $\lambda\beta-\epsilon\Psi(\lambda)>\alpha+H(\epsilon)$. In both these two cases, $A\cap(F^{n_{0}}\times B_{\sqrt{n\beta}}(\mathbf{v}_{0}))$ has probability at least $e^{-n(\alpha+H(\epsilon)+o(1))}$.
3.5.2. Step 2: Covering Argument
We next quantize the ball $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ by a covering argument. Specifically, we cover the ball $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ by using a collection of small balls of radius $\sqrt{n\delta}$ for a given small $\delta>0$. In the case that the condition (1.9) holds, $\beta$ can be arbitrarily small. Thus, in this case, we can set $\beta\leq\delta$ and $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ is covered by itself.
We next deal with the case that the condition (1.10) holds. In this case, we need the covering bound assumption (Condition 4). By this assumption, for any $\delta>0$, there is a cover of $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ consisting of balls $B_{\sqrt{n\delta}}(\mathbf{v}_{i})$ with $\mathbf{v}_{i}\in\mathcal{X}^{n_{1}}$ and $1\leq i\leq N$, where $N=e^{n\epsilon^{\prime}}$ with $\epsilon^{\prime}=\epsilon^{\prime}_{n,\epsilon,\beta,\delta}$ satisfying
$\lim_{\epsilon\to 0}\limsup_{n\to\infty}\epsilon^{\prime}_{n,\epsilon,\beta,\delta}=0,\;\forall\beta>\delta>0.$
Thus, there is at least one small ball $B:=B_{\sqrt{n\delta}}(\mathbf{v}_{i^{*}})$ for some $i^{*}$ such that
$\tilde{A}:=A\cap(F^{n_{0}}\times B)$
has probability at least $e^{-n(\alpha+H(\epsilon)+\epsilon^{\prime}+o(1))}$. Note that for fixed $\beta,\delta$, the exponent of this probability approaches $\alpha$, as $n\to\infty$ first and $\epsilon\to 0$ then.
Denote $\tilde{A}_{0}:=\{\mathbf{u}\in F^{n_{0}}:\exists\mathbf{v},(\mathbf{u},\mathbf{v})\in\tilde{A}\}$ as the the projection of $\tilde{A}$ on the first $n_{0}$ components. Then, $\tilde{A}_{0}\times B$ is a good isoperimetric approximation of $A$ in the sense that $\tilde{A}_{0}\times B$ has probability at least $e^{-n(\alpha+H(\epsilon)+\epsilon^{\prime}+o(1))}$ (since $\tilde{A}\subseteq\tilde{A}_{0}\times B$) and
$A^{\sqrt{n\tau}}\supseteq\tilde{A}^{\sqrt{n\tau}}$ $=\bigcup_{\mathbf{x}\in\tilde{A}}B_{\sqrt{n\tau}}(\mathbf{x})$
$\supseteq\bigcup_{(\mathbf{u},\mathbf{v})\in\tilde{A}}B_{\sqrt{n\tau_{0}}}(\mathbf{u})\times B_{\sqrt{n\tau_{1}}+2\sqrt{n\delta}}(\mathbf{v})$
$\supseteq\bigcup_{(\mathbf{u},\mathbf{v})\in\tilde{A}}B_{\sqrt{n\tau_{0}}}(\mathbf{u})\times B_{\sqrt{n\tau_{1}}+\sqrt{n\delta}}(\mathbf{v}_{i^{*}})$
(3.15) $=\tilde{A}_{0}^{\sqrt{n\tau_{0}}}\times B^{\sqrt{n\tau_{1}}},$
where $\tau_{0},\tau_{1}$ are arbitrary nonnegative numbers such that $\tau_{0}+\left(\sqrt{\tau_{1}}+2\sqrt{\delta}\right)^{2}\leq\tau$. Let $\alpha_{0,n}=-\frac{1}{n}\ln\nu^{\otimes n_{0}}(\tilde{A}_{0})$ and $\alpha_{1,n}=-\frac{1}{n}\ln\nu^{\otimes n_{1}}(B)$. By passing to subsequence, we can assume that $\alpha_{0,n}\to\alpha_{0}$ and $\alpha_{1,n}\to\alpha_{1}$ as $n\to\infty$ for some nonnegative numbers $\alpha_{0},\alpha_{1}$ such that $\alpha_{0}+\alpha_{1}\leq\alpha+H(\epsilon)+\epsilon^{\prime}$.
3.5.3. Step 3: Information-Theoretic Bound
We next use information-theoretic techniques to derive bounds. Denote $\nu_{F}=\nu(\cdot|F)$. Applying Theorem 7 with substitution $(n,\alpha,\tau,\nu_{X},\nu_{Y},A)\leftarrow(n_{0},\alpha_{0},\tau_{0},\nu_{F},\nu,\tilde{A}_{0})$ yields that given any $\delta>0$, there is a probability measure $\pi_{XW}\in\mathscr{P}_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ induced by the set $\tilde{A}_{0}$ such that $D(\pi_{X|W}\|\nu_{F}|\pi_{W})\leq\frac{\alpha_{0}}{1-\epsilon}+\delta$ and for sufficiently large $n$,
(3.16) $\frac{1}{n_{0}}\ln\nu^{\otimes n_{0}}(\tilde{A}_{0}^{\sqrt{n\tau_{0}}})\leq\inf_{\begin{subarray}{c}\pi_{Y|XW}:\mathbb{E}[d^{2}(X,Y)]\leq\tau_{0}-\delta,\\ \mathbb{E}[\mathrm{Var}(d^{2}(X,Y)|X,W)]\leq\kappa\end{subarray}}(1+\delta)D(\pi_{Y|W}\|\nu|\pi_{W})+\delta.$
We choose $\delta$ here as the same $\delta$ in the radius of the small ball described above.
We next derive a multiletter bound for exponent of probability of $B^{\sqrt{n\tau_{1}}}$. Let $\mu_{\mathbf{X}}=\nu^{\otimes n_{1}}(\cdot|B)$ be the conditional probability measure of $\nu^{\otimes n_{1}}$ given $B$. Then,
(3.17) $D(\mu_{\mathbf{X}}\|\nu^{\otimes n_{1}})=-\log\nu^{\otimes n_{1}}(B).$
Let $\mu_{\mathbf{Y}|\mathbf{X}}$ be a transition probability defined from $\mathcal{X}^{n_{1}}$ to $\mathcal{Y}^{n_{1}}$ such that given each $\mathbf{x}$, $\mu_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}$ is concentrated on the cost ball $B_{\sqrt{n\tau_{1}}}(\mathbf{x}):=\{\mathbf{y}:d(\mathbf{x},\mathbf{y})\leq\sqrt{n\tau_{1}}\}$. Then, we have that $\mu_{\mathbf{Y}}:=\mu_{\mathbf{X}}\circ\mu_{\mathbf{Y}|\mathbf{X}}$ is concentrated on $B^{\sqrt{n\tau_{1}}}$, which implies that $\frac{1}{n_{1}}\ln\nu^{\otimes n_{1}}(B^{\sqrt{n\tau_{1}}})\leq\frac{1}{n_{1}}D_{0}(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})\leq\frac{1}{n_{1}}D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})$. Here $D_{0}(\mu\|\nu):=-\ln\nu\{\frac{\mathrm{d}\mu}{\mathrm{d}\nu}>0\}$ is the Rényi divergence of order 0, which is no greater than the relative entropy $D(\mu\|\nu)$ [71]. Since $\mu_{\mathbf{Y}|\mathbf{X}}$ is arbitrary, we have that
(3.18) $\log\nu^{\otimes n_{1}}(B^{\sqrt{n\tau_{1}}})\leq\inf_{\mu_{\mathbf{Y}|\mathbf{X}}:d(\mathbf{X},\mathbf{Y})\leq t_{1}\;\mu_{\mathbf{XY}}\textrm{-a.e.}}D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}}).$
We now construct a feasible solution to the infimization above. Given $M,\tau_{1},\delta$, and $\mu_{\mathbf{X}}$, let $\mu_{Y|X}$ be a transition probability such that the induced joint distribution $\mu_{\mathbf{X}\mathbf{Y}}=\mu_{\mathbf{X}}\mu_{Y|X}^{\otimes n_{1}}$ satisfies
(3.19) $\theta(\mathbf{x}):=\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},Y_{i})|x_{i}]\leq n(\tau_{1}-\delta)\;\mu_{\mathbf{X}}\textrm{-a.e.}\textrm{ and }\;\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},Y_{i})|x_{i})\leq nM\;\mu_{\mathbf{X}}\textrm{-a.e.}$
In the following, we modify $\mu_{\mathbf{Y}|\mathbf{X}}$ by truncating it into some high-probability set, in order to make it satisfy the constraint $d(\mathbf{X},\mathbf{Y})\leq\sqrt{n\tau_{1}}$ $\mu_{\mathbf{XY}}\textrm{-a.e.}$
Define
$\theta(x)=\mathbb{E}_{\mu}\left[d^{2}(x,Y)|x\right].$
Then, $\theta(\mathbf{x})=\sum_{i=1}^{n_{1}}\theta(x_{i}).$
By Chebyshev’s inequality, it holds that
$\mu\{\mathbf{Y}\notin B_{\sqrt{n\tau_{1}}}(\mathbf{x})|\mathbf{x}\}$ $=\mu\{d(\mathbf{x},\mathbf{Y})>\sqrt{n\tau_{1}}|\mathbf{x}\}\leq\frac{\mathbb{E}_{\mu}\left[(d^{2}(\mathbf{x},\mathbf{Y})-\theta(x_{i}))^{2}|\mathbf{x}\right]}{(n\tau_{1}-\theta(\mathbf{x}))^{2}}$
$=\frac{\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}\left[(d^{2}(x_{i},Y_{i})-\theta(x_{i}))^{2}|x_{i}\right]}{(n\tau_{1}-\theta(\mathbf{x}))^{2}}$
$=\frac{\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},Y_{i})|x_{i})}{(n\tau_{1}-\theta(\mathbf{x}))^{2}}\leq\frac{M}{n\delta^{2}}=:\epsilon_{n}.$
Here given $M,\delta$, the sequence $\epsilon_{n}$ vanishes if we let $n\to\infty$.
Denote $\hat{\mu}_{\mathbf{Y}|\mathbf{X}}$ as a transition probability given by
$\hat{\mu}_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}=\mu_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}(\cdot|B_{\sqrt{n\tau_{1}}}(\mathbf{x})),\;\forall\mathbf{x}.$
That is, $\hat{\mu}_{\mathbf{Y}|\mathbf{X}}(A|\mathbf{x})=\frac{\mu_{\mathbf{Y}|\mathbf{X}}(A\cap B_{\sqrt{n\tau_{1}}}(\mathbf{x})|\mathbf{x})}{\mu_{\mathbf{Y}|\mathbf{X}}(B_{\sqrt{n\tau_{1}}}(\mathbf{x})|\mathbf{x})}$ for all measurable $A$. (Given measurable $A$, the measurability $\mathbf{x}\mapsto\hat{\mu}_{\mathbf{Y}|\mathbf{X}}(A|\mathbf{x})$ is well known; see e.g., [16, Theorem 18.3].) We can split $\mu_{\mathbf{Y}|\mathbf{X}}$ as a mixture:
$\mu_{\mathbf{Y}|\mathbf{X}}$ $=(1-\epsilon_{n})\hat{\mu}_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}+\epsilon_{n}\tilde{\mu}_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}$
for some transition probability $\tilde{\mu}_{\mathbf{Y}|\mathbf{X}}$. For the same input distribution $\mu_{\mathbf{X}}$, the output distributions of channels $\mu_{\mathbf{Y}|\mathbf{X}}$, $\hat{\mu}_{\mathbf{Y}|\mathbf{X}}$, and $\tilde{\mu}_{\mathbf{Y}|\mathbf{X}}$ are respectively denoted as $\mu_{\mathbf{Y}}$, $\hat{\mu}_{\mathbf{Y}}$, and $\tilde{\mu}_{\mathbf{Y}}$, which satisfy
$\mu_{\mathbf{Y}}=(1-\epsilon_{n})\hat{\mu}_{\mathbf{Y}}+\epsilon_{n}\tilde{\mu}_{\mathbf{Y}}.$
Denote $J\sim\mu_{J}:=\mathrm{Bern}(\epsilon_{n})$, and $\mu_{\mathbf{Y}|J=1}=\hat{\mu}_{\mathbf{Y}},\mu_{\mathbf{Y}|J=0}=\tilde{\mu}_{\mathbf{Y}}$. Then,
$\mu_{\mathbf{Y}}=\mu_{J}(1)\mu_{\mathbf{Y}|J=1}+\mu_{J}(0)\mu_{\mathbf{Y}|J=0}.$
Observe that
$D(\mu_{\mathbf{Y}|J}\|\nu^{\otimes n_{1}}|\mu_{J})$ $=(1-\epsilon_{n})D(\hat{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\epsilon_{n}D(\tilde{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}})$
$\geq(1-\epsilon_{n})D(\hat{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}}).$
On the other hand,
$D(\mu_{\mathbf{Y}|J}\|\nu^{\otimes n_{1}}|\mu_{J})$ $=D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+D(\mu_{\mathbf{Y}|J}\|\mu_{\mathbf{Y}}|\mu_{J})$
$\leq D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\ln 2,$
where the inequality follows since $D(\mu_{\mathbf{Y}|J}\|\mu_{\mathbf{Y}}|\mu_{J})=I_{\mu}(J;\mathbf{Y})\leq H_{\mu}(J)\leq\ln 2.$ Hence,
(3.20) $D(\hat{\mu}_{\mathbf{Y}}\|\nu^{\otimes n_{1}})\leq\frac{D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\ln 2}{1-\epsilon_{n}}.$
By choosing $Q_{\mathbf{Y}|\mathbf{X}}$ in (3.18) as the feasible solution $\hat{Q}_{\mathbf{Y}|\mathbf{X}}$, we then have that
(3.21) $\log\nu^{\otimes n_{1}}(B^{\sqrt{n\tau_{1}}})\leq\inf\frac{D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})+\log 2}{1-\epsilon_{n}},$
where the infimum is taken over all transition probabilities $\mu_{Y|X}$ satisfying (3.19).
Combining (3.16) and (3.21) yields that given $\delta>0$, for sufficiently large $n$ and sufficiently small $\epsilon$, it holds that
$-\log\nu^{\otimes n}(\tilde{A}_{0}^{\sqrt{n\tau_{0}}}\times B^{\sqrt{n\tau_{1}}})$
(3.22) $\leq\inf(1+\delta)\left(n_{0}D(\pi_{Y|W}\|\nu|\pi_{W})+D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})\right)+n\delta$
for some $(\pi_{XW},\mu_{\mathbf{X}})$ such that
$n_{0}D(\pi_{X|W}\|\nu_{F}|\pi_{W})+D(\mu_{\mathbf{X}}\|\nu^{\otimes n_{1}})\leq n\alpha^{\prime},\quad\alpha^{\prime}:=\alpha+2\delta,$
where the infimum is taken over all distributions $(\pi_{Y|XW},\mu_{Y|X})$ satisfying
$\mathbb{E}_{\pi}[d^{2}(X,Y)]$ $\leq\tau_{0}-\delta,$
$\mathbb{E}_{\pi}[\mathrm{Var}_{\pi}(d^{2}(X,Y)|X,W)]$ $\leq\kappa,$
and (3.19). Since $\tau_{0},\tau_{1}$ are arbitrary nonnegative numbers such that $\tau_{0}+\left(\sqrt{\tau_{1}}+2\sqrt{\delta}\right)^{2}\leq\tau$, optimizing the bound over all such $\tau_{0},\tau_{1}$ yields that the infimum can be taken over all distributions $(\pi_{Y|XW},\mu_{Y|X})$ satisfying
(3.23) $n_{0}\mathbb{E}_{\pi}[d^{2}(X,Y)]+\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},Y)|x_{i}]$ $\leq n\tau^{\prime}\;\mu_{\mathbf{X}}\textrm{-a.e.}$
(3.24) $\mathbb{E}_{\pi}[\mathrm{Var}_{\pi}(d^{2}(X,Y)|X,W)]$ $\leq\kappa,$
(3.25) $\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},Y)|x_{i})$ $\leq nM\;\mu_{\mathbf{X}}\textrm{-a.e.}$
where $\tau^{\prime}=\tau-5\delta-4\sqrt{\delta\tau}$.
3.5.4. Step 4: Heat Flow
We next construct a feasible solution to (3.22) by using heat flow. Let $\mu_{X_{t}|X=x}$ be the heat flow starting from the point $x$, which forms a transition probability. Let
$\pi_{WXX_{t}}=\pi_{WX}\mu_{X_{t}|X}$
and let $\pi_{X_{t}|W}$ the conditional probability induced by this joint probability. Let
(3.26) $\pi_{WXX_{t}}^{*}=\pi_{WX}\pi_{X_{t}|XW}^{*}$
where $\pi_{X_{t}|XW}^{*}$ is the optimal transport map attaining $\mathsf{W}_{2}(\pi_{X|W},\pi_{X_{t}|W}|\pi_{W})$. Let $\mu_{\mathbf{X}\mathbf{X}_{t}}=\mu_{\mathbf{X}}\mu_{X_{t}|X}^{\otimes n_{1}}$.
We choose $(\pi_{Y|XW},\mu_{Y|X})=(\pi_{X_{t}|XW}^{*},\mu_{X_{t}|X})$ and choose the time $t$ properly to satisfy (3.23). Let $\omega$ be a function given by
$\mathsf{W}_{2}^{2}(t):=n_{0}\mathsf{W}_{2}^{2}(\pi_{X|W},\pi_{X_{t}|W}|\pi_{W})+\mathsf{W}_{2}^{2}(\mu_{\mathbf{X}},\mu_{\mathbf{X}_{t}}),$
which is continuous in $t\geq 0$. By Remark 7, $\nu$ satisfies a log-Sobolev inequality with a constant $K>0$. Let $t_{0}=\frac{1}{2K}\ln\frac{\alpha^{\prime}}{\delta}$. Let $\tau^{\prime\prime}$ be a fixed number $\leq\tau$ which will be specified later. We choose the time $t=t_{0}$ if $\mathsf{W}_{2}(t_{0})\leq\sqrt{n\tau^{\prime\prime}}$, and choose $t$ as the smallest time $t_{1}$ such that $\mathsf{W}_{2}(t_{1})=\sqrt{n\tau^{\prime\prime}}$ if $\mathsf{W}_{2}(t_{0})>\sqrt{n\tau^{\prime\prime}}$. Hence, $\mathsf{W}_{2}(t)\leq\sqrt{n\tau^{\prime\prime}}$.
We next verify the two requirements—the expectation requirement and variance requirement in (3.23) and (3.24) for $(\mu_{X_{t}|X}\circ\mu_{X|W},\mu_{X_{t}|X})$, where $\mu_{X_{t}|X}\circ\mu_{X|W}$ denotes the resulting transition probability by cascading $\mu_{X|W}$ and $\mu_{X_{t}|X}$. We first focus on the expectation requirement. By the triangle inequality, we obtain for any $\mathbf{x},\mathbf{x}^{\prime}\in B$,
$\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\delta_{\mathbf{x}}\right)$ $\leq\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}}\right)+\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}},\delta_{\mathbf{x}}\right)$
(3.27) $\leq\sqrt{n\delta}+\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}},\delta_{\mathbf{x}}\right),$
where $\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}^{\prime}}\right)\leq\mathsf{W}_{2}\left(\delta_{\mathbf{x}},\delta_{\mathbf{x}^{\prime}}\right)\leq\sqrt{n\delta}$ follows by contraction of Wasserstein metric under $\mathrm{RCD}(0,\infty)$ given in [7, (6.1)] or [8, Corollary 3.18]. Taking expectation $\mathbb{E}_{\mathbf{X}^{\prime}\sim\mu_{\mathbf{X}}}$ and by Jensen’s inequality,
$\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\delta_{\mathbf{x}}\right)$ $\leq\mathbb{E}_{\mathbf{X}^{\prime}\sim\mu}[\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{X}^{\prime}},\delta_{\mathbf{x}}\right)]+\sqrt{n\delta}$
$\leq\sqrt{\mathbb{E}_{\mathbf{X}^{\prime}\sim\mu}[\mathsf{W}_{2}^{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{X}^{\prime}},\delta_{\mathbf{x}}\right)]}+\sqrt{n\delta}$
$=\sqrt{\mathbb{E}_{\mathbf{X}\sim\mu}\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})|X_{i}]}+\sqrt{n\delta}$
$=\sqrt{\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})]}+\sqrt{n\delta}$
$=\mathsf{W}_{2}\left(\delta_{\mathbf{x}},\mu_{\mathbf{X}_{t}}\right)+\sqrt{n\delta}.$
Therefore,
$\sqrt{\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})|x_{i}]}$ $=\mathsf{W}_{2}\left(\mu_{\mathbf{X}_{t}|\mathbf{X}=\mathbf{x}},\delta_{\mathbf{x}}\right)\leq\mathsf{W}_{2}\left(\delta_{\mathbf{x}},\mu_{\mathbf{X}_{t}}\right)+\sqrt{n\delta}$
$\leq\mathsf{W}_{2}\left(\mu_{\mathbf{X}},\mu_{\mathbf{X}_{t}}\right)+2\sqrt{n\delta},$
where the last inequality follows by the triangle inequality again. Moreover, $\mathbb{E}_{\pi^{*}}[d^{2}(X,Y)]=\mathsf{W}_{2}^{2}(\pi_{X|W},\pi_{X_{t}|W}|\pi_{W})$. Therefore, combining these with $\mathsf{W}_{2}(t)\leq\sqrt{n\tau^{\prime\prime}}$ yields
$n_{0}\mathbb{E}_{\pi}[d^{2}(X,Y)]+\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},Y)|x_{i}]\leq n(\tau^{\prime\prime}+4\delta+4\sqrt{\delta\tau})\;\mu_{\mathbf{X}}\textrm{-a.e.}$
Choosing $\tau^{\prime\prime}=\tau-9\delta-8\sqrt{\delta\tau}$ yields the expectation requirement in (3.23).
We next focus on the variance requirement. Under $\mathrm{CD}(0,\infty)$, the gradient flow (heat flow) $\rho_{t}$ (given in (2.4)) of Cheeger’s energy on the Hilbert space $L^{2}(\mathcal{X},\nu)$ is identical to the gradient flow $\mu_{t}$ of relative entropy on the Wasserstein metric space in the sense that $\mu_{t}=\rho_{t}\nu$ when $\mu_{0}=\rho_{0}\nu$ [6, Theorem 9.3]. In this case, denote $P_{t}$ as the Markov (heat) semigroup operator given by $P_{t}:\rho_{0}\mapsto\rho_{t}$. If in addition, the space satisfies $\mathrm{RCD}(0,\infty)$, then these two flows satisfy the equation $\frac{\mathrm{d}}{\mathrm{d}t}\rho_{t}=\Delta_{\mathcal{E}}\rho_{t}$ for all $t>0$ where $\Delta_{\mathcal{E}}$ is the infinitesimal generator of $P_{t}$ and $\rho_{t}=P_{t}\rho_{0}$ [7, Theorem 2.13 and Section 6.1], and moreover, in this case, Bakry-Émery curvature-dimension condition $\mathrm{BE}(0,\infty)$ is satisfied by the semigroup $P_{t}$ [8, Theorem 4.17].
By the local Poincaré inequalities in [8, Corollary 2.3, Definition 2.4, Theorem 4.17][10, Theorem 4.7.2], under $\mathrm{RCD}(0,\infty)$ (equivalently, $\mathrm{BE}(0,\infty)$), it holds that for every $f\in W^{1,2}(\mathcal{X},d,\nu)$ with the Sobolev space $W^{1,2}(\mathcal{X},d,\nu)$ defined around (2.3) and for $t>0$,
$P_{t}(f^{2})-(P_{t}f)^{2}\leq 2tP_{t}(|\nabla f|_{*}^{2}),\;\nu\textrm{-a.e. in }\mathcal{X},$
where $|\nabla f|_{*}$ is the minimal weak gradient defined below (2.3) and $f\mapsto|\nabla f|_{*}^{2}$ forms the Carré du champ $\Gamma$ induced by the Dirichlet form $\mathcal{E}=2\mathrm{Ch}$. That is, for every $f\in W^{1,2}(\mathcal{X},d,\nu)$,
(3.28) $\mathrm{Var}_{\mu}(f(X_{t})|X=x)\leq 2t\mathbb{E}_{\mu}[|\nabla f|_{*}^{2}(X_{t})|X=x],\;\nu\textrm{-a.e. }x\in\mathcal{X}.$
Let $f(y)=d^{2}(x,y)$ and for $N>0$, let $f_{N}(y)=\min\{f(y),N^{2}\}=\left(\min\{d(x,y),N\}\right)^{2}$. Then, $|\nabla f_{N}|\leq 2\min\{d(x,y),N\}$ and hence $f_{N}\in W^{1,2}(\mathcal{X},d,\nu)$. Substituting this function into (3.28) and noting that $|\nabla f_{N}|_{*}\leq|\nabla f_{N}|\leq 2d(x,y)$ $\nu$-a.e. $y$, we obtain that
$\mathrm{Var}_{\mu}(f_{N}(X_{t})|X=x)\leq 8t\mathbb{E}_{\mu}[d^{2}(x,X_{t})|X=x],\;\nu\textrm{-a.e. }x\in\mathcal{X}.$
Letting $N\to\infty$ and by the monotone convergence theorem, for $\nu\textrm{-a.e. }x\in\mathcal{X}$,
(3.29) $\mathrm{Var}_{\mu}(d^{2}(x,X_{t})|X=x)\leq 8t\mathbb{E}_{\mu}[d^{2}(x,X_{t})|X=x].$
Using the triangle inequality, for any $x_{0},x\in F$,
$\sqrt{\mathbb{E}_{\mu}[d^{2}(x_{0},X_{t})|X=x]}$ $=\mathsf{W}_{2}(P_{t}\delta_{x},\delta_{x_{0}})$
$\leq\mathsf{W}_{2}(P_{t}\delta_{x},\nu)+\mathsf{W}_{2}(\nu,\delta_{x_{0}})$
$\leq\mathsf{W}_{2}(\delta_{x},\nu)+\mathsf{W}_{2}(\nu,\delta_{x_{0}})$
$\leq\mathsf{W}_{2}(\delta_{x},\delta_{x_{0}})+2\mathsf{W}_{2}(\delta_{x_{0}},\nu)$
(3.30) $\leq\mathrm{diam}(F)+2\mathsf{W}_{2}(\delta_{x_{0}},\nu)=:C,$
where the second inequality follows by $\mathsf{W}_{2}(P_{t}\delta_{x},\nu)\leq\mathsf{W}_{2}(\delta_{x},\nu)$, contraction of Wasserstein metric in [7, (6.1)]. Noting that $C<\infty$, $\mathbb{E}_{\mu}[d^{2}(x_{0},X_{t})|X=x]$ is upper bounded by the finite number $C$ uniformly for all $x\in F$. We then obtain from (3.29) that
(3.31) $\mathbb{E}_{\mu}[d^{4}(x_{0},X_{t})|X=x]\leq C^{4}+8tC^{2}.$
Moreover, by triangle inequality and Minkowski inequality, for any $x_{0}\in F$,
$\mathbb{E}_{\pi^{*}}[d^{4}(X,X_{t})]^{1/4}$ $\leq\mathbb{E}_{\pi^{*}}[d^{4}(x_{0},X_{t})]^{1/4}+\mathbb{E}_{\pi^{*}}[d^{4}(X,x_{0})]^{1/4}$
$\leq\mathbb{E}_{\pi}[d^{4}(x_{0},X_{t})]^{1/4}+\mathrm{diam}(F),$
where the last line follows since $\pi_{X_{t}}^{*}=\pi_{X_{t}}$, according to (3.26), and $d(x,x_{0})\leq\mathrm{diam}(F)$ for all $x\in F$. By (3.31), $\mathbb{E}_{\pi}[d^{4}(x_{0},X_{t})]\leq C^{4}+8tC^{2}.$ Thus,
$\mathbb{E}_{\pi^{*}}[d^{4}(X,X_{t})]^{1/4}\leq(C^{4}+8tC^{2})^{1/4}+\mathrm{diam}(F).$
Note that we have truncated the time $t$ up to the finite value $t_{0}$, implying $t\leq t_{0}$. Hence, the first variance requirement is satisfied. This gives the reason why we truncate $t$.
We next consider the second variance requirement. Substituting $x=x_{i}$ into (3.29) and taking summation over $i$, we obtain that for $\mu_{\mathbf{X}}$-a.e. $\mathbf{x}$,
$\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},X_{i,t})|X_{i}=x_{i})\leq 8t\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},X_{i,t})|X_{i}=x_{i}]\leq 8tn(\tau-\delta),$
where the last inequality above follows by the expectation condition in (3.23). Since $t\leq t_{0}$, the second variance requirement is satisfied.
Since all requirements are satisfied by the transition probability induced by heat flow, we obtain a feasible solution to (3.22).
3.5.5. Step 5: Final Bound
We next analyze the relative entropy induced by the heat flow. Let
$D(t):=n_{0}D(\pi_{X_{t}|W}\|\nu|\pi_{W})+D(\mu_{\mathbf{X}_{t}}\|\nu^{\otimes n_{1}}).$
We first show that $D(t)\leq n\delta$ if $t=t_{0}$. Since $\nu$ admits a log-Sobolev inequality with constant $K$, by exponential decay in entropy [10, Theorem 5.2.1], we have that in this case,
(3.32) $D(t_{0})\leq e^{-2Kt_{0}}D(0)\leq e^{-2Kt_{0}}n\alpha=n\delta,$
where the last equality is due to the choice of $t_{0}$.
We next upper bound the relative entropy for $t=t_{1}$. By the chain rule for relative entropy and Fisher information and the fact that conditioning increases relative entropy, the following kind of tensorization property can be proven.
Theorem 9.
[59] Assume that $\nu$ admits a $\theta$-nonlinear log-Sobolev inequality for some convex function $\theta$. Then, $\nu^{\otimes n}$ admits a $\theta_{n}$-nonlinear log-Sobolev inequality where $\theta_{n}(t)=n\theta(t/n),\forall t\geq 0$.
By this theorem and Theorem 6, we obtain that
(3.33) $\sqrt{n\tau^{\prime\prime}}=\mathsf{W}_{2}(t)$ $\leq\int_{D(t)}^{D(0)}\frac{1}{\sqrt{n\theta(r/n)}}{\text{\rm d}}r=\sqrt{n}\int_{D(t)/n}^{D(0)/n}\frac{1}{\sqrt{\theta(t)}}{\text{\rm d}}t=\sqrt{n}\Upsilon\left(D(t)/n,\alpha\right).$
Substituting (3.32) and (3.33) into (3.22) and letting $\delta\to 0$ yield that $\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\bar{\psi}(\alpha,\tau).$
3.6. Interpretation from Concentration of Measure
The isoperimetric problem with thick boundary admits a natural interpretation from the perspective of concentration of measure. Concentration of measure principle states that for any set of not too small probability, slightly enlarging this set will always have probability close to one. Formally, let $r(a)$ be the minimum number such that the $r$-enlargement of any set of measure $a$ will always have probability close to one. Denote $a=e^{-n\alpha}$ and $r^{2}(a)=n\tau(\alpha)$. By our results, for sufficiently large $n$, $\tau(\alpha)$ is the minimum $\tau$ such that $\psi(\alpha,\tau)=0$. Solving it yields $\tau\geq\mathsf{W}_{2}(\pi_{X|W},\nu|\pi_{W})$ for all $\pi_{XW}$ such that $D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha$. Hence, $\tau(\alpha)$ satisfies
$\tau(\alpha)$ $=\sup_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}\mathsf{W}_{2}(\pi_{X|W},\nu|\pi_{W}),$
which is the upper concave envelope of $\tau_{0}$ given below
$\tau_{0}(\alpha):$ $=\sup_{\pi:D(\pi\|\nu)\leq\alpha}\mathsf{W}_{2}(\pi,\nu).$
The curves $\tau$ and $\tau_{0}$ respectively characterize the best convex and nonconvex tradeoffs between the relative entropy and the Wasserstein metric. Thus, they determine the optimal nonlinear transport inequalities.
Recall that $K_{C}$ is the optimal transport constant in the following inequality: $\mathsf{W}_{2}(\pi,\nu)\leq\sqrt{\frac{2}{K_{C}}D(\pi\|\nu)},\;\forall\pi.$ Hence, $\tau(\alpha)\leq\frac{2\alpha}{K_{C}},\;\forall\alpha>0,$ and the constant $K_{C}$ is optimal. This implies that for sufficiently large $n$, enlarging any set of measure $\exp(-\frac{K_{C}r^{2}}{2})$ by the distance $r$ will always make it have probability close to one, and moreover, the constant $K_{C}$ is the optimal one satisfying this property. In contrast, another (but indeed equivalent) interpretation of $K_{C}$ in the concentration of measure is that for all $A$ such that $\nu_{n}(A)=1/2$, it always holds that $\nu_{n}((A^{r})^{c})\leq\exp(-\frac{K_{C}r^{2}}{2}),\;\forall r>0$. Here, the constant $K_{C}$ cannot be improved either.
4. From Variant Isoperimetry to Standard Isoperimetry
We now turn back to the standard isoperimetric problem in the Riemannian manifold, in which the boundary is the infinitely thin shell. We now connect the isoperimetric problem with thick shell boundary considered in the last section to this setting.
Recall that $\mathcal{M}$ is a $k$-dimensional Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol},V\in C^{2}(\mathcal{M})$ and the induced geodesic distance $d$. Recall the boundary measure defined in (1.2). The boundary measure can be alternatively expressed as
$\nu_{n}^{+}(A)$ $=\liminf_{r\downarrow 0}\frac{\nu_{n}(A^{r})-\nu_{n}(A)}{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}\cdot\frac{\ln\nu_{n}(A^{r})-\ln\nu_{n}(A)}{r}$
(4.1) $=\nu_{n}(A)\liminf_{r\downarrow 0}\frac{\ln[\nu_{n}(A^{r})/\nu_{n}(A)]}{r}.$
However, to apply Theorem 8, we need keep $r$ to be large (in order of $\sqrt{n}$), and hence, we need remove $\liminf_{r\downarrow 0}$ in the last line. To this end, we need some nice properties of isoperimetric minimizers from geometric measure theory.
Here we adopt the notation used in [52]. For a smooth hypersurface $S$, the second fundamental form is by $\varPi_{S,p}^{\sigma}(u,v)=g(D_{u}v,v)$ for $u,v\in T_{p}S$, where $D$ is the covariant derivative. According to this definition, the second fundamental form of the sphere in Euclidean space with respect to the outer normal $\sigma$ is positive definite. The total curvature $H_{S}^{\sigma}(p)$ of $S$ at a point $p$ is the trace of the second fundamental form $\varPi_{S,p}^{\sigma}$. (The mean curvature is obtained if it is divided by the dimension of $S$.)
Definition 3.
The $\nu$-total curvature of a smooth hypersurface $S$ at a point $p\in S$ with respect to the unit normal vector $\sigma$, denoted $H_{S,\nu}^{\sigma}(p)$, is defined as
$H_{S,\nu}^{\sigma}(p):=H_{S}^{\sigma}(p)-\langle\nabla V(p),\sigma(p)\rangle.$
Bayle [15] and Morgan [55, 56] showed that isoperimetric minimizers always exist and the isoperimetric profile is concave. See an explicit explanation in [52, p. 216].
Theorem 10 (Morgan [54, 55, 56] and Bayle [15]).
Under Condition 2, the following hold.
  1. (1)
    For any $a\in(0,1)$, there always exists an open isoperimetric minimizer $A$ in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a$ such that: (a) its boundary $\partial A$ consists of disjoint regular part and singular part; (b) the regular part is a $C^{2}$-smooth hypersurface (since $V\in C^{2}(\mathcal{M})$) which has a constant $\nu_{n}$-total curvature, denoted by $H_{\nu_{n}}(A)$; (c) the singular part has dimension strictly lower than that of the regular part, more precisely, lower than or equal to the dimension of the regular part minus $7$.
  2. (2)
    The isoperimetric profile $I_{n}$ is a concave function on $[0,1]$.
  3. (3)
    For the isoperimetric minimizer given in Statement 1, the constant $\nu_{n}$-total curvature $H_{\nu_{n}}(A)$ of the regular part of its boundary satisfies
    $\lim_{\epsilon\downarrow 0}\frac{I_{n}(a+\epsilon)-I_{n}(a)}{\epsilon}\leq H_{\nu_{n}}(A)\leq\lim_{\epsilon\downarrow 0}\frac{I_{n}(a)-I_{n}(a-\epsilon)}{\epsilon}.$
    In particular, if $I_{n}$ is differentiable at $a$, then $H_{\nu_{n}}(A)=I_{n}^{\prime}(a)$.
Theorem 11 (Generalized Heintze-Karcher due to Morgan [55, Theorem 2, Remark 3]).
Assume that for $\kappa\in\mathbb{R}$,
$\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq\kappa g.$
Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 10. Let $S$ denote the regular part of $\partial A$, and let $H_{\nu_{n}}(A)$ denote the constant $\nu_{n}$-total curvature of $S$ with respect to the outer unit normal vector field $\sigma$ on $S$. Then for any $r>0$,
$\nu_{n}(A^{r})-\nu_{n}(A)\leq\nu_{n}^{+}(A)\int_{0}^{r}\exp\left(H_{\nu_{n}}(A)t-\frac{\kappa}{2}t^{2}\right){\text{\rm d}}t.$
Note that the sign of the total curvature $H_{S}^{\sigma}(p)$ here is opposite to the one in [55, Theorem 2, Remark 3]. Based on the two theorems above, we build a bridge between the standard isoperimetric problem and the isoperimetric problem with thick boundary.
Theorem 12.
Assume that Condition 1 holds. Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 10. Then, for any $r>0$,
$\frac{I_{n}(a)}{a}=\frac{\nu_{n}^{+}(A)}{\nu_{n}(A)}\geq\frac{1}{r}\ln\frac{\nu_{n}(A^{r})}{\nu_{n}(A)}\geq\frac{1}{r}\ln\frac{\Gamma_{n}(a,r)}{a}.$
Proof.
Let $S$ denote the regular part of $\partial A$, and let $H_{\nu_{n}}(A)$ denote the constant $\nu_{n}$-total curvature of $S$ with respect to the outer unit normal vector field $\sigma$ on $S$. Let
$P(r)$ $:=\nu_{n}^{+}(A)\exp\left(H_{\nu_{n}}(A)r\right),$
$V(r)$ $:=\nu_{n}(A)+\int_{0}^{r}P(t){\text{\rm d}}t.$
Define g(r):=V(r)P(0)P(r)V(0),g(r):=V(r)P(0)-P(r)V(0), whose derivative satisfies
(4.2) $g^{\prime}(r)=P(r)P(0)-P^{\prime}(r)V(0)=P(r)P(0)-H_{\nu_{n}}(A)P(r)V(0).$
From Theorem 10, it is known that
$H_{\nu_{n}}(A)\leq\lim_{\epsilon\downarrow 0}\frac{I_{n}(a)-I_{n}(a-\epsilon)}{\epsilon},$
where $a=\nu_{n}(A)$. Since $I_{n}$ is concave,
$\lim_{\epsilon\downarrow 0}\frac{I_{n}(a)-I_{n}(a-\epsilon)}{\epsilon}\leq\frac{I_{n}(a)}{a}=\frac{P(0)}{V(0)}.$
Hence, $H_{\nu_{n}}(A)\leq\frac{P(0)}{V(0)},$ which, combined with (4.2), implies $g^{\prime}(r)\geq 0.$
Observe that $g(0)=0$. So, for all $r\geq 0$, it holds that $g(r)\geq 0,$ i.e., $\frac{P(r)}{V(r)}\leq\frac{P(0)}{V(0)}.$ Noting that $P(r)=V^{\prime}(r)$, we then have that $\ln\frac{V(r)}{V(0)}\leq\frac{P(0)}{V(0)}r.$ From Theorem 11, it is known that $V(r)\geq\nu_{n}(A^{r})$. We hence conclude that
$\frac{\nu^{+}(A)}{\nu(A)}=\frac{P(0)}{V(0)}\geq\frac{1}{r}\ln\frac{V(r)}{V(0)}\geq\frac{1}{r}\ln\frac{\nu_{n}(A^{r})}{\nu_{n}(A)}.$
5. Putting Everything Together: From Log-Sobolev to Standard Isoperimetry
In this section, we combine all the tools built in Sections 24 to prove the lower bound in our large deviations theorem, i.e., to deduce standard isoperimery from nonlinear log-Sobolev inequalities. As a consequence of Theorem 8 and Theorem 12, we can lower bound the asymptotics of the isoperimetric profile $I_{n}$ in terms of $\breve{\Theta}$. Recall that $\breve{\Theta}$ defined in (1.7) is the function characterizing the optimal tradeoff in the nonlinear log-Sobolev inequality. We obtain the lower bound part of Theorem 1. Recall $\underline{\Lambda}$ defined in (1.5).
Theorem 13 (From Nonlinear Log-Sobolev to Standard Isoperimetry).
Under Conditions 1 and 2, for any $\alpha>0$,
(5.1) $\underline{\Lambda}(\alpha)$ $\geq\sqrt{\breve{\Theta}(\alpha)}.$
Proof.
Let $A_{n}$ be an isoperimetric minimizer of size $e^{-n\alpha}$ in the $n$-dimensional space. From Theorem 12 with $r=\sqrt{n\tau}$, we see that
$\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}\geq\frac{\alpha-E_{n}(\alpha,\tau)}{\sqrt{\tau}}.$
Letting $n\to\infty$, from Theorem 8 we obtain for any $\tau>0$, $\underline{\Lambda}(\alpha)\geq\frac{\alpha-\bar{\psi}(\alpha,\tau)}{\sqrt{\tau}}.$ Letting $\tau\to 0$, we obtain $\underline{\Lambda}(\alpha)\geq\limsup_{\tau\downarrow 0}\frac{\alpha-\Upsilon_{\alpha}^{-1}(\sqrt{\tau})}{\sqrt{\tau}}$.
Denote $t=\sqrt{\tau}$ and $s(\alpha,t)=\Upsilon_{\alpha}^{-1}(t)$. Then,
$t=\int_{s(\alpha,t)}^{\alpha}\frac{1}{\sqrt{\breve{\Theta}(r)}}{\text{\rm d}}r.$
Note that $s(\alpha,0)=\alpha$. Taking derivative w.r.t. $t$ at $t=0$ yields $\partial_{t}s(\alpha,t)\big|_{t=0}=\sqrt{\breve{\Theta}(\alpha)}$. Therefore, $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$
6. The Other Direction: From Standard Isoperimetry to Log-Sobolev
We now prove the other direction (i.e., the upper bound) in Theorem 1. Recall $\bar{\Lambda}$ defined around (1.5).
Theorem 14 (From Isoperimetry to Nonlinear Log-Sobolev).
Under Condition 1, for any $\alpha>0$,
(6.1) $\bar{\Lambda}(\alpha)$ $\leq\sqrt{\breve{\Theta}(\alpha)}.$
We provide two distinct proofs for this theorem. One is to use the tools of the variant isoperimetry and the famous HWI inequality. The other one is to apply the coarea formula and the large deviations theory.
First Proof of Theorem (14).
We first prove the first inequality in (6.1). Let $A=A_{n}$ be a set approximately attaining $\Gamma_{n}(e^{-n\alpha},n\tau)$ (given in (3.1)) within a factor $1-\epsilon$ where $\epsilon>0$, i.e., $A$ has probability $e^{-n\alpha}$ and $\nu_{n}(A^{\sqrt{n\tau}})\geq(1-\epsilon)\Gamma_{n}(e^{-n\alpha},n\tau)$. Let $r:=\sqrt{n\tau}$. Since $f(t):=\ln\nu_{n}(A^{t})$ is monotone, it is differentiable almost everywhere on $(0,r)$. As a basic property of monotone functions, $f^{\prime}$ is integrable over $[0,r]$ and $\int_{0}^{r}f^{\prime}(t){\text{\rm d}}t\leq f(r)-f(0)$ (see e.g., [66, Corollary 4 on p.113]), which implies that there is some $r_{0}\in(0,r)$ such that $\frac{f(r)-f(0)}{r}\geq f^{\prime}(r_{0}),$i.e.,
$\frac{1}{r}\ln[\nu_{n}(A^{r})/\nu_{n}(A)]\geq\frac{\nu_{n}^{+}(A^{r_{0}})}{\nu_{n}(A^{r_{0}})}.$
Since $\nu_{n}^{+}(A^{r_{0}})\geq I_{n}(\nu_{n}(A^{r_{0}}))$ and $I_{n}(a)/a$ is non-increasing in $a$, it holds that
$\frac{1}{r}\ln[\nu_{n}(A^{r})/\nu_{n}(A)]\geq\frac{I_{n}(\nu_{n}(A^{r_{0}}))}{\nu_{n}(A^{r_{0}})}\geq\frac{I_{n}(\nu_{n}(A^{r}))}{\nu_{n}(A^{r})}.$
Noting that $r=\sqrt{n\tau}$ and letting $n\to\infty$, by Statement 2 in Theorem 7, we obtain for any $\tau>0$,
(6.2) $\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$ $\geq\lim_{n\to\infty}\frac{\ln[\nu_{n}(A_{n}^{\sqrt{n\tau}})/\nu_{n}(A_{n})]}{n\sqrt{\tau}}\geq\limsup_{n\to\infty}\frac{I_{n}(a_{n,\tau})}{a_{n,\tau}\sqrt{n}},$
where $a_{n,\tau}=\nu_{n}(A_{n}^{\sqrt{n\tau}})$ and $\psi$ is defined in (3.12).
Since $A$ approximately attains $\Gamma_{n}(e^{-n\alpha},n\tau)$ within a factor $1-\epsilon$, by Statement 2 in Theorem 7 again, it holds that $\liminf_{n\to\infty}-\frac{1}{n}\ln a_{n,\tau}\geq\psi(\alpha,\tau).$ Hence, (6.2) implies that for any $\delta>0$, $\bar{\Lambda}(\psi(\alpha,\tau)-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$ Taking $\tau\downarrow 0$ yields
(6.3) $\limsup_{\tau\downarrow 0}\bar{\Lambda}(\psi(\alpha,\tau)-\delta)\leq\xi(\alpha)$ $:=\limsup_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$
We can assume that $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$, since otherwise, $\xi(\alpha)=\infty$ and nothing is needed to prove. Hence, given any $\delta>0$, we assume that $\psi(\alpha,\tau)>\alpha-\delta$ for all sufficiently small $\tau$, which, combined with the inequality above, implies
(6.4) $\bar{\Lambda}(\alpha-2\delta)\leq\xi(\alpha).$
We next prove $\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$. Substituting the expression of $\psi(\alpha,\tau)$ into the bound above yields that
$\xi(\alpha)$ $=\limsup_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$
$\leq\lim_{\tau\downarrow 0}\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}$
$\qquad\sup_{\pi_{Y|W}:\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\sqrt{\tau}}\frac{\alpha-D(\pi_{Y|W}\|\nu|\pi_{W})}{\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})}$
$\leq\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})=\alpha}\lim_{\tau\downarrow 0}$
$\qquad\sup_{\pi_{Y|W}:\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\sqrt{\tau}}\frac{D(\pi_{X|W}\|\nu|\pi_{W})-D(\pi_{Y|W}\|\nu|\pi_{W})}{\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})},$
where in the last line, we swap the infimization and the limit, and we restrict the inequality constraint in the infimization to be an equality.
Denote $\alpha_{i}=D(\pi_{X|W=i}\|\nu),\beta_{i}=D(\pi_{Y|W=i}\|\nu)$, and $\gamma_{i}=\mathsf{W}_{2}(\pi_{X|W=i},\pi_{Y|W=i})$, $i\in\{0,1\}$. Then, by the Cauchy–Schwarz inequality,
$D(\pi_{X|W}\|\nu|\pi_{W})-D(\pi_{Y|W}\|\nu|\pi_{W})=\mathbb{E}_{\pi}[\alpha_{W}-\beta_{W}]$
$\qquad\leq\sqrt{\mathbb{E}_{\pi}[\left(\frac{\alpha_{W}-\beta_{W}}{\gamma_{W}}\right)^{2}]\mathbb{E}_{\pi}[\gamma_{W}^{2}]}.$
So,
(6.5) $\xi(\alpha)$ $\leq\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})=\alpha}\lim_{\tau\downarrow 0}\sup_{\pi_{Y|W}:\mathbb{E}_{\pi}[\gamma_{W}^{2}]\leq\tau}\sqrt{\mathbb{E}_{\pi}[\left(\frac{\alpha_{W}-\beta_{W}}{\gamma_{W}}\right)^{2}]}.$
Since $\pi_{Y|W=i}=\pi_{X|W=i}$ is always a feasible solution to the supremum above, we have $\beta_{i}\leq\alpha_{i},i\in\{0,1\}$.
Define the descending slope of the relative entropy w.r.t. the Wasserstein metric as
$|\partial^{-}D|(\mu\|\nu)$ $:=\limsup_{\pi\to\mu}\frac{[D(\mu\|\nu)-D(\pi\|\nu)]^{+}}{\mathsf{W}_{2}(\pi,\mu)}$
$=\lim_{\tau\downarrow 0}\sup_{\pi:0<\mathsf{W}_{2}(\pi,\mu)\leq\tau}\frac{[D(\mu\|\nu)-D(\pi\|\nu)]^{+}}{\mathsf{W}_{2}(\pi,\mu)}.$
Using this notation, we rewrite (6.5) as
$\xi(\alpha)$ $\leq\inf_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})=\alpha}\sqrt{\mathbb{E}_{W\sim\pi_{W}}[\left(|\partial^{-}D|(\pi_{X|W}\|\nu)\right)^{2}]}.$
We now need the HWI inequality for the Riemannian manifold setting [73, Corollary 20.13] [58].
Theorem 15 (HWI Inequality [73, Corollary 20.13]).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that Condition 1 holds. Then, for any two probability measures $\mu_{0}=\rho_{0}\nu,\mu_{1}=\rho_{1}\nu$ in $\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\mu_{1}\|\nu)<\infty$ and $\rho_{0}$ is Lipschitz, it holds that
$D(\mu_{0}\|\nu)-D(\mu_{1}\|\nu)\leq\mathsf{W}_{2}(\mu_{0},\mu_{1})\sqrt{I(\mu_{0}\|\nu)}.$
This theorem implies $|\partial^{-}D|(\mu\|\nu)\leq\sqrt{I(\mu\|\nu)}.$ In fact, this is an equality even for Polish metric probability measure space satisfying $\mathrm{\mathrm{C}D}(K,\infty)$; see [6, Theorem 9.3]. We restrict $\pi_{X|W=i},i\in\{0,1\}$ to be in $\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$. Therefore,
$\xi(\alpha)$ $\leq\inf_{\begin{subarray}{c}\pi_{XW}:\pi_{X|W=i}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),\forall i,\\ D(\pi_{X|W}\|\nu|\pi_{W})=\alpha\end{subarray}}\sqrt{I(\pi_{X|W}\|\nu|\pi_{W})}=\sqrt{\breve{\Theta}(\alpha)}.$
Combining this with (6.4) yields that $\bar{\Lambda}(\alpha-2\delta)\leq\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$ That is, for $\alpha,\delta>0$, $\bar{\Lambda}(\alpha)\leq\xi(\alpha+\delta)\leq\sqrt{\breve{\Theta}(\alpha+\delta)}.$ Since $\breve{\Theta}(t)$ is continuous in $t>0$, letting $\delta\downarrow 0$ yields $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$
Second Proof of Theorem (14).
By an approximation argument, we may assume that the probability measures $\mu$ in the nonlinear log-Sobolev inequality have smooth density $f:={\text{\rm d}}\mu/{\text{\rm d}}\nu$. Obviously, $\int f\,d\nu=1$. Let $f^{\otimes n}$ be the $n$-fold product of $f$ with itself, which satisfies ${\text{\rm d}}\mu_{n}/{\text{\rm d}}\nu_{n}=f^{\otimes n}$ with $\mu_{n}=\mu^{\otimes n}$. Then, the coarea formula (see e.g., [25, Exercise III.12]) states:
(6.6) $\int|\nabla f^{\otimes n}|\,d\nu_{n}=\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$
where $\varphi_{n}(\mathbf{x})=e^{-\sum_{i=1}^{n}V(x_{i})}$ be the density function of $\nu_{n}$ w.r.t. the volume form on the manifold $\mathcal{M}^{n}$, $C_{s}=\{\mathbf{x}\in\mathcal{M}^{n}:f^{\otimes n}(\mathbf{x})=s\}$ be the level set of the function $f^{\otimes n}$, ${\text{\rm d}}\mathrm{\mathrm{area}}$ be the induced volume form (Hausdorff measure of dimension equal to the dimension of $\mathcal{M}^{n}$ minus $1$) on the level set $C_{s}$, and $|\nabla f^{\otimes n}(\mathbf{x})|$ be the norm of the gradient of $f^{\otimes n}$ at the point $\mathbf{x}\in\mathcal{M}^{n}$. For brevity, we denote $D:=D(\mu\|\nu)$.
Observe that
$\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s\geq\int_{e^{n(D-\delta)}}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$
(6.7) $\qquad=\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}^{+}\{f^{\otimes n}\geq s\}{\text{\rm d}}s\geq\eta_{n}\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s,$
where
$\eta_{n}:=\inf_{\begin{subarray}{c}s\geq e^{n(D-\delta)}:\\ \nu_{n}\{f^{\otimes n}\geq s\}>0\end{subarray}}\frac{\nu_{n}^{+}\{f^{\otimes n}\geq s\}}{\nu_{n}\{f^{\otimes n}\geq s\}}.$
Note that
$\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s=\int_{0}^{\infty}\int 1\{f^{\otimes n}\geq s,\,s\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}{\text{\rm d}}s$
(6.8) $\qquad=\int(f^{\otimes n}-e^{n(D-\delta)})\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}=\gamma_{1,n}-\gamma_{2,n},$
where
$\gamma_{1,n}$ $:=\int f^{\otimes n}\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}=\mu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\},$
and
$\gamma_{2,n}:=e^{n(D-\delta)}\nu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\}.$
By the law of large numbers, $\gamma_{1,n}\to 1$ as $n\to\infty$. By the large deviations theory (or the asymptotic equipartition property in information theory),
$\limsup_{n\to\infty}-\frac{1}{n}\ln\gamma_{2,n}=\sup_{t>0}t\left(D_{1-t}(\mu\|\nu)-(D(\mu\|\nu)-\delta)\right)>0,$
where $D_{1-t}(\mu\|\nu)=-\frac{1}{t}\ln\int f^{1-t}d\nu,\;f={\text{\rm d}}\mu/{\text{\rm d}}\nu$ is the (1t)(1-t)-Rényi divergence of $\mu$ w.r.t. $\nu$. Therefore, substituting these into (6.8) yields $\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s\to 1,\textrm{ as }n\to\infty,$ which, combined with (6.7), implies
$\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$ $\geq(1+o(1))\eta_{n}.$
We now lower bound $\eta_{n}$. Observe that
$\eta_{n}$ $\geq\inf_{\begin{subarray}{c}s\geq e^{n(D-\delta)}:\\ \nu_{n}\{f^{\otimes n}\geq s\}>0\end{subarray}}\frac{I_{n}(\nu_{n}\{f^{\otimes n}\geq s\})}{\nu_{n}\{f^{\otimes n}\geq s\}}=\frac{I_{n}(a_{n})}{a_{n}},$
where $a_{n}:=\nu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\}$, and the equality follows since $I_{n}(a)/a$ is non-increasing in $a$. Hence,
$\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$ $\geq(1+o(1))\frac{I_{n}(a_{n})}{a_{n}}.$
One can see that
$a_{n}$ $=\int_{\frac{{\text{\rm d}}\mu_{n}}{{\text{\rm d}}\nu_{n}}\geq e^{n(D-\delta)}}{\text{\rm d}}\nu_{n}\leq\int_{\frac{{\text{\rm d}}\mu_{n}}{{\text{\rm d}}\nu_{n}}\geq e^{n(D-\delta)}}e^{-n(D-\delta)}{\text{\rm d}}\mu_{n}\leq e^{-n(D-\delta)}=:b_{n}.$
By again the fact that $I_{n}(a)/a$ is non-increasing in $a$, we have
$\limsup_{n\to\infty}\frac{I_{n}(a_{n})}{a_{n}\sqrt{n}}\geq\limsup_{n\to\infty}\frac{I_{n}(b_{n})}{b_{n}\sqrt{n}}=\bar{\Lambda}(D-\delta).$
Therefore, for any $\delta>0$,
(6.9) $\limsup_{n\to\infty}\frac{1}{\sqrt{n}}\int_{0}^{\infty}\left(\int_{C_{s}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$ $\geq\bar{\Lambda}(D(\mu\|\nu)-\delta).$
On the other hand, by the Cauchy–Schwarz inequality,
(6.10) $\int|\nabla f^{\otimes n}|\,{\text{\rm d}}\nu_{n}\leq\sqrt{\int f^{\otimes n}\,{\text{\rm d}}\nu_{n}\,\cdot\int\frac{|\nabla f^{\otimes n}|^{2}}{f^{\otimes n}}\,{\text{\rm d}}\nu_{n}}=\sqrt{nI(\mu\|\nu)}.$
Combining (6.6), (6.9), and (6.10) yields that $\sqrt{I(\mu\|\nu)}\geq\bar{\Lambda}(D(\mu\|\nu)-\delta),\,\forall\mu.$ Noting that $\bar{\Lambda}$ is non-decreasing, we have $\sqrt{\Theta(\alpha)}\geq\bar{\Lambda}(\alpha-\delta).$ That is,
(6.11) $\bar{\Lambda}(\alpha)\leq\sqrt{\Theta(\alpha+\delta)}.$
By substitution $\nu\leftarrow\nu^{\otimes k}$, we define
$\bar{\Lambda}_{k}(\alpha):=\limsup_{n\to\infty}\frac{I_{kn}(e^{-kn\alpha})}{e^{-kn\alpha}\sqrt{n}}.$
Noting that both $n\mapsto I_{n}(a)$ and $a\mapsto I_{n}(a)/a$ are non-increasing, it holds that for $kn\leq n^{\prime}<k(n+1)$,
$\frac{I_{k(n+1)}(e^{-kn\alpha})}{e^{-kn\alpha}\sqrt{k(n+1)}}\leq\frac{I_{n^{\prime}}(e^{-n^{\prime}\alpha})}{e^{-n^{\prime}\alpha}\sqrt{n^{\prime}}}\leq\frac{I_{kn}(e^{-k(n+1)\alpha})}{e^{-k(n+1)\alpha}\sqrt{kn}}.$
Taking $n\to\infty$ yields that for any $\delta>0$, $\frac{\bar{\Lambda}_{k}(\alpha-\delta)}{\sqrt{k}}\leq\bar{\Lambda}(\alpha)\leq\frac{\bar{\Lambda}_{k}(\alpha+\delta)}{\sqrt{k}}.$ Applying (6.11) to $(\mathcal{M}^{k},\nu^{\otimes k})$ yields that for all $k$,
(6.12) $\bar{\Lambda}(\alpha)\leq\frac{\bar{\Lambda}_{k}(\alpha+\delta)}{\sqrt{k}}\leq\sqrt{\frac{\Theta_{k}(k(\alpha+2\delta))}{k}},$
where $\Theta_{k}$ is defined for $(\mathcal{M}^{k},\nu^{\otimes k})$ and given by
$\Theta_{k}(\alpha):=\inf_{\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu^{\otimes k})\geq\alpha}I(\mu\|\nu^{\otimes k}).$
We now claim that
(6.13) $\lim_{k\to\infty}\frac{\Theta_{k}(k\beta)}{k}=\breve{\Theta}(\beta).$
This is because, on one hand, by the tensorization property of nonlinear log-Sobolev inequality, for all $k$ and $\beta\geq 0$, $\frac{1}{k}\Theta_{k}(k\beta)\geq\breve{\Theta}(\beta);$ on the other hand, by setting the function $f$ in the $k$-dimensional nonlinear log-Sobolev inequality to be the product form $f(x)=\prod_{i=1}^{j}f_{1}(x_{i})\prod_{i=j+1}^{k}f_{2}(x_{i})$ for some functions $f_{1}$ and $f_{2}$ defined on $\mathcal{M}$, the lower bound $\breve{\Theta}$ can be asymptotically approached.
Combining (6.12) and (6.13) yields $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha+2\delta)}$. Letting $\delta\downarrow 0$ yields the desired result $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$. ∎
Combining this theorem with Theorem 13 yields that $\underline{\Lambda}(\alpha)=\bar{\Lambda}(\alpha)=\sqrt{\breve{\Theta}(\alpha)},$ completing the proof of our main result, Theorem 1.
Appendix A: Proof of Statement 2 of Theorem 1
Our proof for $\Lambda(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$ still works. We now show that if Condition 2 does not hold, then $\breve{\Theta}(\alpha)=0$ for any $\alpha>0$. We prove this by contradiction.
Suppose that $\breve{\Theta}(\alpha_{0})>0$ for some $\alpha_{0}>0$. Then, $\breve{\Theta}(\alpha)\geq C\alpha$ for $\alpha\geq\alpha_{0}$ where $C=\frac{\breve{\Theta}(\alpha_{0})}{\alpha_{0}}>0$. Thus,
(6.14) $C(D(\mu\|\nu)-\alpha_{0})\leq I(\mu\|\nu),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}).$
Equivalently, for any smooth function $f\in L^{2}(\mathcal{M},\nu)$:
$\int f^{2}\ln f^{2}{\text{\rm d}}\nu-\left(\int f^{2}{\text{\rm d}}\nu\right)\ln\left(\int f^{2}{\text{\rm d}}\nu\right)-\alpha_{0}\int f^{2}{\text{\rm d}}\nu\leq\frac{4}{C}\int|\nabla f|^{2}{\text{\rm d}}\nu$
We now use Herbst’s argument to derive exponential integrability from this log-Sobolev inequality. Let $F(x)=d(x,x_{0})$ be the distance function. By Rademacher’s theorem, $|\nabla F|\leq 1$ almost everywhere. Define the moment generating function for $F$: $\psi(\lambda)=\int e^{\lambda F}{\text{\rm d}}\nu$. We assume $\psi(\lambda)<\infty$ for all $\lambda\geq 0$.
Apply the LSI to the function $f^{2}=e^{\lambda F}$. Then $|\nabla f|^{2}=\frac{\lambda^{2}}{4}e^{\lambda F}|\nabla F|^{2}$. Substituting this into the LSI:
$\int\lambda Fe^{\lambda F}{\text{\rm d}}\nu-\psi(\lambda)\ln\psi(\lambda)-\alpha_{0}\psi(\lambda)\leq\frac{4}{C}\int\frac{\lambda^{2}}{4}e^{\lambda F}|\nabla F|^{2}{\text{\rm d}}\nu.$
Since $|\nabla F|\leq 1$, we have:
$\frac{\lambda\psi^{\prime}(\lambda)-\psi(\lambda)\ln\psi(\lambda)-\alpha_{0}\psi(\lambda)}{\lambda^{2}\psi(\lambda)}\leq\frac{1}{C}.$
Notice that the left-hand side is the derivative of the function $\lambda\mapsto\frac{\ln\psi(\lambda)+\alpha_{0}}{\lambda}$, i.e., $\frac{{\text{\rm d}}}{{\text{\rm d}}\lambda}\left(\frac{\ln\psi(\lambda)+\alpha_{0}}{\lambda}\right)\leq\frac{1}{C}$. Integrating from a constant $\lambda_{0}$ to $\lambda$ yields $\frac{\ln\psi(\lambda)+\alpha_{0}}{\lambda}-C_{1}\leq\frac{\lambda}{C}$ where $C_{1}=\frac{\ln\psi(\lambda_{0})+\alpha_{0}}{\lambda_{0}}$. Thus,
$\psi(\lambda)\leq e^{\lambda^{2}/C+C_{1}\lambda-\alpha_{0}}.$
Note that in the proof above, we assume $\psi(\lambda)<\infty$ for all $\lambda\geq 0$. This is in fact unnecessary since we can obtain the same inequality by substituting $F\leftarrow F_{n}=\min\{F,n\}$ in the proof above and letting $n\to\infty$.
Using Markov’s inequality, we obtain the tail estimate for any $r>0$:
$\nu(F\geq r)=\nu(e^{\lambda F}\geq e^{\lambda r})\leq\frac{\mathbb{E}[e^{\lambda F}]}{e^{\lambda r}}\leq e^{\lambda^{2}/C+C_{1}\lambda-\alpha_{0}-\lambda r}.$
To get the tightest bound, we minimize the exponent with respect to $\lambda$. Setting $\lambda=C(r-C_{1})/2$ gives $\nu(F\geq r)\leq e^{-\frac{C(r-C_{1})^{2}}{4}-\alpha_{0}}$.
We now use this tail estimate to evaluate $\mathbb{E}_{\nu}[e^{\alpha F^{2}}]$. We use the layer-cake representation $\mathbb{E}_{\nu}[e^{\alpha F^{2}}]=\int_{0}^{\infty}\nu(e^{\alpha F^{2}}>t){\text{\rm d}}t$. Let $t=e^{\alpha r^{2}}$, which implies $r=\sqrt{\frac{\ln t}{\alpha}}$. Changing variables $\mathbb{E}_{\nu}[e^{\alpha F^{2}}]=1+\int_{0}^{\infty}\nu(F>r)\cdot 2\alpha re^{\alpha r^{2}}{\text{\rm d}}r$. Substitute the tail bound into the integral:
$\mathbb{E}_{\nu}[e^{\alpha F^{2}}]\leq 1+\int_{0}^{\infty}2e^{-\frac{C(r-C_{1})^{2}}{4}-\alpha_{0}}\cdot 2\alpha re^{\alpha r^{2}}{\text{\rm d}}r=1+4\alpha\int_{0}^{\infty}re^{\alpha r^{2}-\frac{C(r-C_{1})^{2}}{4}-\alpha_{0}}{\text{\rm d}}r.$
Obviously, the integral converges to a finite value for any $\alpha<\frac{C}{4}$. This contradicts with that Condition 2 does not hold. Hence, $\breve{\Theta}(\alpha)=0$ for any $\alpha>0$.
Appendix B: Proof of Theorem 2
Observe that $\{\nu_{m}^{\otimes n}\}$ converges to $\nu^{\otimes n}$ in total-variation distance and in addition $\nu_{m}^{\otimes n}(B)\geq(1-1/m)^{n}\nu^{\otimes n}(B)$ for any Borel set $B$. E. Milman [51, Theorem 6.10] showed that in this case, for all $a\in[0,1]$, $I_{\nu^{\otimes n}}(a)=\lim_{m\to\infty}I_{\nu_{m}^{\otimes n}}(a),$ and consequently, $I_{\nu^{\otimes n}}$ is concave. Let $A$ be an isoperimetric minimizer for $\nu_{m}^{\otimes n}$ in the standard sense. By Theorem 12, for any $r>0$,
$\frac{I_{\nu_{m}^{\otimes n}}(a)}{a}\geq\frac{1}{r}\ln\frac{\nu_{m}^{\otimes n}(A^{r})}{a}\geq\frac{1}{r}\ln\frac{(1-1/m)^{n}\nu^{\otimes n}(A^{r})}{a}\geq\frac{1}{r}\ln\frac{(1-1/m)^{n}\Gamma_{\nu^{\otimes n}}(a,r)}{a}.$
Taking the limit as $m\to\infty$, we obtain $\frac{I_{\nu^{\otimes n}}(a)}{a}\geq\frac{1}{r}\ln\frac{\Gamma_{\nu^{\otimes n}}(a,r)}{a}.$ Note that $\mathrm{RCD}(0,\infty)$ is satisfied by $\nu$, and thus, if Condition 2 is additionally satisfied, then Theorem 8 is valid for $\nu$. The remaining part of the proof is exactly the same as Theorem 1.
As for the other direction, noting that $I_{\nu^{\otimes n}}$ is concave for every $n$ and $(\mathcal{M},d,\nu)$ is a $\mathrm{RCD}(0,\infty)$-space, our first proof given in Section 6 still works for this space, but with the HWI inequality replaced by [6, Theorem 9.3]. Hence, $\Lambda=\sqrt{\breve{\Theta}}$ holds.
If Condition 2 is not satisfied, then $\Lambda\leq\sqrt{\breve{\Theta}}$ still holds, and thus, by the same proof in Appendix A, $\Lambda=\sqrt{\breve{\Theta}}=0$.
Appendix C: Proof of Theorem 3
In this section, we now prove the remaining unproven parts of Theorem 3.
Appendix C-1. Inequality (b)
We first prove Inequality (b) in Theorem 3. Due to Theorem 11, or more precisely, due to (3.1) and (4.2) in [52] but with $K_{CD}\geq 0$, it holds that for any $n$,
$\frac{1}{2}-a\leq I_{n}(a)\int_{0}^{r(a)}\exp\left(\frac{I_{n}(a)}{a}t-\frac{K_{CD}}{2}t^{2}\right){\text{\rm d}}t,$
where $r(a)=\sqrt{\frac{2}{K_{C}}\ln\frac{1}{a}}$. Taking limits as $n\to\infty$, we have
(6.15) $\frac{1}{2}-a\leq I_{\inf}(a)\int_{0}^{r(a)}\exp\left(\frac{I_{\inf}(a)}{a}t-\frac{K_{CD}}{2}t^{2}\right){\text{\rm d}}t.$
By definition, $\sqrt{K_{IS}^{-}}=\liminf_{a\to 0}\frac{I_{\inf}(a)}{a\sqrt{2\ln\frac{1}{a}}}$, which implies that for any $\epsilon>0$, there is a decreasing sequence $a_{k}$ converging to 0 as $k\to\infty$ such that for all sufficiently large $k$,
(6.16) $I_{\inf}(a_{k})\leq a_{k}\sqrt{2(K_{IS}^{-}+\epsilon)\ln\frac{1}{a_{k}}}.$
Substituting $a_{k}$ and the upper bound in (6.16) into (6.15) yields
$\frac{1}{2a_{k}}-1\leq\sqrt{2(K_{IS}^{-}+\epsilon)\ln\frac{1}{a_{k}}}\int_{0}^{r(a_{k})}\exp\left(\sqrt{2(K_{IS}^{-}+\epsilon)\ln\frac{1}{a_{k}}}t-\frac{K_{CD}}{2}t^{2}\right){\text{\rm d}}t.$
Denote $\alpha_{k}=\ln\frac{1}{a_{k}}$, which diverges to $\infty$. Substituting it to the inequality above and performing change of variables, we then have
$\frac{1}{2}e^{\alpha_{k}}-1$ $\leq\sqrt{2(K_{IS}^{-}+\epsilon)}\alpha_{k}\int_{0}^{\sqrt{\frac{2}{K_{C}}}}\exp\left(\sqrt{2(K_{IS}^{-}+\epsilon)}\alpha_{k}t-\frac{K_{CD}}{2}\alpha_{k}t^{2}\right){\text{\rm d}}t.$
Note that $K_{C}\geq K_{IS}^{-}\geq K_{CD}$, which implies $K_{C}\left(K_{IS}^{-}+\epsilon\right)>K_{CD}$. By Laplace’s method, under this condition, the last line above is
$\sim\frac{\sqrt{\left(K_{IS}^{-}+\epsilon\right)K_{C}}}{\sqrt{\left(K_{IS}^{-}+\epsilon\right)K_{C}}-K_{CD}}\exp\left((2\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{C}}}-\frac{K_{CD}}{K_{C}})\alpha_{k}\right)$
as $k\to\infty$. It is asymptotically larger than $\frac{1}{2}e^{\alpha_{k}}-1$ only if $2\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{C}}}-\frac{K_{CD}}{K_{C}}\geq 1$, i.e., $K_{IS}^{-}\geq\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}$ with letting $\epsilon\to 0$.
Appendix C-2. Inequality (k)
Inequality (k) is exactly the following inequality: for any $n\geq 1$,
(6.17) $\liminf_{a\to 0}\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq\sqrt{K_{LS}^{+}}.$
Given any $n$, to prove (6.17), it suffices to show that for any sequence $\{a_{i}\}_{i\in\mathbb{N}}$ of positive numbers that converges to 0, it holds that
(6.18) $\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}\geq\sqrt{K_{LS}^{+}}.$
Let $\alpha>0$. For each $a_{i}$, one can find an integer $n_{i}$ such that $e^{-(n_{i}+1)\alpha}<a_{i}\leq e^{-n_{i}\alpha}$. As $i\to\infty$, it holds that $n_{i}\to\infty$. So, by the monotonicity of $t\mapsto\frac{I_{n}(t)}{t}$ [52, Proposition 3.1] and $t\mapsto\sqrt{\ln\frac{1}{t}}$,
$\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}\geq\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2(n_{i}+1)\alpha}}.$
Taking $\liminf_{i\to 0}$, we obtain
$\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}$ $\geq\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2(n_{i}+1)\alpha}}=\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}$
$\geq\liminf_{i\to 0}\frac{I_{n_{i}}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}\geq\sqrt{\frac{\breve{\Theta}(\alpha)}{2\alpha}},$
where the second inequality is due to that $I_{n}(a)\geq I_{m}(a)$ for all $m\geq n$ and any $a$. Since $\alpha>0$ is arbitrary, we obtain (6.18).
Appendix D: Proof of Theorem 4
We next compare the dimension-free isoperimetric profile $I_{\inf}$ and the infinite-dimensional isoperimetric profile $I_{\infty}$. Recall that the dimension-free isoperimetric profile $I_{\inf}(a):=\lim_{n\to\infty}I_{n}(a)=\inf_{n\geq 1}I_{n}(a),$ and $I_{\infty}$ is the isoperimetric profile for the space $(\mathcal{M}^{\infty},\nu^{\otimes\infty})$.
Theorem 16.
Let $(\mathcal{X},d,\nu)$ be a locally compact Polish probability metric space. For $n\in\mathbb{N}\cup\{\infty\}$, let
(6.19) $I_{n}(a):=\inf_{\mathrm{closed}\,A:\nu_{n}(A)=a}\nu_{n}^{+}(A),\;a\in[0,1]$
be the isoperimetric profile for the $n$-product space $(\mathcal{X}^{n},\nu^{\otimes n})$, where the enlargement of a set is defined under $d_{n}(\mathbf{x},\mathbf{y}):=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$. If $I_{\inf}$ given above is continuous at $a$, then $I_{\inf}(a)=I_{\infty}(a).$
Remark 9.
We restrict the set $A$ in (6.19) to be closed, in order to ensure that $A^{r}\to A$ and hence $\nu_{n}(A^{r})\to\nu_{n}(A)$ as $r\downarrow 0$. This restriction usually does not affect the isoperimetric profile, e.g., for the smooth Riemannian manifold setting with an absolutely continuous reference measure [54].
In our Riemannian manifold setting, under Condition 1, the isoperimetric profiles $I_{n},n\in\mathbb{N}$ are all concave, which implies their pointwise minimum $I_{\inf}$ is concave as well and hence also continuous on $(0,1)$. So, Theorem 16 implies $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in(0,1)$. Moreover, obviously $I_{\inf}(a)=I_{\infty}(a)=0$ for $a\in\{0,1\}$. Thus, $I_{\inf}$ is continuous on $[0,1]$. Theorem 4 is a consequence of Theorem 16. Hence it remains to prove Theorem 16.
Appendix D-1. Proof of Theorem 16
In the following, we also use $\mathcal{X}^{\mathbb{N}}$ to denote $\mathcal{X}^{\infty}$, the countably infinite product of $\mathcal{X}$ with itself. Recall that $\nu_{n}=\nu^{\otimes n}$ with $n\in\mathbb{N}\cup\{\infty\}$.
Observe that $I_{\infty}(a)\leq I_{\inf}(a)$, since for any $n$ and $B\subseteq\mathcal{X}^{n}$, $B\times\mathcal{X}^{\mathbb{N}}$ in $\mathcal{X}^{\mathbb{N}}$ has the same measure and the same boundary as those of $B$ in $\mathcal{X}^{n}$. We next prove that $I_{\infty}(a)\geq I_{\inf}(a)$. We follow similar proof steps for Gaussian measures in [34, Theorem 2.2.4]. We first make the following claim.
Claim 1.
Let $\pi_{n}\colon\mathcal{X}^{\mathbb{N}}\to\mathcal{X}^{n}$ be the projection $\pi_{n}(\mathbf{x})=\pi_{n}(x_{k}:k\in\mathbb{N})=(x_{1},\dots,x_{n})$. Then it holds that: (a) $\nu_{n}=\nu_{\infty}\circ\pi_{n}^{-1}$, (b) if $K\subseteq\mathcal{X}^{\mathbb{N}}$ is compact in the product topology, then $K=\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K))$, (c) $K^{r}:=\bigcup_{\mathbf{x}\in K}B_{r]}(\mathbf{x})$ is compact in the product topology if $K$ is, where $B_{r]}(\mathbf{x})$ is the closed ball with center $\mathbf{x}$ and radius $r$ under the metric $d_{\infty}$.
Proof of Claim 1.
(a) is obvious.
(b) Inclusion $K\subseteq\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))$: By definition, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$ for any $\mathbf{x}\in K$, so $\mathbf{x}\in\pi_{n}^{-1}(\pi_{n}(K))$ for all $n$. Thus, $\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
Inclusion $\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))\subseteq K$: Suppose $\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K))$. Then for each $n$, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$, so there exists $\mathbf{x}^{(n)}\in K$ such that $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$. Since $K$ is compact in the product topology of $\mathcal{X}^{\mathbb{N}}$, it is closed and sequentially compact. The sequence $(\mathbf{x}^{(n)})$ lies in $K$. The projections $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$ imply that the first $n$ coordinates of $\mathbf{x}^{(n)}$ match those of $\mathbf{x}$. This implies $\mathbf{x}^{(n)}\to\mathbf{x}$ in the product topology. Since $K$ is closed, $\mathbf{x}\in K$. Hence, $K=\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
(c) We first note that $d_{\infty}$ might be infinity between two points, and it does not induce the product topology (topology of coordinate-wise convergence) on $\mathcal{X}^{\mathbb{N}}$. In fact, the convergence in $d_{\infty}$ implies the convergence in product topology.
Let $K\subseteq\prod_{i=1}^{\infty}X_{i}$ be compact, and define
$C_{i}:=\pi_{i}(K)\subseteq X_{i},$
which is compact in $X_{i}$ since $\pi_{i}$ is continuous. For any $\mathbf{y}\in K^{r}$, there exists $\mathbf{x}\in K$ such that $\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}\leq r^{2}.$ Hence, for each coordinate $i$, $d(y_{i},C_{i})\leq d(y_{i},x_{i})\leq r,$ because $x_{i}\in C_{i}$. This implies
$y_{i}\in C_{i}^{r}:=B_{r]}(C_{i}):=\{z\in X_{i}:d(z,C_{i})\leq r\},$
which is compact since $\mathcal{X}$ is locally compact. Therefore, $K^{r}\subseteq\prod_{i=1}^{\infty}C_{i}^{r},$ and by Tychonoff’s theorem, $\prod_{i=1}^{\infty}C_{i}^{r}$ is compact in the product topology.
We now prove that $K^{r}$ is closed in the product topology. Fix $\mathbf{x}\in K$. For each $N\in\mathbb{N}$, define
$B_{r]}^{(N)}(\mathbf{x}):=\left\{\mathbf{y}:\sum_{i=1}^{N}d(x_{i},y_{i})^{2}\leq r^{2}\right\}\textrm{ and }K_{N}^{r}:=\bigcup_{\mathbf{x}\in K}B_{r]}^{(N)}(\mathbf{x}).$
Obviously, $K_{N}^{r}$ is non-increasing in $N$, and $K^{r}\subseteq K_{N}^{r}$. So, $K^{r}\subseteq G:=\bigcap_{N=1}^{\infty}K_{N}^{r}$.
On the other hand, let $\mathbf{y}\in G$. It holds that for each $N$, there is $\mathbf{x}^{(N)}\in K$ such that $\sum_{i=1}^{N}d(x_{i}^{(N)},y_{i})^{2}\leq r^{2}$. Since $K$ is compact in the product topology, there is a subsequence $\mathbf{x}^{(N_{k})},k\in\mathbb{N}$ which converges in the product topology to some point $\mathbf{z}\in K$. By the definition of product topology, for each $i$, $d(x_{i}^{(N_{k})},y_{i})\to d(z_{i},y_{i})$ as $k\to\infty$. Given any $m\in\mathbb{N}$, for sufficiently large $k$ such that $N_{k}\geq m$, by triangle inequality and Minkowski inequality, it holds that
$\sum_{i=1}^{m}d(z_{i},y_{i})^{2}\leq\sum_{i=1}^{m}(d(x_{i}^{(N_{k})},y_{i})+d(x_{i}^{(N_{k})},z_{i}))^{2}$
$\qquad\leq\Big(\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},y_{i})^{2}}+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}}\Big)^{2}\leq\Big(r+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}}\Big)^{2},$
Taking limits for both sides as $k\to\infty$, we obtain that $\sum_{i=1}^{m}d(z_{i},y_{i})^{2}\leq r^{2}.$ Letting $m\to\infty$ yields $\sum_{i=1}^{\infty}d(z_{i},y_{i})^{2}\leq r^{2}.$ That is, $\mathbf{y}\in K^{r}$. So, $G\subseteq K^{r}$. Hence, $K^{r}=G=\bigcap_{N=1}^{\infty}K_{N}^{r}$.
By a similar convergence argument, the set $K_{N}^{r}$ is closed in the product topology. Note that $K^{r}=\bigcap_{N=1}^{\infty}K_{N}^{r}$ an intersection of closed sets, hence $K^{r}$ is closed in the product topology. Since $K^{r}$ is closed in the compact set $\prod_{i=1}^{\infty}C_{i}^{r}$, it follows that $K^{r}$ is compact. ∎
We now turn back to prove $I_{\infty}(a)\geq I_{\inf}(a)$. Let $A\subseteq\mathcal{X}^{\mathbb{N}}$ be a compact set. Let $A_{n}=\pi_{n}^{-1}(\pi_{n}(A))$. Then, by the claim above, $A=\lim_{n\to\infty}A_{n}$, and $A^{r}$ is compact as well. By the claim above again, we have $A^{r}=\lim_{n\to\infty}B_{n}$, where $B_{n}=\pi_{n}^{-1}(\pi_{n}(A^{r}))=\pi_{n}^{-1}(\pi_{n}(A)^{r})=A_{n}^{r}$. So, by the continuity of probability measures,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $=\lim_{n\to\infty}\frac{\nu_{\infty}(A_{n}^{r})-\nu_{\infty}(A_{n})}{r}=\lim_{n\to\infty}\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r},$
where $E_{n}:=\pi_{n}(A)$.
Since $f(t):=\nu_{n}(E_{n}^{r})$ is monotone, it is differentiable almost everywhere on $(0,r)$. Hence, by the mean value theorem, there is some $t_{n}\in(0,r)$ such that $f$ is differentiable at $t_{n}$, and moreover, $\frac{f(r)-f(0)}{r}\geq f^{\prime}(t_{n}),$ i.e., $\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r}\geq\nu_{n}^{+}(E_{n}^{t_{n}}).$ Hence,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $\geq\limsup_{n\to\infty}\nu_{n}^{+}(E_{n}^{t_{n}})\geq\limsup_{n\to\infty}I_{n}(\nu_{n}(E_{n}^{t_{n}}))\geq\limsup_{n\to\infty}I_{\inf}(\nu_{n}(E_{n}^{t_{n}})).$
Observe that $\nu_{\infty}(A_{n})=\nu_{n}(E_{n})\leq\nu_{n}(E_{n}^{t_{n}})\leq\nu_{n}(E_{n}^{r})=\nu_{\infty}(B_{n}).$ Taking $n\to\infty$, $\nu_{\infty}(A)\leq\limsup_{n\to\infty}\nu_{n}(E_{n}^{t_{n}})\leq\nu_{\infty}(A^{r}).$ Therefore,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $\geq\inf_{\nu_{\infty}(A)\leq a\leq\nu_{\infty}(A^{r})}I_{\inf}(a).$
Taking $r\downarrow 0$ and by the continuity of $I_{\inf}$ at $a$, $\nu_{\infty}^{+}(A)\geq I_{\inf}(a).$ So, Theorem 16 holds for compact sets.
We now extend the theorem to closed sets. Let $A\subseteq\mathcal{X}^{\mathbb{N}}$ be a closed set. Since $\mathcal{X}^{\mathbb{N}}$ equipped with the product topology is Polish, it follows that $\nu_{\infty}$ is tight. That is, given any $\epsilon>0$, there is a compact set $K_{\epsilon}\subseteq\mathcal{X}^{\mathbb{N}}$ such that $\nu_{\infty}(K_{\epsilon}^{c})\leq\epsilon$. Let $A_{\epsilon}=A\cap K_{\epsilon}$ which is compact. So, $\nu_{\infty}(A_{\epsilon})\leq\nu_{\infty}(A)\leq\nu_{\infty}(A_{\epsilon})+\epsilon$ and $\nu_{\infty}(A_{\epsilon}^{r})\leq\nu_{\infty}(A^{r})\leq\nu_{\infty}(A_{\epsilon}^{r})+\epsilon$. So,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}\geq\frac{\nu_{\infty}(A_{\epsilon}^{r})-\nu_{\infty}(A_{\epsilon})-\epsilon}{r}$
$\qquad\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A_{\epsilon})\leq a\leq\nu_{\infty}(A_{\epsilon}^{r})}I_{\inf}(a)\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A)-\epsilon\leq a\leq\nu_{\infty}(A^{r})}I_{\inf}(a).$
Letting first $\epsilon\to 0$ and then $r\to 0$, by the continuity of $I_{\inf}$ at $a$, we obtain $\nu_{\infty}^{+}(A)\geq I_{\inf}(a).$ Hence, $I_{\infty}(a)\geq I_{\inf}(a)$.
Appendix E: Proof of Theorem 5
By an approximation argument of Otto and Villani [58, Appendix A], without affecting the optimal constant in (1.20), it suffices to consider $f\in W^{1,2}(\mathcal{M})$ satisfying the following smoothness and boundedness conditions: $f$ is bounded away from 0 and $\infty$ on $\mathcal{M}$, and $f$ is smooth and $|\nabla f|^{2}$ is bounded on $\mathcal{M}$. Without loss of generality, we can further assume $\mathbb{E}_{\nu}\left[f\right]=0$ and $\mathbb{E}_{\nu}\left[f^{2}\right]=1$, since otherwise we can normalize $f$ which neither affects the optimal constant in (1.20). Among all this kind of functions, in the following we let $f$ be the approximately optimal function attaining the optimal constant in (1.20), i.e., $\mathbb{E}[|\nabla f|^{2}]\leq(K_{P}+\delta)\cdot\mathrm{Var}_{\nu}(f)=K_{P}+\delta$ for $\delta>0$.
For such $f$, we consider the set $A_{t}=H_{t\sqrt{n}}(f)=\left\{\mathbf{x}:g_{n}(\mathbf{x})\leq t\right\}$, where $g_{n}(\mathbf{x})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f(x_{i})$. We next determine the asymptotics of the volume and the perimeter of this set.
By the coarea formula,
(6.20) $\int_{|g_{n}-t|\leq\delta}|\nabla g_{n}|\,d\nu_{n}=\int_{t-\delta}^{t+\delta}\left(\int_{g_{n}=s}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$
where $\varphi_{n}(\mathbf{x})=e^{-\sum_{i=1}^{n}V(x_{i})}$ is the density of $\nu_{n}$, ${\text{\rm d}}\mathrm{\mathrm{area}}$ is the induced volume form on the level set $\{g_{n}=s\}$, and $|\nabla g_{n}(\mathbf{x})|$ be the norm of the gradient of $g_{n}$ at the point $\mathbf{x}\in\mathcal{M}^{n}$. Thus, by the mean value theorem, the RHS above is larger than or equal to
(6.21) $2\delta\cdot\int_{g_{n}=t_{n}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\geq 2\delta\cdot\nu_{n}^{+}(A_{t_{n}}),$
for some $t_{n}\in[t-\delta,t+\delta]$. Denoting $\mathbf{X}\sim\nu_{n}$, the LHS above is equal to
(6.22) $\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}\times\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right].$
By the central limit theorem, as $n\to\infty$,
(6.23) $\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}\to\Phi(t+\delta)-\Phi(t-\delta).$
We next determine the limit of $\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right]$. Note that $|\nabla g_{n}(\mathbf{X})|^{2}=\frac{1}{n}\sum_{i=1}^{n}|\nabla f|^{2}(X_{i})$. Let $\beta:=\mathbb{E}[|\nabla f|^{2}]\leq K_{P}+\delta.$ By the weak law of large number, $\lim_{n\to\infty}\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\}\to 1.$ By basic probabilistic inequalities, for $\delta>0$,
$\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta,|g_{n}(\mathbf{X})-t|\leq\delta\}$ $\geq\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}-\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|>\delta\},$
which implies
$\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\big||g_{n}(\mathbf{X})-t|\leq\delta\}$ $\geq 1-\frac{\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|>\delta\}}{\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}}.$
By taking limit as $n\to\infty$, $\lim_{n\to\infty}\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\big||g_{n}(\mathbf{X})-t|\leq\delta\}=1.$ Since $|\nabla f|^{2}$ is bounded by assumption (and thus, so is $|\nabla g_{n}|^{2}$), we obtain
(6.24) $\lim_{n\to\infty}\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right]$ $\leq\sqrt{\beta+\delta}\leq\sqrt{K_{P}+\delta}+\delta.$
Combining (6.20), (6.21), (6.22), (6.23), and (6.24) yields
$\limsup_{n\to\infty}\nu_{n}^{+}(A_{t_{n}})\leq\frac{\Phi(t+\delta)-\Phi(t-\delta)}{2\delta}\left(\sqrt{K_{P}+\delta}+\delta\right).$
Letting $\delta\to 0$ yields
(6.25) $\limsup_{\delta\to 0}\limsup_{n\to\infty}\nu_{n}^{+}(A_{t_{n}})\leq\sqrt{K_{P}}\phi(t).$
By the central limit theorem, $g_{n}(\mathbf{X})$ converges in distribution to $Z\sim\mathcal{N}(0,1)$. Thus, the volume $\nu_{n}(A_{t_{n}})$ is asymptotically sandwiched between $\Phi(t-\delta)$ and $\Phi(t+\delta)$ as $n\to\infty$, which implies
(6.26) $\lim_{\delta\to 0}\lim_{n\to\infty}\nu_{n}(A_{t_{n}})=\Phi(t).$
Equations (6.25) and (6.26) implies that there is a sequence of sets $B_{n}\subseteq\mathcal{M}^{n}$ such that $\lim_{n\to\infty}\nu_{n}^{+}(B_{n})\leq\sqrt{K_{P}}\phi(t)$ and $\lim_{n\to\infty}\nu_{n}(B_{n})=\Phi(t)$. By continuity of $I_{\inf}$, it holds that $I_{\inf}(a)\leq\sqrt{K_{P}}\cdot I_{\mathrm{G}}(a)$.
Acknowledgements
This work was supported by the National Key Research and Development Program of China under grant 2023YFA1009604 and the NSFC under grant 62101286.